A Complete Bibliography of Publications in Isis, 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

26 September 2023 Version 1.09

Title word cross-reference

$149.99 [689]. $155 [215]. $29.95 [187]. $34.95 [695]. $40 [188, 694]. $41.95 [692]. $45 [189]. $75 [157]. $80 [157]. p [112].

.} [456, 453].

129 [323]. 1300 [488]. 1300-1700.} [488]. 150 [689]. 1560-1700 [707]. 1600-1800 [491]. 1616 [37]. 1650 [238]. 1650-1750. [238]. 1660 [390]. 1660-1732.} [390]. 1665 [172]. 1700 [107]. 1700.} [488]. 1720 [731]. 1720-92 [391]. 1732.} [390]. 1750 [495, 498, 494, 241]. 1750-1820.} [494]. 1750-1850.} [495]. 1750-1950.} [498, 241]. 1750. [238]. 1784 [446]. 1784-1838.} [446]. 1800 [332, 549, 97]. 1800-2000.} [332]. 1815 [341]. 1815-1920.} [341]. 1820.} [494]. 1834 [555]. 1838.} [446]. 1850 [217, 546, 33]. 1850.} [495]. 1860-1910 [453]. 1860-1914 [740]. 1864 [397]. 1864-1914.} [397]. 1865 [550]. 1870 [20]. 1890 [502]. 1890-1930 [710]. 1890-1940.} [502]. 19 [202, 814, 807, 799]. 1900 [402, 403]. 1900-1926.} [402]. 1900-1940.} [403]. 1900s [520]. 1909 [294]. 1910 [50, 184]. 1911 [16]. 1914 [683, 557]. 1914.} [397]. 1916-1956 [456]. 1920 [245]. 1920.} [341]. 1920s [74, 324]. 1925 [118]. 1926.} [402]. 1930s [521, 520, 74, 126]. 1940 [180]. 1940.} [502, 403]. 1940s [763]. 1940s-1960s [763]. 1945 [646]. 1950 [788, 688]. 1950.} [498, 241]. 1960s [521, 763, 514, 256]. 1966-1989 [313]. 1970 [378]. 1980 [217, 255, 132].

2000 [648, 158]. 2000.} [332]. 2005 [745]. 2015.} [359]. 2018 [63]. 2019 [77].

3 [352].

5 [689].

6 [694, 695]. 62 [308].

7 [692]. 73 [52].

8 [215, 449].

978 [689, 352, 215, 692, 694, 695, 449]. 978-0-19-088293-8 [449]. 978-0-262-54290-6 [694]. 978-1-138-69761-4 [215]. 978-1-4214-3984-6 [695]. 978-1-4724-2039-8 [215]. 978-1-77199-265-7 [692]. 978-3-030-67909-5 [689]. 978-3-319-96835-3 [250]. 978-981-12-0701-3 [352].

Īsā [228].

A. [483]. Aaron [151]. Abayomi [467]. Abbate [750]. ’Abd [385]. Abena [362]. Abi [603]. Abi-Rached [603]. Abigail [217, 111]. Abortion [440]. Abroad [248]. Abstractions [750]. Abundance [521]. Academia [517]. Academie [493]. Academy [85]. accademiche [38]. Access [431]. Acclimatization [176]. Acid [257, 412]. Ackermann [165]. Acknowledgments [798]. Acoustic [713]. Acoustical [419]. Acoustics [761]. across [570, 494, 617]. Action [150]. Active [189]. Activist [300]. Actores [610]. Ada [45]. Adanson [663]. Addams [297]. Addresses [181]. Adelman [240]. Administration [667]. Adrian [45]. Adriana [102]. Adventures [187, 733]. Advice [426]. Advocacy [701]. Aesthetic [731]. Aesthetics [302]. Affair [614]. Affective [223]. Affiliated [526]. Afkhami [244]. Aflāk [385]. Africa [362, 546, 204, 529, 565]. African [56, 481, 467, 576, 663]. after [362, 695, 290]. Afterlives [565, 459]. Aftermath [197]. against [664, 745]. Age [392, 39, 49, 244, 205, 217, 332, 73, 133, 782, 682, 337, 642, 485, 599, 433, 61]. Age.} [485]. Agency [530]. Agents [158, 400, 172]. Ages [287, 229, 543, 726]. Ágnes [633]. Agricultural [576]. Agriculture [622, 693, 680, 389]. Agustí [180]. Ahab [178]. Ahmed [437]. AI [482]. AIDS [807]. Air [507]. Akteure [339]. Aktouarios [542]. Al [23, 385, 486]. al-Aflāk [385]. al-Dīn [486]. al-Idrāk [385]. al-Jabbār [385]. al-Kharaqī [385]. al-Mu’īniyya [486]. Al-Rawi [23]. al-Risāla [486]. al-Tūsi [486]. Alī [228]. Alamos [307]. Albany [157]. Albert [157, 90]. Alberta [692]. Alchemical [634, 439]. Alchemist [99]. Alchemy [286, 634, 488, 636, 573, 164]. Aldous [68]. Alex [638, 656, 15]. Alexander [243, 353, 731, 162, 690, 779, 125]. Alexandra [455]. Alfred [394, 110]. Algol [133]. Algorithmic [410]. Alida [777]. Aligned [425]. Alisha [544]. Alison [286, 243]. Allan [258]. Allen [697]. Alliance [131]. Allied [695]. Allure [46]. Almanac [334]. Alonso [711]. Alps [505]. Álvaro [711]. Amanda [86]. Ambasciano [517]. Ambassadors [774]. Amelia [177]. America [503, 446, 360, 701, 246, 688, 655, 431, 816, 596, 345, 405, 281, 379, 382, 60, 357, 599, 591, 66]. American [647, 559, 245, 554, 121, 188, 49, 69, 448, 645, 788, 744, 247, 61, 52, 451, 550, 713, 468, 248, 710, 513]. Amílcar [576]. Amir [244]. among [137]. Amos [384]. Amy [328]. Ana [610]. Analogy [736]. Analysis [138]. Anatolia [436]. Anatomies [388]. Anatomy [707, 492, 740]. ancienne [25]. Ancient [286, 333, 674, 724, 162, 630, 107, 810, 754]. Anctil [161]. Andra [360]. André [305, 407]. Andrea [237]. Andreas [434, 28, 239, 448, 593]. Andrei [503, 685]. Andrew [215, 227, 114, 186, 751, 645, 382]. Ane [101]. Anecdotal [633]. Angela [111, 42, 224]. Angelica [169]. Angola [785]. Animal [123, 678, 185, 242, 228, 451]. Animals [727, 759, 169, 111, 16]. Ankeny [602]. Anker [411]. Ann [26]. Anna [19, 539, 444, 280, 78, 47, 636]. Anne [334, 503, 775, 104, 339]. Annotated [46]. Annotations [28]. Annual [77, 771, 580]. Answers [419]. Antarctic [521]. Antarctica [296]. Anthology [497]. Anthony [156]. Anthropocene [529, 485, 530, 527, 329, 454, 643]. Anthropology [664, 195, 693, 630, 748]. Anti [664, 95, 37, 688]. Anti-Asian [688]. Anti-Copernican [37]. Anti-Mannerism [95]. Anti-Racism [664]. Antibiotic [64]. antike [90]. Antiquity [480, 675, 164, 552]. Antoine [447, 451]. Antonio [410]. Anxieties [813]. Anxiety [645, 597, 51]. Anxious [177]. Anya [190]. Apocalyptic [61]. Apollo [708]. app [352, 692]. Appetite [498]. Approach [147, 145, 630]. Approaches [526, 801]. Arabic [384]. Archaeological [630]. Archaeologies [353]. Archaeology [26, 107]. Archäologie [630]. Archimedes [142]. Architectural [353]. Architecture [416, 332, 507, 514]. Archives [323]. Archiving [714]. Arctic [499]. Argentina [453]. Argument [736]. Aristotle [334, 540]. Arizona [137]. Ark [36]. Arkansas [502]. Arleen [395]. Armstrong [110]. Arne [402]. Aronova [608]. Arredondo [291]. Arribas [165]. Arrizabalaga [215]. Art [444, 159, 416, 775, 288, 685, 101, 94, 169, 256, 623, 286]. Arte [777, 711]. Artefacts [630]. Arthur [288, 128]. Artificial [550, 736]. Artisan [32, 447]. Artisanal [524, 464, 148]. Artist [235]. Artists [159]. Arts [652]. Arunabh [356]. Asare [362]. Ascent [628]. ‘Asfūriyyeh [603]. Ashley [453].

Ashworth [597]. Asia [814, 813, 812, 787, 224]. Asian [688, 430]. Asimov [304]. Aspaas [391]. Aspiration [321]. Assembling [347]. Assessment [428]. Association [370]. Astral [541]. Astrolabes [165]. Astrologer [232]. Astrological [445, 167]. Astrology [425, 606]. Astronomer [738]. Astronomical [383, 229, 615, 728]. Astronomy [184, 229, 438, 225]. Asylum [368, 48]. Atlantic [635, 546, 491, 404, 494, 234, 369, 270, 663]. Atlas [690]. Atmosphere [458]. Atmospheric [742]. Atom [282, 655]. Atomic [307, 408, 187, 362, 747]. Atomistic [157]. Atoms [442]. Atri [54]. Aubin [118]. Audiences [759]. Audra [65]. Audubon [733]. Augurello [634]. Augustine [252]. Augustus [467]. Aurality [455]. Aurelio [634]. Australia [176, 41]. Australian [295]. Austria [298]. Authorities [595]. Authority [681, 382, 544]. Authorship [15]. Autistic [611]. Autobiography [130, 167]. Automobiles [124]. autre [113]. available [157, 215, 187, 157, 188, 250, 189]. Avicenna [384]. Axelle [29]. Axolotl [397]. Ayres [399].

B [601, 57, 360, 129, 44, 506, 33]. B. [426]. B.C.E [419]. Babylonia [163]. Babylonian [383]. Bachelard [157, 157]. Bacon [386, 422]. Bacterial [560]. Bacteriology [73, 710]. Badia [410]. Baggio [630]. Baigent [407]. Bailey [247]. Bailly [34]. Balance [206]. Ball [537]. Ballistic [322]. Baltic [734]. Baltimore [695]. Banchetti [442]. Banchetti-Robino [442]. Banerjee [415]. Bangham [651]. Banned [37]. Banu [191]. Barak [351]. Baranski [789]. Barba [711]. Barbara [216, 153, 639]. Barber [507]. Barcia [650]. Barnett [290]. Barr [632]. Barré [614]. Barré-Lebiez [614]. Barres [130]. Bárth. [496]. Basic [194]. Bass [775]. Bate [167]. Battimelli [562]. Battle [697, 544]. Battles [561]. Bauer [291]. Baum [119]. Be [646, 7, 214, 786]. Bear [277]. Beasts [242]. Beauty [75, 264, 648]. Became [411, 219, 46]. Become [538]. Bedeviled [725]. Bedlam [682]. Bedor [703]. Beeley [490]. before [108, 507, 555]. Begleitband [630]. behind [587, 691, 746, 361]. Beitrag [47]. Beiträge [359]. Belgian [740]. Belgium [343]. Belhoste [392]. Beliefs [489, 61]. Bellamy [600]. Ben [130]. beneath [525]. Bengal [646]. Benjamin [690, 490, 553, 174, 337]. Bennett [27, 400]. Benson [435]. Benyovsky [378]. Beretta [361]. Berlin [689, 302]. Bernard [630]. Bernstein [190]. Bert [648]. Bertolacci [384]. Bertoloni [170]. Besky [354]. Bessels [52]. Bettina [340]. Betts [115]. Between [531, 321, 87, 779, 163]. Beyond [791, 378, 699, 692, 282]. Bharat [604]. Bian [441]. Bias [410]. bibl [187, 157, 188, 352, 250, 449, 215, 692, 189, 694, 695]. Bibliographical [817]. Bibliography [260, 461, 659]. Bibliothèque [156]. Big [134, 653, 206, 755]. Bikes [686]. Bildungspolitik [43]. Bile [227]. Bill [555]. Binzley [87]. Biobibliographical [407]. Biochemistry [139, 139]. Biocultures [306]. Biodiversity [579]. Biographie [401]. Biographies [250]. Biography [414, 413]. Biological [416]. Biologie [132, 50, 359]. Biologische [57]. Biology [602, 57, 281, 134, 723, 185]. Biometry [664]. Biopolitics [191, 242]. Biotechnology [310]. Biotic [688]. Bird [159]. Birds [247]. Birdsong [185]. Birth [405, 779, 237]. Birthed [232]. Bisbee [68]. Bittel [218]. Black [565, 227, 576]. Blades [267]. Blay [330]. Blind [679]. Blood [109, 636, 651, 739]. Bloomers [686]. Board [74]. Bod [772]. Boddice [557]. Bodies [230, 683, 667, 554]. Body [168, 108, 387, 669, 675, 48]. Bohemia [303]. Bohm [67]. Bologna [586]. Bolshevik [608]. Boltzmann [51]. Bomb [307, 747]. Bonea [177]. Bones [666]. Book [628, 647, 304, 792, 638, 221, 85, 776, 741, 102, 539, 161, 297, 495, 487, 158, 752, 213, 390, 175, 232, 81, 590, 509, 689, 294, 783, 516, 29, 410, 156, 111, 601, 286, 444, 38, 241, 698, 635, 159, 686, 559, 551, 596, 174, 549, 123, 501, 778, 31, 345, 743, 395, 238, 96, 734, 27, 732, 293, 21, 126, 168, 84, 435, 306, 187, 656, 602, 291, 585, 16, 730, 458, 157, 392, 550, 637, 588, 598, 82, 702, 738, 230, 784, 305, 434, 484, 124, 22, 503, 122, 347, 302, 583, 789, 245]. Book [536, 677, 329, 540, 483, 114, 508, 748, 350, 773, 502, 416, 678, 295, 498, 699, 646, 186, 393, 134, 775, 55, 118, 554, 130, 86, 396, 165, 489, 731, 629, 793, 45, 504, 333, 67, 28, 790, 121, 480, 116, 20, 649, 335, 385, 486, 657, 50, 64, 547, 177, 252, 57, 298, 83, 397, 188, 216, 352, 182, 606, 310, 389, 399, 736, 131, 172, 566, 173, 285, 703, 441, 39, 89, 35, 179, 243, 92, 584, 445, 438, 296, 499, 358, 359, 171, 401, 254, 414, 127, 610, 307, 751, 218, 412, 109, 348]. Book [595, 128, 653, 454, 190, 278, 355, 184, 654, 605, 49, 452, 226, 442, 69, 633, 25, 448, 40, 250, 93, 288, 257, 772, 280, 449, 120, 299, 383, 104, 496, 301, 697, 279, 239, 386, 42, 244, 645, 753, 437, 723, 565, 215, 692, 249, 362, 62, 791, 742, 108, 253, 634, 674, 220, 726, 26, 727, 447, 183, 593, 683, 180, 185, 542, 283, 346, 349, 586, 733, 380, 745, 643, 788, 515, 189, 361, 492, 729, 340, 339, 564, 562, 545, 400, 740, 409, 56, 546, 231, 217, 103, 541, 632, 780, 693, 413, 737, 459]. Book [511, 747, 455, 287, 641, 59, 510, 353, 507, 440, 680, 405, 609, 436, 481, 402, 387, 744, 247, 415, 133, 68, 513, 456, 281, 408, 603, 592, 491, 644, 724, 685, 328, 432, 557, 681, 457, 117, 587, 63, 600, 248, 101, 277, 229, 99, 342, 87, 631, 34, 94, 98, 176, 292, 560, 675, 110, 394, 594, 181, 162, 227, 694, 407, 47, 17, 32, 439, 488, 636, 112, 446, 691, 384, 258, 782, 191, 125, 23, 537, 500, 360, 308, 651, 556, 284, 19, 682, 233, 97, 337, 24, 354, 251, 582, 754, 95]. Book [391, 336, 746, 132, 555, 344, 411, 18, 388, 282, 543, 563, 363, 701, 630, 37, 30, 242, 338, 493, 41, 453, 289, 379, 240, 129, 786, 91, 164, 222, 642, 497, 51, 728, 228, 382, 219, 482, 695, 166, 443, 749, 43, 170, 589, 60, 106, 44, 735, 403, 290, 485, 113, 777, 46, 357, 36, 781, 404, 687, 787, 107, 450, 246, 398, 794, 236, 225, 494, 684, 54, 178, 639, 650, 548, 167, 690, 505, 558, 160, 58, 163, 750, 100, 330, 725, 223, 331, 696, 224, 700, 679, 538, 774, 341, 608]. Book [544, 739, 599, 90, 604, 607, 506, 378, 105, 652, 48, 115, 169, 517, 235, 343, 785, 591, 88, 597, 33, 640, 514, 61, 779, 53, 356, 66, 52, 451, 65, 309, 512, 381, 755, 490, 688, 300, 234, 303, 256, 655, 561, 406, 119, 351, 676, 433, 553, 552, 237, 611]. Books [187, 490]. Boole [181]. Boos [88]. Border [570]. Borderlands [33, 687]. Borders [688]. Born [402, 432]. Born-Digital [432]. Børsen [509]. Botanical [452]. Botanik [340]. Botany [741, 149, 466, 676]. Boundaries [549, 773, 430, 378]. Bouras [216, 542]. Bouras-Vallianatos [216, 542]. Bowen [502]. Bowie [89]. Bowler [304]. Box [737]. Boy [174]. Boyle [442]. Braider [173]. Brain [264]. Brandt [681]. Bravo [160]. Brazilian [10]. Bredekamp [95]. Breed [109]. Breeding [727, 575, 648]. Breen [337]. Brendan [181]. Brent [107]. Bresalier [111]. Brescius [782]. Bretelle [24]. Bretelle-Establet [24]. Brett [400]. Breyfogle [44]. Brian [409]. Bridging [332]. Brightening [139]. Brill [216]. Bringing [147]. Britain [399, 459, 548, 759, 784, 177, 243, 109, 595, 406]. British [294, 238, 40, 683, 103, 491, 308, 242, 599]. Brockmann [446]. Broecke [167]. Broglio [242]. Brooks [232]. Brothers [782]. Brotons [730]. Brown [783, 129]. Brownian [605]. Bruce [439]. Bruno [94, 392, 134, 721]. Bruton [55]. Buberl [448]. Bucher [677]. Buddhism [497, 283]. Buell [772]. Buffon [292]. Build [307]. Built [253]. Burden [242]. Bureaucratizing [569]. Burnett [165]. Burton [492, 696]. Buss [61]. Bwamba [4]. Byrne [540]. Byzantine [27, 542, 573].

C [157, 637, 39, 120, 409, 780, 457, 342, 129, 60, 100, 20, 542]. C. [91, 497]. Cable [404]. Cabral [576]. Cabré [777]. Cage [781]. Cahan [213]. Cain [357]. Caitlin [691]. Calcutta [54]. Calendars [726, 726]. California [767]. Cambridge [674, 162, 694]. Came [445]. Cameron [33]. Camille [743]. Campaign [745, 755]. Canada [39]. Canadelli [361]. Canal [700]. Canales [725]. Cancer [188, 415, 749]. Cannabis [594]. Can’t [410]. Canton [524]. Capital [392]. Capture [451]. Carbon [251, 787]. Carbon-14 [251]. Carbonization [351]. Cares [204]. Caribbean [308, 431, 763]. Carl [601]. Carla [218]. Carlos [393, 777, 167]. Carmen [689]. Carnegie [248]. Carolina [78]. Caroline [433]. Carolyn [485]. Carson [68]. Cartesianism [237]. Cartographic [335]. cartographie [29]. Cartography [734, 62]. Case [814, 352, 301, 737]. Cassidy [111]. Cast [713]. Catalogue [115]. Catastrophic [454]. Cateforis [256]. Caterina [606]. Catharina [514]. Cathedral [728]. Catherine [734, 246]. Catholic [779]. Causality [67]. Causation [230, 141]. Cause [188]. Cave [482]. CE [217]. Cécile [630]. Ceglia [387]. Celebrity [754]. Celestial [385]. celik [353]. Cellar [59]. Celsus [90]. Censorship [545]. Census [28]. Center [72]. Centering [209]. Central [446]. centrales [610]. Centuries [288, 617, 445]. Century [776, 524, 590, 635, 596, 784, 393, 177, 389, 243, 595, 448, 496, 622, 103, 432, 594, 650, 652, 343, 597, 616, 469, 664, 271, 713, 614, 419, 523, 369, 760, 553, 102, 444, 499, 414, 42, 753, 249, 593, 349, 361, 513, 685, 15, 631, 450, 548]. Certainty [287]. Ceruzzi [694]. Cesarean [121]. Chabás [229]. Chadarevian [657]. Chair [738]. Chakanetsa [56]. Chakrabarti [552]. Chakrabarty [485]. Challenges [382]. Cham [250]. Chance [67]. Change [158, 254, 485, 92, 400]. Changed [539, 122, 737]. Changing [143]. Channell [219]. Chantal [172, 172]. Character [98]. Characteristic [112]. Charles [174, 165, 46, 450, 427, 394]. Chassagnette [29]. Chastain [360]. Che [224]. Chemical [414, 442, 120, 780, 144]. Chemicals [158]. Chemistry [509, 689, 504, 144, 590]. Chernobyl [459]. Cheryl [786].

Chicago [188, 189, 194]. Chicago/London [188, 189]. Chicken [317]. Chief [52]. Children [140]. Chile [615]. China [516, 403, 419, 74, 6, 622, 269, 563, 620, 621, 667, 572, 741, 441, 630, 356, 119]. Chinese [16, 279, 269]. Chinmay [642]. Cholera [563, 244]. Chomsky [697]. Chris [67]. Christ [445]. Christa [500]. Christer [281]. Christian [583, 397, 388]. Christina [682]. Christine [218, 700]. Christoph [733]. Christopher [49, 540, 45, 173]. Christopoulos [440]. Chromosomes [657]. Chronometers [115]. Chrysopoeia [634]. Church [761]. Chymical [442]. Chymistry [493]. Ciccotti [562]. Ciencia [777, 585]. Cigarette [653]. Cinchona [494]. Cinematographica [325, 322]. Cinematography [318]. circa [465, 525, 466, 267]. Circle [615, 597]. Circuit [599]. Circulation [366, 617, 19]. Citizens [453]. City [348, 180]. Civil [396, 409]. Civilization [466]. Claes [740]. Claire [781]. Clapperton [56]. Clara [611]. Clarke [308, 141]. Class [423, 632]. Classic [422]. Classification [157]. Classifying [270]. Classroom [722, 645]. Claudia [775, 53]. Cleft [465]. Clements [69]. Climate [509, 583, 526, 254, 742, 763, 507, 485, 175, 39, 643]. Clinic [569, 140]. Clips [323]. Clive [479]. Clock [728]. Closing [619]. cloth [689, 187, 157, 250, 449, 215, 189]. Co [149]. Co-teaching [149]. Coal [351]. Coast [269]. Coastal [733]. Cocain [47]. Coddington [396]. Coen [175]. Coercion [548]. Coghe [785]. Cohen [95]. Coinfection [803]. Cold [296, 128, 360, 357, 404, 608, 65, 324]. Colder [299]. Cole [467]. Collaboration [367, 105, 256, 751]. Collard [394]. Collecting [102, 76, 41, 134, 582]. Collection [490, 432]. Collections [156, 49]. Collins [166]. Colonial [646, 649, 39, 346, 176, 344, 41, 785, 640, 210]. Colonialism [241]. Color [345]. Come [206]. comete [38, 172]. Coming [206, 547]. Commemorative [689]. Commentary [634, 91]. Commerce [548]. Commercial [366]. Committee [790]. Commodification [609, 676]. Commodities [763]. Common [609, 770]. Communication [645, 706]. Communications [699]. Communities [784, 120]. Community [74]. Comorbidity [803]. Companion [216, 674, 394]. Comparison [196]. Compatibility [428]. Compendium [383]. Competitions [75]. Compiler [168]. Completed [275]. Complicated [327]. Compound [724]. Comprehension [385]. Comprehensive [385]. Compromise [623]. Computational [229]. Computer [45, 133, 562]. Computers [410]. Computing [750, 694]. Concept [271, 82]. Conceptual [170]. Conceptualizing [101]. Concerns [419]. Conchophilia [775]. Conclusion [819]. Conclusions [314]. Condition [451]. Conditioning [507]. Condon [60]. Confirmed [117]. Conflict [342]. Conflicting [266]. Congo [570]. Conjecture [616]. Connecting [270]. Connection [556]. Connor [17]. Conscious [763]. Consciousness [423, 547]. Consequence [121]. Conservation [255]. Conserved [182]. Constant [300]. Constructing [297, 195]. Construction [75]. Constructions [637]. Contagion [244]. Contagious [188]. Contemporary [157, 793, 190, 497]. Contents [259, 797, 460, 658]. Contests [75]. Context [215, 402]. Contexts [434].

Continued [181]. Continuity [491]. Contrast [196]. Control [779]. Controversial [701]. Conversations [336]. Cooking [464]. Cool [653]. Cooley [727]. Cooperation [505, 309]. Cooperative [60]. Copernican [93, 37]. Copernicus [37]. Cornel [238]. Corpse [564]. Corpses [387, 740]. Corpus [285]. Correspondance [172, 172]. Correspondence [710, 182]. Corrupts [423]. Cosmetics [403]. cosmología [730]. Cosmological [419]. Cosmology [385, 83, 384]. Cosmos [283, 668]. Costa [394]. Cotta [34]. Coulborn [624]. Count [356]. Country [55]. cour [172]. Courageous [194]. Course [146, 194]. Court [169]. Courtenay [598]. Courtney [596]. Courts [588]. Cover [1, 70, 135, 192, 261, 311, 364, 417, 462, 518, 567, 612, 660, 704, 756, 795]. Covered [456]. COVID [199, 202, 814, 807, 799]. COVID-19 [202, 814, 807, 799]. Cowles [501]. Craciun [102]. Craft [792, 377]. Crafting [377]. Cranial [428]. Crash [146, 323]. Crawford [234]. Crawshaw [378]. Creating [4, 569, 283]. Creation [597, 46]. Creators [46]. Creatures [161, 357]. Cremona [182]. Crime [596]. Cristiano [32]. Critical [157, 82, 789, 634, 711, 309, 801]. Critique [330]. critiques [172]. Cross [291, 163]. Cross-Cultural [291, 163]. Crossroads [696]. Cruelty [759]. Cruz [777]. Csiszar [15]. Cultural [260, 461, 291, 738, 659, 248, 163]. Culture [441, 128, 466, 794, 169, 343, 678, 278, 160]. Cultures [732, 143, 165]. Cuneiform [23]. Cunningham [215]. Cupers [514]. Curanderos [687]. Cure [604]. Cures [380]. Curiosity [775]. Curious [236, 102]. Current [260, 461, 654, 659]. Curriculum [149]. Curtis [449]. Cut [700]. Cycle [686]. Cycles [265]. Cyclone [571]. Cynthia [68].

D [22, 359, 683, 332, 754, 129, 303]. Dīn [486]. Dag [384]. Dagmar [16]. Dahlia [40]. Dangerous [746]. Dangers [269]. Daniel [783, 217, 507, 117, 28, 496]. Daniela [301]. Danielle [412]. dans [29]. Danysz [366]. Darab [633]. D’Arcy [296]. Daring [616]. Dark [550, 525]. Darker [479]. Darwin [501, 454, 680, 555, 655, 80, 736, 449, 452]. Darwinian [614]. Darwinism [614]. Dasgupta [133]. Data [656, 134, 76]. Daughter [611]. Daughters [444, 793]. David [402, 394, 19, 46, 167, 213, 27, 118, 67, 179, 454, 541, 594, 219, 166, 105, 256, 553]. Davids [637]. Dawn [693, 32]. Day [480]. DDR [359]. Dead [41, 667, 740]. Deadly [221, 306]. Deafness [380]. Dean [456]. Dear [536]. Death [190, 346, 415]. Deb [294]. Debate [566]. Debates [348]. Debating [770]. Deborah [175, 209]. Debye [402]. Decisions [794]. Decline [709]. Decolonization [576]. Decree [37]. Deep [702, 697, 515, 456, 700]. Defense [557]. Defining [773]. Degenerations [525]. del [730, 38]. Delhi [415]. Delivering [710]. della [584]. Demands [526]. Democracy [297, 405]. Demographics [76]. Demography [108, 785]. Demon [650]. Demons [725]. Dengue [73]. Dennis [310]. Dentinger [111]. Department [182]. Dependencies [416]. Depends [215]. Depicting [269]. Depiction [711]. Depictions [283]. Deprivation [315]. Derr [344]. Descriptive [28].

Design [507, 353]. Destined [246]. Destiny [20, 606]. Determinants [632]. Development [80, 456, 140, 308, 404]. Devriese [675]. Dewey [501]. DGGTB [359]. Diabetes [12, 395]. Diagnosing [502]. Diagnosis [415]. Diana [487]. Diarmid [599]. Dick [750, 178]. Dickson [177]. Dictionary [168]. Didier [405]. Dietz [340]. Difference [42, 103]. Diffusion [614]. Digging [389]. Digital [703, 377, 430, 432, 433]. Dihal [482]. Dijksterhuis [637, 434]. Dillon [482]. Dinge [339]. Dinosaur [347, 248]. Dinosaurierfragmente [63]. Dinosaurs [691]. Dipesh [485]. Diplodocus [248]. Diplomacy [131]. Directors [479]. Disabilities [701]. Disability [679]. Disalienation [743]. Disaster [9, 459]. Disasters [10]. Disciplines [526, 617]. Discontents [498, 186]. discours [447]. Discover [296]. Discoveries [551]. Discovery [34, 433, 537]. Discussion [76]. Disease [522, 22, 279, 215, 103, 344, 650, 558, 209, 801, 395]. Diseases [206, 801]. Disenchantment [791]. Disharmony [774]. Disorders [607]. Disparities [632]. Displaying [361]. Distinguished [8]. Disturbed [578]. Diverse [719]. Diversity [732, 454]. Divide [31]. Divided [523]. Dixon [236]. DNA [754]. Do [512]. Doctor [752, 786]. Doctors [48]. Documents [352, 254]. Dodds [778]. Dodman [221]. Dog [109]. Dogs [245]. Dolan [389]. Dolezel [239]. Dolores [252]. Dome [140]. Domenico [170]. Domesticity [236]. Don [687]. Donahue [691]. Donald [632]. Donato [106]. Don’t [512]. Dooley [181]. Dora [584, 301]. Dorn [408]. Dorner [548]. Dose [215]. dossier [172]. Doubt [554, 117]. Dougal [98]. Douglas [695]. Dove [362]. Down [333]. Dracone [91]. Drawing [5, 95]. Drug [337]. Drugs [430, 594, 544, 234]. Drunk [522]. Dublin [348]. Dunn [139]. Duplicating [320]. Durba [649]. during [753, 208, 668]. Durrenberger [693]. Dutch [241, 637, 434, 648, 237]. Duval [256]. Dwaipayan [415]. Dyck [68]. Dynasty [623].

E-book [187, 250, 215, 189, 157, 188]. E. [693, 164, 781]. Ear [138]. Earliest [16]. Early [765, 96, 291, 114, 678, 775, 489, 764, 768, 335, 547, 173, 441, 35, 215, 249, 492, 729, 545, 440, 387, 5, 681, 101, 98, 269, 270, 711, 95, 338, 368, 164, 589, 106, 290, 575, 679, 714, 607, 148, 591, 356, 719, 490, 234, 572, 767, 425, 769, 622, 676]. Earth [584, 142, 531, 525, 530, 751, 253]. Earthquakes [489]. Earwitness [8]. East [284, 589, 787, 696, 224, 351, 423, 812, 333, 603]. Eat [498]. Eclipse [117]. Eco [765]. Eco-Prospecting [765]. Ecological [4, 400]. Ecologies [618]. Ecology [44, 600]. Economy [344]. Ecosystem [654]. Ed [46, 293, 588, 677, 67, 57, 182, 172, 89, 35, 254, 127, 120, 386, 249, 674, 726, 183, 675, 181, 240, 164, 497, 166, 106, 44, 90, 88]. Edad [777]. Eddy [46]. Edges [803]. Edinburgh [555, 778]. Editing [566]. Edition [634, 91, 228]. Editions [28]. Editor [534, 662, 274, 375, 154, 376, 14]. Editors [672, 198, 3, 477, 532]. Edmonton [692]. Edney [62]. eds [85, 102, 158, 509, 689, 783, 96, 732, 602, 291, 16, 637, 434, 484, 583, 55, 165, 333, 480, 486, 83, 216, 358, 359, 414, 218, 128, 288, 301, 239, 42, 215, 692, 742, 180, 733, 361, 339, 400, 409, 217, 455, 353, 68, 281, 87, 34, 394, 162, 227, 407, 384, 360, 556, 284, 19, 24, 582, 630, 37, 338, 129, 482, 777, 46, 450, 167, 750, 224, 378, 514, 490, 234, 256, 433, 553]. Eduard [212]. Education [147, 350, 145, 150, 586, 146, 710, 144, 324]. Edward [595, 60]. Edwardian [399]. Effects [267]. Egdell [55]. Egel [386]. Eggs [318]. Egypt [758, 344]. Egyptian [286]. eHarmony [427]. Ehlers [306]. Eighteenth [102, 524, 393, 445, 496, 103, 432, 548, 616, 271]. Eighteenth-Century [524, 393, 496, 103, 432, 616, 271]. Eijk [227]. Einstein [117, 58, 820, 303]. Ekerholm [281]. Ekmeleddin [735]. Elaine [218]. Elder [633]. Electricity [646]. Electrons [404]. Elements [537, 320]. Elena [361, 608]. Elisa [630]. Elisabeth [158]. Elise [696]. Elizabeth [55, 125, 453, 498, 554, 407, 754]. Ellen [108]. Elliot [680]. Elliott [502]. Elmar [43]. Eloge [670, 151, 626, 625, 671, 624, 581, 326, 374, 152, 153, 720, 721, 212]. Else [183]. Elusive [698, 506]. Elwick [644]. Emanuel [392]. Embodiments [750]. Embrace [766]. Embryological [317]. Embryos [317]. Emergence [266, 801, 442]. Emergency [526, 300]. Emergent [369]. Emerging [206, 801]. Emil [52]. Emily [748, 119]. Emma [257, 703]. Emmanuel [405]. Emotion [209, 221]. Emotions [505]. Empire [221, 175, 294, 635, 265, 39, 438, 726, 683, 400, 594, 32, 782, 308, 403, 684, 351, 240, 103]. Empires [582, 35]. Empty [641]. Encounter [663]. Encounters [102, 546]. Encouraging [150]. Encyclopaedia [325, 322]. Encyclopedia [631]. Encyclopedic [633]. End [636, 308, 608]. Endangered [503]. Ended [755]. Ends [391]. Enduring [46, 415]. Energy [787]. Engagement [573]. Engine [105]. Engineering [646, 219, 509]. Engineers [46]. England [98, 590, 230, 547, 558, 597]. English [168, 634, 488, 676, 800]. English-Language [800]. Engraver [157]. Enigma [99]. Enlightenment [778, 678, 680, 446, 391, 392, 682, 100]. Entangled [37]. Entanglement [448]. Enterprise [782]. Entomologists [784]. Entrepreneurship [646]. Entropy [51]. Enumerative [266]. Environment [647, 601, 69, 253, 360, 344, 290]. Environmental [255, 131, 176, 411]. Environmentalisms [435]. Environments [435, 44, 400]. Epidemic [366, 803, 808, 812, 201]. Epidemics [817, 809]. Epidemiology [618, 800, 558]. Epistemic [571, 96, 448, 197, 249, 266, 712]. Epistemological [801]. Epistemologies [424, 148]. epistemologischen [132]. época [730]. Epstein [698]. Equation [51]. Era [64, 788, 61, 433]. Erasmo [352]. Erasmus [680]. Erdöl [690]. Erdur [132]. Eric [82, 283, 51]. Erika [68, 357]. Erling [19]. Erna [692]. Ernst [158, 762]. Erratum [627]. Esch [630]. Esha [223]. espace [29]. española [585]. españolas [610]. Espitia [393]. Essay [157, 817, 15]. Essays [22]. Establet [24]. Establishing [402]. Esteban [285]. Ether [249]. Ethics [297, 306, 197, 288, 509, 790]. Ethik [285]. Ethnicity [632]. Ethology [320]. Etienne [435]. Euclidean [468]. Eugenia [403]. Eugenic [305]. Eugenics [692, 398]. EUR [250]. Eurasian [44]. Euro [713]. Europe [489, 108, 282, 774, 739, 378, 490, 271, 215, 420, 588, 775, 335, 173, 729, 391, 338, 290]. European [213, 215, 387, 631, 755, 815, 707]. Eva [582, 239].

Eva-Maria [582]. Evangelists [349]. Evans [566]. Evelyn [236, 236]. Evelyne [25]. Everest [69]. Everyday [698]. Evidence [314, 387]. Evolution [647, 501, 329, 6, 120, 779, 555, 293]. Evolutionary [297]. Evolutionism [6]. Evolving [780]. Ewald [619]. Ewen [89]. Examination [449]. Examinations [644]. Exchange [217]. Exemplary [325]. Exercise [181]. Exhibit [361]. Exhibitions [295]. Existential [61]. Exorcist [496]. Exotic [169]. Expanses [18]. Expectations [711]. Expedition [63, 52, 69]. Expeditions [49]. Experience [173, 513]. Experiment [4, 761, 464, 171]. Experimental [4, 57, 369, 557, 488, 443, 173]. expérimentateur [447]. Experimentation [197, 305]. Experimenting [549]. Experiments [22, 134, 710, 746, 608, 544]. Expert [468]. Expertise [255, 360]. Experts [124, 308]. Exploitation [695]. Exploration [635, 299]. Explorations [19, 499]. Explorers [702]. Export [524]. Extant [27]. Extinction [454]. Extraordinary [686, 250, 125]. Extreme [69, 299]. Eye [288].

F [251, 219]. Fabbri [37]. Fabrica [28]. Face [421]. Faces [396]. Facility [755]. Faflak [293]. Fagan [81, 723]. Fairs [128, 361]. Faith [598, 246]. Fake [410]. Fakes [630]. Falk [543]. Fälschungen [630]. Famine [579]. Fang [563]. Fantasies [17]. Fantasmas [585]. Far [589]. Farber [670]. Farmlands [755]. Farouk [23]. Farquhar [516]. Farrell [68]. Fascism [664]. Fate [398]. Fateme [486]. Father [506]. Favino [37]. FBI [129]. Fear [504]. Federica [37]. Fee [277]. Feeling [505]. Feijoo [426]. Feingold [338]. Feke [226]. Feldtkeller [448]. Felipe [436]. Fellowship [399]. Feminist [426]. Fermentation [652]. Ferreiro [332]. Festa [678]. Fever [73, 650, 370, 558]. Feynman [250]. fi [385]. Fiction [678]. Fieguth [358]. Field [470, 123, 469, 4, 570, 473, 195, 475, 355, 471, 281]. Fields [474, 469, 578]. Fieldwork [469]. Fight [504, 408]. Fighting [650]. figs [250, 215]. Files [129]. Film [325, 759, 321, 316, 324, 316]. Filming [318]. Films [317, 322]. Filter [197]. Filth [558]. Fin [739]. Final [246]. Financialization [310]. Finding [209, 93]. Findlen [35]. Finn [46]. Finnegan [599]. Finzi [301]. Fire [120, 780, 99, 488, 382]. Fireball [479]. Firm [489]. First [385, 414, 307, 108, 631, 735, 458, 419, 617]. First-Century [419]. Fischer [297]. Fishmeal [654]. Fiss [645]. Fleetwood [684]. Fleming [458, 98]. Flemming [480]. Flight [323]. Flipping [475]. Flocon [157]. Flood [290]. Florence [88, 24, 169]. Florinskii [398]. Flow [324]. Flower [575]. Flu [806]. Fluid [228]. Fly [56, 318]. Fog [331]. Fokko [637, 434]. Folklore [321]. Food [769, 504]. Foods [579]. Footprints [638]. Forbidden [545]. Forces [539]. Forecasting [334]. Foreign [744]. Forensic [748, 781, 667]. Forest [685]. Forgeries [630]. Forgery [630]. Forging [184]. Forgotten [187, 279, 231]. Forks [713]. Form [40, 617]. Formation [761]. Forms [416, 225, 530]. Formularies [27]. Förschler [339]. Forster [104]. Forterre [81]. Forth [445]. Forty [159]. Forum [372, 373]. Fossil [347]. Foster [600]. Foundation [95, 74, 561]. Foundations [59]. Founder [551]. Four [275]. Four-Volume [275]. Fourier [427]. Fourteenth [288]. France [96, 172, 7, 614, 377, 156, 743, 459, 592, 594, 781]. Frances [389, 624]. Francesco [387, 443]. Francis [422]. Franciscan [496]. Francisco [350]. Frank [692]. Frankel [476]. Frankenstein [46]. Frankreich [132]. Franz [583]. Frederick [215]. Freedom [65]. Freidenfelds [379]. French [238, 157, 305, 663]. Freud [301, 47]. Friar [496]. Friberg [23]. Friedrich [630]. Friend [820, 58]. Front [1, 70, 135, 192, 71, 136, 193, 261, 311, 364, 417, 262, 312, 365, 418, 462, 518, 567, 612, 463, 519, 568, 613, 660, 704, 756, 795, 661, 705, 757, 796]. Frontier [562, 246]. Frontiers [33]. Frühen [239, 339]. Fun [587]. Function [728]. Funding [512, 561]. Furchtlos [401]. Furniture [524]. Future [304, 190, 354, 246]. Futures [373].

G [304, 484, 55, 278, 288, 42, 443]. Gabriel [128, 234]. Gabriele [626]. Gabrielsson [514]. Gahtan [582]. Galan [180]. Galen [227, 216]. Galenic [724]. Galilei [235, 196]. Galileo [72, 196, 205, 95, 235, 336]. Galluzzi [289]. Galton [398]. Gamble [655]. Gänger. [494]. Gap [419]. Garden [305]. Gardener [369]. Gardens [680]. Garrett [206]. Garrod [96]. Gary [590]. Gas [509]. Gascoigne [84]. Gaston [157, 157]. Gathering [516]. Gaukroger [275]. Gave [747]. Gaze [673]. Gazetteers [623]. Gelehrtenleben [183]. Gender [301, 681, 453, 569, 224]. Gendered [218]. Gene [566, 223]. Genealogies [584]. General [147, 146]. Generating [474]. Genes [133]. Genetic [565, 696]. Genetics [126, 21, 651]. Genome [657]. Genre [96]. Gentzen [59]. Geoanthropology [528]. Geoffrey [664]. Geographers [407]. Geographical [584, 742]. Geographies [778, 734, 348, 450]. géographique [29]. Geology [10, 552]. Geometrica [5]. Geometry [174, 468, 266]. Geophysical [668]. Georg [104, 138]. George [181, 605]. Georgian [445]. Gerd [57]. Gerhard [59]. German [695, 75, 818, 324, 782]. germanique [29]. Germany [549, 252, 423, 107, 636, 500]. Germs [208]. Gerson [309]. Geschichte [63, 47]. Ghalandari [385]. Ghosh [356]. Ghosts [643]. Giannini [338]. Giant [265, 708]. Gibbs [215]. Gibson [734]. Giffort [412]. Gilded [599]. Gillen [296]. Gillin [595]. GIM [628]. Giordano [94]. Giorgio [182]. Giovanni [562, 634, 562]. Girolami [689]. Giulia [338]. Giunta [689]. Glacier [278]. Glass [792, 717, 520]. Glassware [520]. Glebe [108]. Glen [282]. Global [628, 699, 268, 254, 772, 807, 253, 217, 809, 337, 582, 240, 290, 341, 655, 351, 217]. Globalization [789, 257]. Godel [239]. Golden [49]. Golder [90]. Goldfish [280]. Goldgar [775]. Goldstein [86, 104]. González [393]. Good [609, 770]. Gooday [732, 556]. Goods [719]. Goose [320]. Gordin [303]. Gordon [284]. Gospel [208]. Gottfried [677]. Göttingen [402]. Grade [644]. Graeme [556, 732]. Graf [53]. Graf-Grossmann [53]. Grains [267]. Grand [4]. Grande [118]. Grandin [85, 281]. Gräper [317]. Grass [267]. Gravesande [237]. Gravity [72, 141, 456]. Great [31, 49]. Grèce [25]. Greco [562, 235]. Greece [163]. Greek [674, 164]. Green [789]. Greene [752]. Greening [131]. Greenwich [115, 115]. Greg [744]. Gregory [689, 736, 749]. Greifswald [359]. Grell [172, 215]. Greta [591]. Greylag [320]. Grmek [22]. Grønfeldt [538]. Groom [169]. Grootenboer [775]. Grosseteste [267]. Grossmann [53]. Grote [189]. Ground [514, 489]. Grove [445]. Growth [768]. Guba [594]. Guenzi [606]. Guerre [118, 25]. Guichelaar [58]. Guide [282]. Guidebook [179]. Guillaume [786]. Gulf [33]. Gullberg [438]. Gunpowder [174]. Günther [728]. Guston [46]. Gutmann [153].

H [304, 637, 295, 121, 64, 566, 62, 681, 34, 394, 23, 19]. Hahn [639]. Haigh [694]. Hallonsten [755]. Halpern [746]. Hamblin [655]. Hammar [19]. Hand [157, 95]. Handbook [232, 583, 83]. Hanford [408]. Hanif [385]. Hanna [292]. Hannah [545, 233]. Hanneke [775]. Hanns [152]. Hans [157, 609, 358]. Hans-Jörg [157]. Hans-Jürgen [358]. Happiness [380]. Hardback [188]. Hardin [87]. Harmless [176]. Harmonization [732]. Harmony [427]. Harold [457]. Harris [697, 125]. Harrison [547, 36]. Harry [607]. Hartmann [108]. Hartogsohn [513]. Harun [390]. Harvey [504]. Hasse [384]. Hathor [286]. Hay [370]. hay’a [385]. Hazardous [158]. Healers [231, 681, 586]. Healing [481, 234, 546]. Health [618, 111, 698, 502, 137, 807, 244, 231, 710, 548, 207, 224, 607, 785, 300, 632, 481, 632]. Healthcare [703]. Hearing [455, 380]. Heat [404]. Hebrew [384]. Hediger [570]. Heels [222]. Heggie [299]. Heidenblad [19]. Heimaquarium [397]. Heini [570]. Heinrich [108]. Heinz [43]. Heinz-Elmar [43]. Helen [18]. Helena [514, 281]. Helge [83]. Heligoland [370]. Hell [391, 81]. Hellenistic [163]. Helmholtz [302, 213]. Henry [55, 167, 501, 476, 709]. Hentschel [116]. Hepatitis [746]. Her [611]. Herbals [676]. Herbert [288]. Heredity [657, 21, 696]. Herman [301]. Hermann [302, 383]. Hermetic [286]. Hermeticum [285]. Hermeticum-Wirkungsgeschichte [285]. Herodotus [89]. Heroic [470, 94]. Herzog [479]. Heumann [63]. Hevelius [172]. Hevelius.Volume [172]. Hidalgo [638]. Higashi [30]. Higher [299]. Hiland [703]. Hilary [279]. Hilbert [402]. Hillel [739]. Himalaya [684]. Himalayan [283]. Himmel [359]. Hindenburg} [323]. Hindu [191]. Hippocratic [227]. Hispanic [393]. histoire [113, 330, 292]. Historia [633]. Historian [89]. Historical [732, 805, 818, 145, 449, 120, 692, 112, 582, 107, 804]. Histories [618, 210, 371, 291, 748, 180, 377, 809, 812, 240, 750, 111]. Historiography [577, 803, 800, 807, 530]. History [628, 304, 638, 161, 260, 372, 461, 495, 752, 213, 232, 689, 635, 159, 149, 395, 96, 21, 435, 656, 276, 16, 522, 550, 82, 702, 531, 817, 265, 503, 577, 789, 147, 393, 554, 520, 790, 121, 116, 298, 83, 526, 150, 171, 254, 127, 218, 431, 49, 772, 299, 645, 723, 215, 466, 574, 319, 430, 10, 146, 808, 603, 758, 659, 813, 248, 771, 162, 694, 144, 537, 500, 143, 816, 19, 711, 528, 18, 379, 482, 749, 485, 178, 163, 725, 223, 224, 341, 608, 209, 327, 517, 591, 422, 527, 309, 381]. History [8, 77, 49, 215, 54, 580, 213, 601, 738, 22, 583, 62, 220, 380, 693, 481, 328, 258, 170, 44]. Hobart [31]. Hochadel [180]. Hodge [736]. Hoffenberg [295].

Holbein [774]. Holdridge [763]. Holger [239, 63]. Hollings [45]. Holy [191]. Homberg [493]. Homburg [158]. Home [473]. Homemaker [710]. Hooke [443]. Hopkins [695]. Hopwood [480]. Horizons [436]. Hormones [328]. Horn [758]. Horst [95]. Horwitz [258]. Hospital [378]. Hospitals [27]. Hot [502, 251, 404]. Hour [333]. House [735]. Howard [374]. Howe [254]. HSS [373]. Huancavelica [770]. Hugh [699]. Hugo [126]. Hui [455]. Huib [434]. Human [123, 210, 21, 305, 657, 566, 355, 564, 227, 651, 746, 368, 357, 530, 696]. Humane [557]. Humanism [335, 559]. humanismo [730]. Humanities [485]. Humanity [678]. Humboldt [243, 654, 43]. Humor [139]. Humphry [68]. Hundred [641]. Hunger [383, 788, 641]. Hunt [188, 357, 236]. Hunter [780, 300]. Hunters [347, 749]. Huntington [283]. Hurricane [12]. Hussey [683]. Hutton [174]. Huxley [68]. Hybrid [577, 256]. Hybridity [420]. Hygiene [315]. Hygienic [202]. Hypotheses [314]. Hypothesis [468, 112, 315]. Hysteria [301, 328].

Ibero [635]. Ibero-Atlantic [635]. ibn [228]. Ice [296]. Icon [46]. Icy [499]. Ida [637]. Idea [87]. Ideal [698, 62]. Ideals [266]. Identification [313]. Identity [139, 569, 679]. Ido [513]. Idrāk [385]. Igniting [120]. Ignorance [17, 222]. ihrer [63]. Ihsanoglu [735]. II [197, 84, 744]. IIASA [423]. Ilaria [505]. Ilja [248]. Illness [481]. Illnesses [279]. illus [689, 187, 157]. Iltis [126]. im [397, 340, 47]. Imagery [73]. Images [319, 324]. Imaginary [232, 806]. Imagination [742, 553]. Imaginative [482]. Imagine [389]. Imagining [290]. Immanenz [285]. Immortality [190]. Impact [515]. Imperial [734, 298, 44, 238, 683]. Imperialism [244, 33]. Improvement [105]. Ina [63]. Inca [438]. Inclusive [149]. Independence [362]. index [689, 187, 157, 188, 449, 215, 692, 189, 694, 695]. India [294, 445, 791, 642, 814, 793, 606, 515, 511]. Indian [413, 649]. Indië [50]. Indies [589]. Indigeneity [565]. Indigenous [638, 41]. Indigo [663]. Indonesia [666]. Induction [40]. Industrial [114, 693, 194, 332, 639]. Industrialism [403]. Industrialization [654]. Industry [302, 60]. Inequality [808]. Inequities [808]. Infectious [801]. Influence [72]. Influences [260, 461, 659]. Information [628, 410, 137, 98]. Initiated [448]. Initiatives [131]. Inner [388]. Innovation [137, 310, 515, 491, 60, 403, 542, 227]. Inquiry [148]. Inquisition [773, 106]. Insane [119]. Inscriptions [739, 552]. Insect [688]. Inspirations [371]. Installation [615]. Institute [139, 367, 316]. Institutionalization [338]. Institutions [281]. Instructors [431]. Instrumental [413]. Instruments [156, 168]. Integration [378]. Intellectual [220, 163]. Intelligence [597]. Intelligent [628, 482, 611]. Interdisciplinary [149]. Interlocking [74]. Intermedial [318]. International [255, 745, 505, 668]. Internationalism [505]. Interoceanic [700]. Interpretation [107]. Interpretations [708]. Interpreters [436]. Interrogations [511]. Intersections [157]. Intersex [554]. Interviewing [8]. Interwar [195, 759, 315, 505]. Intimate [413].

Intoxication [337]. Introduction [469, 800, 764, 424, 198, 9, 574, 634, 143, 799, 316, 712, 527, 620, 157, 307, 408]. Intuitions [157]. Inventing [488, 665]. Invention [20, 109, 119]. Inventors [686]. Investigating [48]. Investigations [593]. Invisible [748, 804]. Iowa [693]. Iran [244]. Iranian [367]. Irena [378]. iride [267]. Irmscher [733]. Irrtümer [630]. Isaac [304, 141]. Isabel [414]. ISBN [689, 352, 250, 449, 215, 692, 694, 695]. Isis [260, 259, 461, 460, 658, 659]. IsisCB [799]. Islam [776, 437]. Islamicate [573, 811]. Islands [187]. Isotopic [187]. Israel [182, 579]. Issue [799]. Istanbul [390, 575]. Italian [289, 817]. Italy [26, 425, 769, 817, 545, 231, 440]. Itineraries [360]. ix [188]. Izrailevich [212]. Izumi [224].

J [304, 689, 159, 96, 434, 114, 134, 333, 736, 92, 414, 595, 288, 733, 281, 34, 227, 166, 749, 426, 404, 450, 178, 739, 597, 65]. J. [129]. Jabbār [385]. Jackson [783]. Jacob [100, 558, 655]. Jacobs [790]. Jacqueline [508, 121, 226]. Jacques [34]. Jahre [132]. Jahrestagung [359]. Jahrhundert [340]. Jaime [291]. Jaipreet [380]. James [551, 733, 68, 394, 113, 105, 234, 187, 458, 644, 342, 98, 556, 341]. James’s [179]. Jan [637, 434, 484, 50, 298, 59, 701, 58, 300]. Jane [602, 297, 378]. Janello [32]. Janet [794, 750]. Janusch [104]. Japan [510, 679, 355, 652]. Jar [619]. Jaramillo [687]. Jared [61]. Jason [46]. Jaume [249]. Jayram [604]. Jean [447]. Jean-Antoine [447]. Jeannie [688]. Jeff [87, 631]. Jeffrey [349]. Jenkins [555]. Jenks [751]. Jennifer [177, 416, 488, 344, 687, 774, 48]. Jenny [651]. Jeremy [752, 550, 240]. Jerónimo [730]. Jessica [245]. Jesuit [391, 767]. Jesuits [571, 350]. Jewett [382]. Jewish [565]. Jews [523]. Jimena [725]. Jinee [511]. Jo [86, 745]. Joachim [359, 509]. Joan [483, 549]. Joanne [458]. Jobin [284]. Joel [293]. Joelle [603]. Joffe [28]. Johan [85, 19]. Johannes [172]. John [783, 383, 542, 733, 353, 34, 98, 236, 467, 84, 122, 566, 127, 564, 440, 600, 251, 236]. Johns [695]. Johnson [449, 51]. Johung [416]. Jon [215]. Jonathan [289, 504, 115]. Jones [629, 162, 754]. Jöran [23]. Jörg [157, 250]. Joris [343]. José [229]. Josef [630]. Josefina [165]. Joseph [723, 234, 368]. Joshua [780, 254, 184]. Jost [401]. Journal [15]. Journey [187]. Jr [594]. Juan [585, 393]. Judgment [150]. Judith [516]. Juhani [168]. Jui [607]. Jülich [85]. Julie [109, 230]. Julie-Marie [109]. Junction [362]. Jungnickel [686, 500]. Jürgen [329, 104, 358]. Juste [167]. Justice [363, 408]. Justin [776].

Kaalund [499]. Kaasch [359]. Kadia [355]. Kalle [546]. Kalmbach [459]. Kananoja [546]. Kanjirakkat [284]. Kanta [482]. Kaolin [718]. Karen [408, 788, 433]. Karena [459]. Kari [19]. Karin [581]. Karl [85, 281, 602]. Kärnfelt [85]. Kassandra [333]. Kassell [480]. Kassenova [747]. Kat [686]. Kate [157]. Katelynn [287]. Katharina [484, 335]. Katherine [536]. Kathleen [378]. Kathryn [163]. Katrina [12]. Katrine [499]. Kazakhstan [747]. Keiner [700]. Keith [653, 227]. Kemp [49]. Ken [452]. Kennefick [117]. Kenny [514]. Keogh [737]. Kerr [453]. Kessler [288]. Key [491]. Kharaqī [385]. Ki [224]. Kieval [739]. Kind [526]. Kinds [46]. King [55, 733, 178, 786]. Kingdom [766]. Kingsberg [355]. Kira [586]. Kirchberger [400]. Kitchen [710]. Klaus [116, 358]. Klein [495]. Kleinbub [388]. Kleist [619]. klinischen [47]. Klose [690]. Knight [278]. Know [512, 441]. Knowing [268, 17, 222]. Knowledge [741, 102, 516, 429, 434, 484, 329, 35, 127, 595, 448, 180, 545, 56, 546, 15, 17, 19, 706, 615, 450, 370, 617, 273, 597, 33, 234, 715, 85, 772, 217, 218, 492, 101, 19, 240, 517]. Knowledges [742]. Knowles [128]. Koeppe [588]. Koerber [328]. Koetsier [628]. Kohler [123, 624]. Kolchinskii [212]. kollaborative [340]. Koloski [26]. Koloski-Ostrow [26]. Kontler [391]. Köppen [183]. Korea [464]. Koshatka [687]. Kouremenos [225]. Kragh [83]. Krebs [78]. Krementsov [398]. Krige [127]. Krill [521]. Kristin [654, 683]. Kruif [300]. Krupar [306]. Kücük. [390]. Kühn [484, 8]. Kunst [239]. Kunst- [239]. Kurbegović [692].

L [122, 333, 252, 751, 288, 87, 344, 246, 378]. Lab [147]. Labor [242]. laboratorium [50]. Laboratory [520, 65, 413]. Labortier [397]. Lachenal [786]. Lachlan [684]. LaFleur [591]. Lai [516]. Lakoff [697]. Land [281, 296]. Landscapes [446, 143]. Langston [643]. Language [800, 818, 292]. Lanka [814]. Lannoo [281]. Lapse [318]. Lara [379]. Larrie [332]. Larsson [19]. Lassman [60]. László [391]. Late [230, 348, 753, 215, 231, 164, 623]. Later [726]. Latin [384, 360, 378, 431, 816, 229]. Latitude [572]. Latour [721]. Laue [401]. Laue-Furchtlos [401]. Laura [361]. Lauren [480]. Laurence [640]. Laurie [206]. Lavender [81, 723]. Law [285, 759]. Lawrence [334, 493]. Lawrence-Mathers [334]. Lean [403]. Leap [708]. Learning [200, 147, 42]. Leary [179]. Lebiez [614]. Lecture [599, 8]. Lectures [307]. Lecturing [42]. Lederer [159]. Lee [481, 652, 124]. Legacies [305, 692]. Legacy [55, 452]. Legend [105]. Legitimacy [412]. Lehr [239]. Lehr- [239]. Leila [539]. Leisure [390]. L’élite [118]. Lemmerich [401]. Leonardo [517]. Leong [218]. Leprosy [745, 241]. Lernorte [239]. Leston [772]. Let [646]. Letter [672, 274, 375, 154, 376, 3, 14]. Letters [67, 352, 68]. Leung [224]. Lexical [170]. Ley [61]. Leyden [619]. Liba [162, 674]. Liberty [368]. Life [174, 549, 55, 525, 649, 190, 250, 753, 183, 763, 217, 415, 457, 125, 556, 242, 236, 105, 640, 753, 189, 698, 306, 702, 508, 416, 86, 226, 439]. Life-Making [306]. Life/Wladimir [183]. Light [161, 550, 646, 525, 116, 807, 543]. Lili [516]. Limits [752, 708, 773, 809, 604]. Limoge [288]. Lindee [331]. Lindemann [47]. Linearity [809]. Linguistics [697]. Linnaean [7]. Lion [636]. Lipton [129]. Lisa [61, 675]. Listening [668]. Lister [444]. Literacy [31]. Literature [805, 230, 678, 693]. Littoral [268]. Littorals [271]. Live [211, 200]. Lived [344]. Lives [348, 413]. Livestock [648]. Living [161]. Livingston [204]. Local [615, 403, 623]. Locations [434]. Locomotion [593]. Lodge [556]. LoGart [621]. Logic [59]. Logics [86]. Lokaneeta [511]. London [187, 731, 188, 215, 683, 189, 694]. Long [102, 42, 593, 409, 685, 360, 548, 256]. Longair [83]. Longitude [572]. Look [201]. Looking [288]. Loop [619]. Lorek [360]. Lorraine [357]. Lost [49]. Love [429, 427, 426, 53, 611]. Lovelace [45]. Loveland [631]. Low [510]. LSD [66]. Luca [630]. Lucia [228]. Lucifers [550]. Lüders [499]. Ludwig [317]. Luft [487]. Luigi [182, 185]. Lukas [347]. Luke [737]. Luminous [161]. lunationibus [91]. Lydia [290]. Lynn [678]. LZ [323].

M [286, 501, 21, 547, 736, 400, 603, 34, 488, 284, 18, 493, 695, 234]. M. [302, 331]. Mach [762]. Machine [628]. Machines [189, 511, 587, 628, 289, 482]. Mackenzie [727]. Mackowski [753]. MacLeod [55]. Mad [245, 607]. Made [351]. Madhouse [21, 682]. Madness [793, 603, 119, 17]. Magic [709, 101, 729, 94, 541, 166]. Magical [758]. Magnifying [267]. Mahaffey [187]. Mahony [742]. Maia [582]. Maienschein [602]. Mainz [689]. Maize [263]. Major [34]. Majorana [352]. Making [85, 102, 345, 347, 483, 134, 45, 335, 271, 80, 171, 726, 683, 180, 586, 455, 471, 315, 651, 754, 317, 760, 403, 370, 804, 714, 451, 8, 588, 784, 254, 644, 356, 306]. Malaquias [414]. Malarial [294]. Malcolm [83]. Malta [350]. Malvern [91]. Man [708, 80, 250]. Management [427, 718]. Managing [766, 768]. Mandelbaum [289]. Maner [107]. Manifold [410]. Manipulation [227]. Mannen [50]. Mannerism [95]. Manning [217]. Mantle [584]. Manuel [94, 650]. Mao [563]. Map [620]. Mapped [589]. Maps [638, 269, 538, 433, 623]. Mara [455]. Marcel [53]. Marci [789]. Marcia [395]. Marco [361, 63]. Marcus [545]. Mareike [63]. Marga [611]. Margaret [247, 97, 236, 100]. Margery [277]. Margócsy [28]. Maria [582, 777, 106, 777]. Marie [444, 109, 280]. Marilyn [297, 247]. Marina [442]. Marine [115, 115]. Marisa [775]. Mariss [339]. Maritime [115]. Mark [28, 310, 729, 561]. Market [310]. Marketing [653]. Marking [293]. Marra [251]. Marriage [428]. Marroquín [291]. Mars [184]. Martin [45, 243, 677, 742, 444]. Martina [16]. Martyr [207]. Marvels [588]. Mary [102, 46, 503, 125]. Mason [111]. Mass [184, 694]. Mass./London [694]. Masthead [2]. Mat [217]. Match [424]. Matchmaking [425]. Matemáticas [730]. Material [155, 520, 344, 719]. Materialism [86]. Materializing [199]. Materials [717, 171, 341]. Maternal [315]. Math [645]. Mathematical [138, 266, 23, 490, 156]. Mathematician [174, 214]. Mathematics [232, 182, 226, 645, 468, 788, 5, 404, 225, 490, 59, 53]. mathématiques [118, 30]. Mathers [334]. Mathias [189]. Matlin [602]. Matteo [634]. Matter [71, 136, 193, 262, 312, 365, 418, 463, 519, 568, 613, 661, 705, 757, 796, 540, 189, 37]. Matters [122]. Matthew [784, 62, 780, 457, 66, 234]. Mattsson [514]. Mauelshagen [583]. Mauersberger [358]. Maura [753]. Mauricio [635]. Mavhunga [56]. Max [402, 401]. Maximilian [391]. May [353]. Maya [581]. Mayer [593].

Mayhew [450]. McCormmach [500]. McKittrick [536]. McNeill [539]. McOuat [284]. Meaning [12]. Meanings [459]. Means [819]. Meant [7]. Measles [523]. Measurement [428, 605]. Meat [769]. Mechanical [443]. Mechanism [442, 170]. médecine [25]. Media [184, 381]. Mediated [605]. Medical [487, 367, 168, 74, 412, 586, 546, 592, 760, 370, 343]. Medici [169]. Medicinal [168]. Medicine [294, 111, 210, 230, 265, 245, 498, 790, 177, 310, 215, 683, 542, 545, 103, 680, 557, 681, 233, 746, 497, 781, 569, 679, 27, 188, 101, 106, 279, 387]. Medicines [516, 215, 548, 640]. Medieval [334, 230, 165, 437, 215, 586, 541, 543, 166, 167, 487]. Mediterranean [238, 810, 333, 724, 811]. Mediterraneanizing [420]. Meeting [77, 771, 580]. Meets [157, 10, 562]. Meli [170]. Melissa [177]. Membranes [189]. Memory [94, 264]. Menace [207]. Mental [116, 703, 315, 607]. Mentality [496]. Mentations [421]. Menthol [653]. Menzies [741]. Merchant [491, 485]. Merchants [548]. Mercury [770]. Merian [777]. Meridian [615]. Mertens [94]. Mesoamerican [265]. Mesopotamia [630]. metales [711]. Metamathematics [88]. Metaphor [584]. Meteorologie [183]. Meteorology [334]. Method [501]. Methodological [196]. Methodology [667]. Metropolis [245]. Mexican [682, 263]. Mexico [638]. Michael [111, 281, 630, 129, 232, 31, 345, 359, 109, 160, 303]. Michel [630, 161, 723, 330, 663]. Michelangelo [388]. Michelle [363]. Michotte [319]. Microbe [300]. Microbes [708, 804, 81]. Microbial [652]. microscoop [50]. Microscope [657]. Mid [469]. Mid-Twentieth [469]. Middle [726, 287, 603, 229, 696, 351]. Midwifery [707]. Migration [688]. Migratory [643]. Milam [357]. Milanich [506]. Mildred [748]. Military [108, 125, 512]. Miller [333, 105]. Millions [607]. Mills [455]. Minard [176]. Mind [598, 310, 419, 432, 341]. Mineral [768]. Mines [5]. Mining [768, 526, 770, 711]. Miriam [355]. Mirko [22]. Mirror [503, 98, 675]. Miscarriage [379]. Mission [512]. Missionary [448]. Missions [767]. Mist [456]. Mistress [368]. Mitch [95]. Mitra [649]. mitraille [118]. Mitton [508]. Mixtures [227]. Mobile [56]. Mobility [683]. Mobilizing [324]. Moby [178]. Moby-Dick [178]. Mode [148]. Model [149]. Models [116]. Modern [85, 425, 765, 111, 635, 769, 345, 96, 291, 722, 350, 775, 520, 489, 764, 768, 649, 335, 547, 761, 83, 173, 441, 35, 109, 244, 215, 492, 729, 545, 455, 440, 387, 5, 101, 98, 269, 270, 694, 711, 95, 630, 338, 497, 382, 589, 60, 106, 735, 290, 781, 787, 575, 331, 224, 679, 148, 719, 490, 234, 406, 119, 620, 676, 572, 507]. Moderna [777]. Modernism [266]. Modernity [202, 275, 177, 249, 603, 739, 640, 164]. Modification [366]. Molecular [188, 605, 189, 562]. Molecules [404]. Molella [128]. Money [716]. Monica [630]. Monika [358]. Monnais [640]. Monster [46]. Monstrosity [760]. Montalbani [38]. Montreal [264]. Montserrat [777]. Moon [708]. Moore [151]. Moral [197, 288]. Morality [759]. Moran [439]. Morange [723]. Morant [664]. Morbid [592]. Mordechai [338]. Morgan [749]. Morisy [129]. Moritz [782]. Morley [731]. Morocco [776]. Morphinisten [47]. Morphological [617]. Morris [414, 510]. Morse [247]. Moseley [55]. Most [171, 452, 75]. Mother [611]. Motion [85, 72, 175, 605, 540]. Mount [69]. Mountain [516]. Move [142, 717]. Moved [737]. Movement [220, 409]. Moves [127]. Moving [570, 319, 324, 124]. Mu’īniyya [486]. Mueller [46]. Muhammad [486]. Mühlenbrock [630]. Muise [723]. Mujeres [777]. Muka [792]. Mukherji [54]. Mukhopadhyay [54]. MUL.APIN [383]. Müller [57]. Mulsow [677]. Multidisciplinary [630]. Multilingual [298]. Muñoz [730]. Muntaha [385]. Murder [596, 187, 614, 739]. Murphy [233]. Museum [432, 115]. Museums [503, 361, 41, 49]. Musical [595]. Musicology [196]. Muslim [735]. Mussell [556]. Myriad [741]. Myth [669, 379, 611].

N [187, 28, 335, 449, 227, 23, 688]. N-Rays [187]. Nadine [306]. Nahin [404]. Nakayama [224]. Nall [184]. Nancy [643]. Nanna [499]. Nanoscale [515]. Naomi [512]. Nara [506]. Narration [633]. Narratives [499, 284, 482]. Narrator [89]. Nasīr [486]. NASA [753, 794, 708, 409]. Nasifoglu [490]. Nasim [738]. Natacha [37]. Natality [547]. Nathan [671]. Nation [516]. National [405, 813, 115, 561, 615]. Nationale [156]. Nationalism [191]. Nationhood [734]. Nations [253]. Nativitas [167]. NATO [131]. Natur [340]. Natural [776, 255, 265, 503, 520, 399, 736, 49, 709, 443, 178, 105, 591, 33, 96, 758]. Naturalienkammern [239]. Naturalis [633]. Naturalism [390]. Naturalist [104]. Naturalization [552]. Nature [539, 255, 155, 291, 350, 773, 717, 731, 20, 278, 464, 466, 727, 377, 600, 99, 292, 227, 446, 284, 320, 44, 357, 160, 552, 243, 503]. naturelle [292]. Naturgeschichte [339]. Naval [332]. Navarro [730, 249]. Navigating [97]. Nayar [92]. Nazi [695]. Neale [326]. Near [333]. Necropolis [346]. Nederland [50]. Need [535]. Neeraja [560]. Negotiating [767]. Neil [109, 773]. Nelson [774]. Neoliberalism [514]. Nets [133]. Network [571, 74]. Networks [713, 35, 400, 446, 448]. Neural [133]. Neuroscience [310, 264]. Neuzeit [239, 339]. Neville [676]. Newell [246]. Newhauser [288]. Newly [206]. Newman [99]. News [184, 410]. Newton [141, 314, 99]. Nice [247]. Nicholas [741, 44]. Nick [480]. Nicola [783]. Nicolaidis [164]. Nicolas [79]. Nieto [635, 180]. Nieto-Galan [180]. Nieuwland [248]. Night [438]. Nikfahm [486]. Nikolai [398]. Nikolaus [384, 386]. Nile [344]. Nilsson [19]. Nineteenth [596, 784, 177, 243, 445, 499, 595, 448, 42, 593, 685, 15, 594, 450, 650, 343, 597, 713, 614, 760, 553]. Nineteenth-Century [596, 784, 177, 243, 595, 448, 594, 650, 343, 597, 713, 614, 760, 553]. Nisbet [701]. Niter [464]. No [117]. Noah [565]. Noakes [406]. Nobelpreisträgers [401]. Nobody [740]. Noel [625]. Nollet [447]. Non [468]. Non-Euclidean [468]. Nonhumans [294]. Nonreusable [322]. Noortje [790]. Nordberg [19]. Nordic [75]. Nordlund [281]. Normal [140]. normaliens [118]. Norri [168]. North [606, 499, 52, 160]. Norton [302]. Norway [101, 411].

Nostalgia [221]. Note [534, 662]. Notes [59, 157, 692, 694, 695]. Nothaft [91]. Novaes [407]. Nuclear [187, 656, 252, 362, 510, 655]. nucleares [610]. Numbers [232, 405, 87]. Numeracy [31]. Nummedal [636]. Nuremberg [233]. Nurses [396].

Object [249, 780]. Objects [666, 483]. Objekte [63]. Obligation [197]. Observation [366]. Observatory [615]. Observers [39]. Observing [667]. Occult [406]. Ocean [512]. Oceanic [521]. Oceans [792, 18]. Odom [409]. Oertzen [218]. Oestmann [728]. Ogilvie [247]. Ohm [138]. Ohrvik [101]. Okapi [468]. Olarte [635]. Ole [215]. Olga [26]. Oliver [556, 180]. Olof [755]. Omar [738]. One [111, 708, 669]. One-Sex [669]. Ontological [801]. Onur [132]. op [50]. Open [431]. Open-Access [431]. Opera [553]. Operations [465]. Oppenheimer [479]. Opponent [820, 58]. Opportunity [782]. Optical [267]. Optics [288]. Opus [386]. Oram [66]. Order [233]. Ordering [741]. Ordnen [239]. Ordnen-Vernetzen-Vermitteln [239]. O’Reagan [695]. Oreskes [512]. Organ [596]. Organic [549]. Organism [647]. Organisms [161]. Organization [607]. Organizational [114]. Organizations [794]. Origin [348, 449]. Origins [722, 699, 649, 271, 466, 253, 726, 342, 619, 264, 337]. Orlemanski [230]. Ornamental [176]. Ornithological [159]. Ornithologist [247]. Oropallo [626]. Osmond [68]. Osseo [362]. Osseo-Asare [362]. Östling [19]. Ostrow [26]. O’Sullivan [348]. Other [245, 536, 634, 561]. Our [295, 93, 433]. Ovidio [38]. Owen [217, 222]. Own [295]. Ownership [28]. Oxford [83, 449].

P [559, 254, 128, 409, 374, 227]. Pablos [610]. Pacific [120]. Padrón. [589]. Page [234]. Pain [701]. Painting [780]. Pakistan [814]. Palate [465]. Paleontologists [702]. Palestine [579]. Palgrave [583]. PAll [176]. Palmieri [185]. Pandemic [806, 807, 563, 208]. Pandemics [805, 815, 818, 810, 808, 813, 816, 642, 799, 802]. Pankaj [515, 413]. Paola [442]. Paolo [387, 185, 289]. Paolucci [185]. Papago [137]. Paper [195, 218, 715, 188, 352, 692, 694, 695]. Paperback [157, 187]. Paracelsus [439]. Parallels [560]. Paris [392]. Park [611]. Parshall [788]. Parts [168, 24, 228]. Parvati [793]. Past [389, 354, 254]. Pasteur [367]. Pasts [436, 107]. Pat [129]. Patent [732]. Paternity [506]. Pathological [22]. Pathologies [520]. Patience [281]. Patient [465]. Patienthood [437]. Patients [209, 48]. Patrick [217, 68, 81, 217, 110]. Patronage [74, 194]. Patterns [772, 217]. Patterson [590]. Paul [96, 445, 93, 42, 693, 68, 694, 300, 670, 305, 20, 680, 41, 404]. Paul-André [305]. Paula [35, 724]. Pavuk [779]. Peace [751]. Peaceful [655]. Pearce [647]. Pearls [767]. Pearson [346]. Pecar [239]. Pedagogical [143]. Peder [411]. Pedrito [687]. Pegasus [187]. Pemberton [109]. Pendulum [272]. Penser [30]. People [701, 356]. Pepys [236]. Perception [345, 268, 319, 667]. Pereira [78]. Perfect [379]. Perfection [727].

Performing [645]. Period [505]. Periodic [689]. Periodicals [784]. Peripheral [755]. Periphery [411]. Perkowitz [46]. Perrin [253]. Person [678, 173]. Personae [195]. Perspective [635, 815, 268, 10, 732, 112, 582, 163]. Perspectives [689, 414]. Pete [176]. Peter [414, 288, 215, 402, 300, 304, 295, 399, 278, 152, 112]. Peterson [581]. Petromoderne [690]. Petros [216, 542]. Pettigrew [34]. Pfister [583]. Phantsy [369]. Pharmaceuticals [640]. Pharmacopoeias [234]. Pharmacy [27, 441, 724]. Phenomenology [185]. Phil [778]. Philadelphia [681]. Philip [227, 105, 69, 587, 537, 490]. Philipp [91]. Phillip [491]. Phillips [753]. Philosopher [167]. Philosophical [369, 88]. Philosophie [132]. Philosophizing [314]. Philosophy [743, 157, 226, 442, 369, 105, 237, 647, 443]. Photographic [75, 140, 363]. Photography [780, 363]. Photons [116]. Phrenology [428, 596, 341]. Physical [302, 748, 54, 664]. Physicians [586, 233]. Physicist [500]. Physics [232, 250, 562, 402, 384, 500, 406]. Physik [401]. Physiognomy [675]. Physiology [299, 426]. Pia [106]. Picatrix [541]. Pieces [24]. Piechocki [335]. Pierce [497]. Pietro [562, 235]. Piety [437]. Pig [421]. Pimentel [585]. Pinch [720]. Pingree [541]. Pinto [793]. Pioneer [556, 411]. Piot [447]. Pippin [391]. Pit [174]. Place [93, 471]. Placenta [707]. Placing [742, 622, 209]. Plague [769, 205, 206, 811]. Plaintiffs [408]. Planet [255]. Planetary [618, 276, 529, 485]. Planning [255]. Plant [688]. Plants [452, 737]. Plaster [248]. Plato [59, 225]. Pliny [633]. Plotnick [702]. Plurality [572]. Plutarch [72]. Podolsky [64]. Poems [634]. Poetics [96]. Poetry [783]. Poison [509, 544, 215]. Poisonous [257]. Poisons [215]. Polar [277]. Polaris [52]. Pole [160, 52]. Policing [511]. Policy [305]. políticas [610]. Politics [698, 769, 743, 346, 15, 701, 222, 608, 785]. Pollution [257]. Polymath [79, 637]. Pop [685]. Populace [108]. Popular [762, 146]. Population [785]. Porter [21, 40]. Porträt [358]. Portugal [350]. Poskett [341]. Post [470, 210, 517]. Post-Heroic [470]. Post-Truth [517]. Postcolonial [666]. Postimperial [420]. Postrevolutionary [592]. Postwar [743, 253]. Power [195, 252, 362, 510, 411, 700, 203]. Powering [351]. Pox [393]. pp [689, 187, 157, 188, 352, 250, 449, 215, 692, 189, 694, 695]. Practical [390, 446]. Practice [523, 217, 318, 648, 353]. Practices [218, 558, 256, 48]. Practicing [606]. Pragmatic [526]. Pragmatism [647]. Praise [145]. Pratik [552]. Praxis [210, 47]. Preaching [5]. Precision [713, 715]. Prediction [571]. Preface [449]. Pregnancy [379]. Prenatal [673]. Preparations [168]. Preparing [691]. Presence [423, 448]. Present [480, 252, 514]. Preservation [579]. Press [157, 188, 449, 692, 189, 694, 695]. Priestley [368]. Primary [141]. Primer [307]. primeras [610]. Principe [493]. Pritchard [601]. Pritikin [408]. Priya [405]. Probability [232]. Problem [175, 429, 40]. Production [628, 161, 490, 56]. Professional [468]. Professions [557]. Program [708]. Progress [304, 128]. Project [430, 420]. Prolégomènes [172]. Prometheans [598]. Proof [781]. Proofs [387]. Proper [197]. Properties [228]. Prophet [61]. Prophets [304, 68].

Proposed [142]. Prospecting [765]. Protect [766]. Protestant [448, 342, 29]. Prototyping [464]. Prussia [495]. Psychedelic [412, 513, 66, 68]. Psychiatry [743, 222, 186, 692]. Psychics [406]. Psycho [140]. Psycho-Clinic [140]. Psychoanalysis [315]. psychologie [113]. Psychology [185, 315, 114]. Psychotherapy [66]. Ptolemaic [385]. Ptolemy [226]. PTSD [258]. Public [295, 75, 244, 710, 181, 770, 207, 561]. Pugwash [309]. Punch [522]. Punch-Drunk [522]. Purabi [54]. Pure [504]. Pursuit [64, 231]. Pursuits [451]. Pushing [653]. Putin [309].

Qing [623, 667]. Qualities [354]. Quanta [116]. Quantification [405]. Quantitative [57]. Quantum [232, 402, 456]. Quarantine [708]. querelle [172]. Quest [698, 412, 99, 506, 517]. Quinlan [592].

R [175, 789, 699, 476, 438, 99, 336]. Rabies [245]. Race [421, 395, 664, 263, 699, 75, 296, 727, 103, 632, 420, 760, 341, 126]. Race-Making [760]. Rached [603]. Rachel [111, 602, 257]. Racial [653, 808, 363, 453]. Racism [126, 664, 688]. Radar [125]. Radder [609]. Radiation [349]. Radical [743]. Radick [736]. Radio [125, 668]. Radiobiology [421]. Ragab [437]. Raggetti [228]. Raghav [605]. Raia [598]. Rain [257]. Rainbow [267]. Rainer [239]. Ralph [291]. Ramos [682]. Rampling [488]. Randalls [742]. Randy [697]. Rankin [544]. Raphaële [96]. Rasmus [538]. Rasmussen [120]. Rational [64, 331]. Rawi [23]. Ray [42]. Rays [187]. Re [523, 5]. Reactor [510]. Reader [711]. Reading [490]. Real [443]. Realities [22]. Reality [605, 321, 711]. Reason [181]. Reasoning [790]. Reassessed [79]. Rebecca [480]. Rebekah [481]. Rebel [174]. Recalcitrance [249]. Recami [352]. Recasting [809]. Reception [216, 384, 572]. Recipes [487]. Recollections [52]. Reconsidering [301, 712]. Reconstructing [609, 148]. Recovering [202, 556]. Redrawing [620]. Rees [504]. Reflections [273]. Reform [64]. Reformation [636, 233]. Refrigerator [611]. Regimes [787]. Regius [237]. Regner [78]. Regulating [759]. Regulations [124]. Reid [491]. Reill [152]. Reis [554]. Reiß. [397]. Related [805]. Relations [744, 651]. Relationships [367]. Relativity [117]. Relief [641]. Religion [635, 31, 791, 729, 342, 87, 166, 327]. Religions [517]. Remains [265]. Remaking [190, 319, 189, 684]. Remarkable [444]. Remedies [441, 724]. Remedy [494]. Renacimiento [730]. Renaissance [232, 412, 634, 377, 231, 289, 544, 92, 587, 727, 675]. Renn [329]. Renny [791]. Rens [772]. Reply [478, 533]. Representation [268]. Representing [731]. Reproduction [480]. Republic [345, 356, 368]. Rereading [208]. Resag [250]. Research [367, 520, 137, 197, 753, 319, 194, 560, 316, 381]. Reser [539]. Reservation [137]. Reservations [472]. Resiliency [12]. Resistance [64]. Resisting [126]. Resource [431]. Resources [255, 764, 714, 712, 719]. Respectably [779]. Response [814]. Responses [92, 807]. Restricted [656]. Restructuring [563]. Rethinking [367, 637, 329, 527, 637]. Retracing [342]. Return [813, 600]. Returns [363]. Reusable [322]. Reusing [319, 316, 324]. Revealed [776]. Review [628, 647, 304, 792, 638, 221, 85, 776, 741, 102, 539, 161, 297, 495, 487, 158, 752, 213, 390, 175, 232, 81, 590, 509, 689, 294, 783, 516, 29, 410, 156, 111, 601, 286, 444, 38, 241, 698, 635, 159, 686, 559, 551, 596, 174, 549, 123, 501, 778, 31, 345, 743, 395, 238, 96, 734, 27, 732, 293, 21, 126, 168, 84, 435, 306, 187, 656, 602, 291, 585, 16, 730, 458, 392, 550, 637, 588, 598, 82, 702, 738, 230, 784, 305, 434, 484, 124, 22, 503, 122, 347, 302, 583, 789, 245, 536, 677]. Review [329, 540, 483, 114, 508, 748, 350, 773, 502, 416, 678, 295, 498, 699, 646, 186, 393, 134, 775, 55, 118, 554, 130, 86, 396, 165, 489, 731, 629, 793, 45, 504, 333, 67, 28, 790, 121, 480, 116, 20, 649, 335, 385, 486, 657, 50, 64, 547, 177, 252, 57, 298, 83, 397, 188, 216, 352, 182, 606, 310, 389, 399, 736, 131, 172, 566, 173, 285, 703, 441, 39, 89, 35, 179, 243, 92, 584, 445, 438, 296, 499, 358, 359, 171, 401, 254, 414, 127, 610, 307, 751, 218, 412, 109, 348, 595, 128]. Review [653, 454, 190, 278, 355, 184, 654, 605, 452, 226, 442, 69, 633, 25, 448, 40, 250, 93, 288, 257, 772, 280, 449, 120, 299, 383, 104, 496, 807, 301, 697, 279, 239, 386, 42, 244, 645, 753, 437, 723, 565, 692, 249, 362, 62, 791, 742, 108, 253, 634, 674, 220, 726, 26, 727, 447, 183, 593, 683, 180, 185, 542, 283, 377, 346, 349, 586, 733, 380, 745, 643, 788, 515, 189, 361, 492, 729, 340, 339, 564, 562, 545, 400, 740, 409, 56, 546, 231, 103, 541, 632, 780, 693, 413, 737, 459, 511, 747, 455]. Review [287, 641, 59, 510, 353, 507, 440, 680, 405, 609, 436, 481, 402, 387, 744, 247, 415, 133, 68, 513, 456, 408, 603, 592, 491, 644, 724, 685, 328, 432, 557, 681, 457, 117, 587, 63, 600, 248, 101, 277, 229, 99, 15, 342, 87, 631, 34, 94, 98, 176, 292, 560, 675, 110, 394, 594, 181, 162, 227, 694, 407, 47, 17, 32, 439, 488, 636, 112, 446, 691, 384, 258, 782, 191, 125, 23, 537, 500, 360, 308, 651, 556, 284, 19, 682, 233, 97, 337, 24, 354, 251, 582, 754, 95, 391, 336, 746, 132]. Review [555, 344, 411, 18, 388, 282, 543, 563, 363, 701, 630, 37, 30, 242, 338, 493, 41, 453, 289, 379, 240, 129, 786, 91, 164, 222, 642, 497, 51, 728, 228, 382, 219, 482, 695, 166, 443, 749, 43, 170, 589, 60, 106, 44, 735, 403, 290, 485, 113, 777, 357, 36, 781, 404, 687, 787, 107, 450, 246, 398, 794, 225, 494, 684, 54, 178, 639, 650, 548, 167, 690, 505, 558, 160, 58, 163, 750, 100, 330, 725, 223, 331, 696, 224, 700, 679, 538, 774, 341, 608, 544, 739, 599, 90, 604]. Review [607, 506, 378, 105, 652, 48, 115, 169, 517, 235, 343, 785, 591, 88, 597, 33, 640, 514, 779, 53, 356, 66, 52, 451, 65, 309, 512, 381, 755, 490, 688, 300, 234, 303, 256, 655, 561, 406, 119, 351, 676, 433, 553, 552, 237, 611]. Reviews [157, 49, 215, 217, 281, 46, 236, 61]. Revisiting [205]. Revival [412]. Revolution [84, 654, 251, 665, 608, 789, 93, 32, 639]. Revolutionary [7, 104]. Revolutions [217]. Reyes [407]. Reynolds [780]. Rheinberger [157]. Rhetoric [492]. Rhetorical [328]. Rhinoceros [758]. Rhodes [307]. Ricardo [589]. Rice [45]. Richard [307, 288, 733, 629, 178, 406]. Rickles [456]. Riemann [112]. Rienk [637, 489]. Rieppel [347]. Rift [31]. Rights [409, 368]. Rim [120]. Ring [120]. Risāla [486]. Rise [295, 188, 332, 644, 233, 219, 688]. Risk [158, 708, 121]. Ritual [739]. River [272]. Rob [557]. Robcis [743]. Robert [442, 443, 46, 123, 50, 307, 450, 267, 492]. Robert-Jan [50]. Roberts [286]. Robertson [745]. Robin [188]. Robino [442]. Robinson [310, 287, 46]. Robison [586]. Roch [157]. Rocío [484]. Rocks [125]. Rodger [458]. Rodgers [282]. Rodríguez [165]. Roel [16]. Roger [159, 736, 386, 220]. Rogue [578]. Rohan [294]. Rojas [436]. Roland [783]. Role [465, 614]. Rolling [178]. Roma [182]. Roman [773, 674, 726, 26, 436, 292]. Romantic [86]. Romanticism [293, 40, 242]. Romeiras [350]. Romero [610]. Ron [242]. Ronald [87, 396, 49]. Ronzon [361]. Rood [217]. Roof [684]. Roos [444, 280]. Roots [286]. Roquette [112]. Rosenkrantz [153]. Rosental [305]. Rossi [345]. Rothschild [257]. Roundtable [76]. Routledge [179, 215]. Roy [294, 702, 55]. Royal [85, 731, 493]. Rozwadowski [18]. Rubin [508]. Rudolph [122]. Rule [241]. Rules [314]. Russel [110, 394]. Russell [55, 288, 500]. Russia [190, 398]. Russian [734, 44]. Ryan [165]. Ryukyu [766].

S [689, 435, 602, 151, 396, 83, 736, 724, 88, 61, 309, 8]. Saab [483]. Sabine [308]. Sacco [443]. Sacha [726]. Sacred [761, 166]. Sailing [97]. Sajjad [486]. Sale [366]. Salguero [497]. Sally [177]. Saltpeter [464]. Salvadori [630]. Sam [583]. Samama [25]. Samantha [792]. Samuel [742, 236, 141, 785]. Sand [267]. Sandmo [19]. Sanitation [26]. Sankaran [560]. Santa [687]. Sapienza [182]. Sara [601]. Sarah [333, 482, 793, 354, 676]. Sarkar [646]. Sarukkai [284]. Satellite [699]. Savadi [486]. Savant [392]. Saved [503, 59]. Savelli [217]. Savoir [29]. Scaglia [505]. Scale [4, 450, 175]. Scenes [587, 691]. Schäfer [16]. Schatzberg [82]. Scheffler [188]. Schell [281]. Schirrmacher [402]. Schlagintweit [782]. Schleper [255]. Schmitt [24]. Scholar [183]. Scholarship [523, 799, 804]. School [97, 54]. Schotte [97]. Schulz [358]. Schulz-Fieguth [358]. Schummer [509]. Science [161, 77, 260, 372, 461, 213, 175, 762, 444, 698, 272, 551, 123, 467, 31, 345, 734, 126, 392, 588, 598, 531, 275, 722, 122, 302, 263, 536, 329, 540, 350, 295, 423, 498, 55, 520, 86, 75, 731, 196, 790, 145, 761, 526, 80, 131, 499, 171, 127, 431, 348, 128, 427, 250, 466, 574, 622, 593, 180, 377, 515, 361, 759, 729, 545, 146, 609, 133, 281, 659, 685, 457, 99, 87, 771, 270, 162, 446, 782, 191, 360, 143, 97, 391, 528, 760, 338, 453, 368, 576, 219, 695, 324, 60, 46, 781, 450, 575]. Science [794, 530, 331, 696, 700, 341, 739, 599, 652, 327, 422, 527, 33, 755, 256, 561, 648, 620, 621, 580, 8, 237, 390, 84, 187, 69, 40, 709, 791, 744, 342, 308, 284, 41, 382, 684, 512, 539, 635, 123, 291, 773, 295, 674, 162, 251, 754, 95, 543, 725, 544, 65]. Sciences [616, 792, 85, 776, 371, 469, 399, 753, 217, 240, 368, 735, 719, 561, 406, 680, 493, 54, 330]. Scientific [741, 559, 174, 501, 84, 784, 147, 352, 35, 74, 595, 93, 666, 217, 511, 15, 32, 125, 24, 282, 316, 382, 139, 665, 714, 517, 343, 779, 309, 553, 608, 219]. Scientist [89, 223, 52, 130, 45]. Scientist-Subject [223]. Scientists [150, 355, 46, 599, 129]. Scott [128, 49, 641, 46, 64]. Scott-Smith [641]. Scull [186]. Sea [602, 271, 733, 178, 792]. Sean [592, 432]. Search [467, 502, 399, 751, 420]. Seas [332]. Seb [543]. Sebastian [381]. Second [133, 500, 695]. Secondary [817, 141]. Secrecy [656]. Secret [187, 99]. Secrets [707, 296, 377]. Section [121]. Secular [100]. Sediments [272]. See [535, 483]. Seeds [577, 574, 578, 576, 575, 267]. Segal [374]. Segregation [378]. Sekhsaria [515, 413]. Sela [167]. Selcer [253]. Selection [736]. Selves [173, 266]. Seman [687]. Semi [346]. Semi-Colonial [346]. Sen [405]. Senegal [663]. Sense [483, 220, 287]. Seow [787]. Sepkoski [454]. Serber [307]. Series [157, 275, 373]. ses [172]. Set [513]. Seth [605, 103, 120]. Sets [229]. Setting [513, 589]. Seventeenth [776, 590, 444, 389, 445, 616, 523, 369]. Seventeenth- [616]. Seventeenth-Century [776, 590, 389, 523, 369]. Sewers [26]. Sex [649, 669, 453]. Sexton [68]. Sexual [698, 502]. Sexuality [649, 591]. Shack [668]. Shackleton [11]. Shadow [117, 725]. Shah [223]. Shaky [489]. Shallow [269, 270, 273]. Shaped [784, 503, 499, 642, 512]. Shapes [719]. Shaping [111, 275, 310, 266, 794]. Sharon [231]. Shawn [363]. Shea [336]. Sheila [92]. Shelley [46]. Shells [775]. Shepherd [256]. Sher [309]. Shiloh [306]. Shin [30]. Shindell [457]. Shinozuka [688]. Ship [491]. Shlomo [167]. Shock [701]. Shooting [317]. Short [800, 333, 34, 258]. Shorthand [59]. Shuffled [323]. Shuttleworth [177]. Siam [346]. Sibylla [777]. Sickness [203]. Sicknesses [168]. Sidney [46]. Siebert [16]. Siècle [739, 30]. siècle.} [447]. Sifilografía [393]. Sigmund [47]. Significance [438, 728]. Silke [165, 339]. Silver [432]. Simon [508, 138, 171]. Simone [255, 131]. Simple [194, 737]. Simpson [458]. Simulation [562]. since [469, 744, 107]. Singapore [352]. Singer [227]. Singular [494]. Sinister [166]. Sir [780, 139]. Sitter [58, 820]. Sivin [671]. Sixteenth [635]. Sixteenth-Century [635]. Sketch [449]. Skies [257]. Skill [97]. Skills [313]. Skulls [453]. Sky [438]. Slave [650]. Sleeping [264]. Slicing [313]. Slonosky [39]. Slotten [699]. Slugnuts [522]. Small [708]. Smallpox [523]. Smeall [786]. Smell [287]. Smith [96, 157, 605, 279, 220, 641, 394, 363]. Smithson [551]. Smithsonian [551]. Snelders [241]. Social [297, 195, 649, 298, 150, 427, 632, 808, 125, 240, 561]. Socialism [600]. Societies [343]. Society [77, 372, 734, 245, 731, 790, 515, 771, 563, 119, 8, 580, 85, 302, 609, 750]. Solar [93]. Solovey [561]. Solve [410]. Solveig [85]. Sombor [496]. Someday [206]. Somos [28]. Sonderausstellung [630]. Sonic [313]. Sonorous [761]. Sophie [446]. Soranzo [634]. Soraya [657]. Soul [675, 65]. Sound [313, 138, 714, 595]. Sounding [713]. Source [287]. Sources [817, 497]. sous [118]. South [565, 809, 33, 814, 516]. Southeast [813]. Southeastern [378]. Southern [755]. Sovereign [346]. Soviet [44, 309]. Space [699, 751, 753, 61, 298]. Spaces [761, 268]. Spaceship [253]. Spain [589, 724]. Spanish [806, 32]. Sparked [351]. Spatial [271]. Spatiality [450, 621]. Speaker [313]. Special [799]. Species [49, 643, 176, 555, 449, 503]. Spectacle [347]. Speculazioni [38]. Spheres [385, 774]. Spirit [559, 94]. Spisak [68]. Splendor [588]. Springer [689, 250]. Springs [502]. Sputnik [699, 668]. Sri [814]. Stabilizing [615]. Stable [271]. Stahnisch [692]. Stamhuis [637]. Standardization [520]. Standardized [644]. Stars [425, 246]. Stasi [313]. State [84, 569, 300, 119, 790]. Statecraft [356]. States [632, 760, 656, 124]. Stationers [676]. Statistical [715]. Statistics [108, 356, 648, 405]. Steadman [587]. Steam [105]. Stearns [776]. Steel [713, 167]. Steele [383, 408]. Steere [558]. Steere-Williams [558]. Stefanie [494]. Steigerwald [549]. Steiner [256]. Steininger [690]. Steno [79]. Step [708]. Stéphane [24]. Stephanie [34, 750]. Stephen [28, 409, 241, 559, 275, 482]. Steppe [747]. Sterckx [16]. Stern [726]. Steven [732, 167, 256, 698, 551, 438, 17]. Stevens [163, 378]. Stevin [637]. Stewart [227]. Stimpson [49]. Stob [42]. Stoecker [63]. Stomach [641]. Stones [666]. Stories [123, 536]. Story [653, 93, 281, 746, 543, 611]. Storytelling [145]. Strang [33]. Strange [109]. Strasbourg [728]. Strasser [134]. Strazzoni [237]. Strengths [807]. Strocchia [231]. Structure [697]. Struggle [65]. Students [431, 586]. Studies [814, 526, 692, 166, 490, 407, 36]. Study [27, 602, 657, 605, 430, 630, 91, 228, 301]. Sturge [157]. Subaltern [420]. Subcultures [592]. Subdivision [385]. Subject [223]. Subjectivity [320]. Subjects [294, 230]. Subramaniam [191]. Subrata [133]. Substitutes [718]. Success [465]. Suchtmitteln [47]. Suman [103]. Sumillera [484]. Sun [438, 589]. Sundar [284]. SUNY [157]. Supernatural [598]. Supplement [486]. Supplies [716]. Surgery [465]. Suriname [241]. Surman [484, 298]. Surprising [543]. Surroundings [435]. Surveillance [129]. Survey [818, 576]. Susan [392, 331, 492, 681]. Sustainability [20, 485]. Suvobrata [646]. Swahili [706]. Swan [775]. Swarm [381]. Sweden [755]. Swedish [85]. Sweet [86]. Sweetness [203]. Swerdlow [625]. Swing [369]. Switzerland [250]. Sydney [746]. Symbols [685]. Symons [333]. Symposium [689]. Symptomatic [230]. Syndemics [803]. Syndrome [11]. Syphilis [502]. Syria [448]. System [531, 93, 340, 530]. Systems [137, 807, 26]. Szpiech [165].

T [414, 231, 408, 394, 439]. Tūsi [486]. Table [689]. Tables [229, 215]. Taking [252, 204, 695]. Talbot [67]. Tale [560]. Tamarkin [565]. Tamborini [63]. Taming [594]. Tan [679]. Tank [792]. Tanya [348]. Taqāsīm [385]. Tara [636]. Tarrant [773]. Tasting [354]. Taub [674, 162]. Taxidermists [503]. Taxonomy [263]. Tea [354]. Teach [122]. teaching [149]. Teams [794]. Tech [703]. Technics [353]. technique [447]. Technocracy [252, 787]. Technological [92, 61]. Technologies [717, 403, 564]. Technology [752, 195, 121, 137, 526, 128, 349, 515, 361, 609, 360, 219, 695, 331, 256, 601, 82, 639, 127, 491, 655]. Technoscience [495]. tecnologías [610]. Telehealth [752]. temperamentis [227]. Temporalities [521]. Temporally [780]. Tendaguru [63]. Tendaguru-Expedition [63]. Tenorth [43]. Teratology [760]. Teresa [81, 723, 687]. Terrall [102].

Terraqueous [268]. Tertium [386]. Testimonies [352]. Testing [455, 644]. Teun [628]. Text [526]. Texts [487, 445, 23, 24]. Their [72, 686, 371, 537]. Theodore [21]. Theokritos [225]. Theoretical [57, 562, 500]. Theories [141, 555]. Theorizing [297, 804]. Theory [157, 605, 117, 227]. Theory-Mediated [605]. Therapeutics [64]. Therapy [349, 66, 703]. There [752, 646]. Theunissen [648]. Thibaud [113]. Thing [678]. Things [741, 176, 535]. Thinking [501, 454, 95, 482, 489, 42]. Third [201]. Thirteenth [288]. Thomas [791, 221, 694, 60, 36, 8]. Thompson [596, 452]. Those [426]. Thought [157, 633, 432, 649]. Threats [61]. Three [67, 530]. Thrifty [171]. Tides [272]. Tiere [339]. Timberlake [93]. Time [221, 293, 702, 333, 80, 318, 363, 799, 433]. Time-Lapse [318]. Times [16, 177, 321, 636]. Timothy [360, 547, 445]. Tinne [740]. Tkaczyk [455]. Tobacco [653]. Tobias [630]. Todd [93, 34]. Together [790, 42]. Togzhan [747]. Toilets [26]. Tom [509, 641]. Tools [712]. Torriani [32]. Torsten [359]. Tour [452, 377]. Traces [436]. Tracing [758, 378]. Trade [676, 337, 650]. Tradition [342, 227, 228, 88]. Traditional [741]. Trail [638]. Trais [346]. Traisnel [451]. Trajectories [560]. Trans [404, 81, 392, 108, 185, 405, 95, 289]. Trans-Atlantic [404]. Transatlantic [305, 733, 711, 650]. Transcribing [667]. Transformation [525, 703, 176, 514]. Transformations [217, 469]. Transformative [80]. Transformed [698, 279]. Transfusion [651]. Transgender [130]. Transimperialism [785]. Translated [157, 243, 403, 772, 723]. Translating [523, 291]. Translation [524, 634, 185, 217, 706, 91, 228, 484, 484]. Translational [310]. Transmutation [555]. Transmutations [493]. Transnational [127]. Transpacific [589, 688]. Transparency [717]. Transwar [355]. Transzendenz [285]. Travel [499]. Traveling [282]. Traversing [430]. Treating [502]. Treatise [486, 288, 541]. Treatment [701]. Treder [358]. Trees [766]. treu [401]. Trevor [647, 720]. Trial [34]. Trials [544, 66]. Tribe [137]. Trinity [597]. Trip [510, 513]. Trippett [553]. Trisha [408]. Trochu [113]. Troelenberg [582]. Tropical [571, 458]. Tropics [785]. Trotter [748]. Troubling [210]. Troyer [564]. Truth [511, 685, 370, 517]. Tsetse [56]. Tuber [371]. Tuchman [395]. Tug [666]. Tug-of-War [666]. Tumbe [642]. Tumor [560, 749]. Tuning [713]. Turchetti [131]. Turn [349]. Turnbull [41]. Turner [156, 551]. Tuscan [235]. Twelve [13]. Twelve-Year [13]. Twentieth [753, 249, 622, 361, 513, 652, 469, 664, 617]. Twentieth-Century [622, 652, 664]. Twenty [414, 631, 617]. Twenty-First [414, 631, 617]. Two [641, 560, 143]. Tycoons [347]. Tyndall [783]. Typhoid [558].

U.S [309]. U.S. [368, 324]. U.S.-Soviet [309]. Uganda [4]. Ulrike [400]. Ultimate [385]. Umgang [47]. Unbuilt [700]. Uncertainty [349]. Uncover [621]. Undercurrents [592]. Underground [371]. Understanding [51, 617, 281]. Undisciplined [148]. Ungureanu [342]. Union [255]. United [656, 124, 523, 253, 632, 760]. Universal [98]. Universe [283, 232].

Università [182]. Universitätsreform [43]. Universities [298]. University [188, 449, 189, 695, 735, 194]. Unknown [21]. Unknowns [238]. Unlock [296]. Unmusical [138]. Unnatural [517]. Untold [653]. Unusual [75]. Upheaval [321]. Urban [348, 677, 180]. Urey [457]. Urge [498]. Urrea [687]. Ursula [45, 495]. Use [325, 438, 490]. Used [719]. Useful [176, 446, 228]. Using [621]. Uta [448].

V [689, 258]. Vaccination [802]. Vallianatos [216, 542]. Value [454]. Values [579]. Vanden [167]. Vandendriessche [343]. Vanessa [299]. Vari [405]. Vasile [49]. Vasilii [398]. Vast [18]. Vaupel [158]. Vehlken [381]. veldtocht [50]. Venkat [604]. Vennen [63]. vento [38]. Vera [689, 508]. Verfahrensweisen [339]. Verhave [300]. Vermij [637, 489]. Vermitteln [239]. Vernacular [522, 403]. Vernetzen [239]. Veronica [584]. Versuchsanstalt [57]. Vertesi [794]. Vesalius [28]. Vicedo [611]. Viceregal [638]. Victor [787, 730]. Victoria [39, 652]. Victorian [686, 109, 348, 644, 558, 48, 422, 598]. Vienna [57]. Vietnam [640]. Views [72, 529]. Viktoria [455]. Vincenzo [196]. Vinchur [114]. Vinsel [124]. Violations [124]. Violence [596, 522, 511]. Viral [73]. Virdi [380]. Virology [749]. Virtual [372, 373]. Virtue [95]. Virus [200, 366, 749]. Viruses [188, 560, 560]. Vision [483, 60]. Visitors [479]. Visual [738, 537, 170]. Visualizing [510]. Vital [416]. Vitality [549]. Vitruvian [32]. Vivarium [57]. Vocabulary [168]. Voice [599]. Voices [408]. Volume [259, 487, 460, 275, 172, 658, 162, 227]. Vongsathorn [378]. Vos [724]. Vox [61]. Voyager [104]. Voyaging [102].

W [738, 69, 215, 692, 326, 68, 450]. Waddell [729]. Wailoo [653]. Walcher [91]. Wale [784]. Walking [593]. Wallace [93, 394, 110]. Wallis [177, 48]. Walton [553]. Wang [245]. Wants [214]. War [221, 84, 396, 128, 197, 666, 744, 603, 357, 65, 695, 324, 108, 360, 331, 608]. Warde [20]. Wardhaugh [174, 490]. Wardian [737]. Warfare [769, 87]. Waring [409]. Warming [254]. Warning [13]. Wars [697, 779]. Was [221]. Wasn’t [752]. Wasp [629]. Water [276, 766, 26, 269]. Waters [270, 273]. Watson [326]. Watt [105]. Way [211, 226]. Wayne [679]. Wealth [770]. Wear [686]. Weather [334, 39, 742]. Weber [434]. Wegener [183]. Wegener-Köppen [183]. Wei [679]. Weldon [559]. Welfare [569]. Wellerstein [656]. Wellington [582]. Wells [304, 492]. Welsh [487]. Werner [479, 90]. Werrett [171]. West [284, 589]. Westerink [301]. Western [423]. Wetlands [765]. Wheat [789]. White [583, 736]. Whiteness [718]. Whitesides [744]. Whither [709]. Who [539, 752, 204, 214, 786, 223]. Whole [72]. Whooley [222]. Whose [473]. Wigen [433]. Wild [187, 369]. Wiley [504]. Wilf [732]. Wilhelm [493, 43]. Wilkins [98]. Will [206]. Wille [50]. Willem [820, 58]. Willes [236]. William [139, 49, 99, 336, 113, 88, 597]. Williams [498, 558]. Willy [61]. Wilson [748]. Wind [766]. Winterer [433]. Wintersteen [654]. Winther [538]. Wirkungsgeschichte [285]. Wisdom [17]. Wise [302]. Wissenschaft [90]. wissenschaftlichen [47]. Wissenskultur [340]. Witchcraft [709, 34]. Withers [450]. Within [370]. Withoos [777]. Without [398, 390, 678]. Wladimir [183, 183]. Wolf [121]. Wolfe [188, 65]. Wolfram [588]. Womack [349]. Woman [458]. Women [539, 686, 793, 231, 399, 681, 67]. Wonder [706, 544]. Wonderful [452]. Wondrous [296]. Wood [296]. Woods [111]. Worboys [109]. Words [606]. Work [385, 321, 691, 556, 139]. Working [367, 218]. Works [818, 243, 34, 227, 422]. Workshop [56]. World [272, 635, 202, 84, 187, 352, 369, 108, 361, 744, 682, 695, 236, 494, 684, 206, 607, 652, 779, 810, 128, 197, 772, 811, 393, 35, 737, 176, 411, 642, 106, 735, 538, 234]. Worlds [687, 479]. Worldwide [28]. Would [786]. Wouldn’t [87]. Wragge [731]. Wragge-Morley [731]. Wretched [655]. Writerly [393]. Writing [127, 5]. Writings [542]. Written [630]. Wu [607]. Wylie [691].

Xiaoping [563]. xii [250]. xiii [215]. xiv [157]. XVIe [30]. xvii [215]. Xviiie [447]. xxiv [157]. xxx [449]. xxxiii [187].

Yale [140]. Yann [447]. Year [13, 668]. Years [689, 114, 641, 607, 201]. Yelda [490]. Yellow [650]. Yi [607]. Yi-Jui [607]. York [187, 215]. York/London [187, 215]. Yorkers [245]. Yu [679]. Yutzy [108].

Zacharias [542]. Zachary [548]. Zallen [550]. Zamparo [630]. Zanetti [32]. Zaunstöck [239]. Zeppelin [323]. Zeuge [448]. Zeuge-Buberl [448]. Zeynep [353]. Zieglerin [636]. Zimring [601]. Zipser [216]. Zones [763]. Zoo [570]. Zoo-Field [570]. Zoonoses [805]. Zootechnologies [381]. Zuidervaart [434]. Zur [63, 359, 47, 630]. Zwierlein [238].

References

Anonymous:2020:FCa

[1] Anonymous. Front cover. Isis, 111(1):??, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2020:M

[2] Anonymous. Masthead. Isis, 111(1):??, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hui:2020:LE

[3] Alexandra Hui and Matthew Lavine. Letter from the Editors. Isis, 111 (1):1–2, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cummiskey:2020:EEG

[4] Julia R. Cummiskey. “An Ecological Experiment on the Grand Scale”: Creating an experimental field in Bwamba, Uganda, 1942–1950. Isis, 111 (1):3–21, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Morel:2020:RGW

[5] Thomas Morel. De Re Geometrica: Writing, drawing, and preaching mathematics in early modern mines. Isis, 111(1):22–45, March 2020. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jin:2020:EEC

[6] Xiaoxing Jin. The evolution of evolutionism in China, 1870–1930. Isis, 111(1):46–66, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Goodman:2020:WIM

[7] Dena Goodman. What it meant to be Linnaean in Revolutionary France. Isis, 111(1):67–85, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

teHeesen:2020:HSS

[8] Anke te Heesen. History of science society distinguished lecture: Thomas S. Kuhn, earwitness: Interviewing and the making of a new history of science. Isis, 111(1):86–97, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Irwin:2020:ID

[9] Julia F. Irwin and Jenny Leigh Smith. Introduction: On disaster. Isis, 111(1):98–103, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lopes:2020:HGM

[10] Maria Margaret Lopes and Silvia Fernanda de Mendonça Figueirôa. The history of geology meets disasters: a Brazilian perspective. Isis, 111(1): 104–111, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Robinson:2020:SS

[11] Michael F. Robinson. Shackleton syndrome. Isis, 111(1):112–119, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mizelle:2020:HKD

[12] Richard M. Mizelle, Jr. Hurricane Katrina, diabetes, and the meaning of resiliency. Isis, 111(1):120–128, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Callison:2020:TYW

[13] Candis Callison. The twelve-year warning. Isis, 111(1):129–137, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nicolai:2020:LE

[14] Roel Nicolai. Letter to the Editor. Isis, 111(1):??, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nyhart:2020:ERA

[15] Lynn K. Nyhart. Essay review: Alex Csiszar. The Scientific Journal: Authorship and the Politics of Knowledge in the Nineteenth Century. Isis, 111(1):139–142, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cao:2020:BRR

[16] Deborah Cao. Book review: Roel Sterckx; Martina Siebert; Dagmar Schäfer, eds. Animals through Chinese History: Earliest Times to 1911. Isis, 111(1):143–144, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pietikainen:2020:BRS

[17] Petteri Pietikainen. Book review: Steven Connor. Madness of Knowledge: On Wisdom, Ignorance and Fantasies of Knowing. Isis, 111(1):144–145, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Robinson:2020:BRH

[18] Sam Robinson. Book review: Helen M. Rozwadowski. Vast Expanses: a History of the Oceans. Isis, 111(1):145–146, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Raj:2020:BRJ

[19] Kapil Raj. Book review: Johan Östling; Erling Sandmo; David Larsson Heidenblad; Anna Nilsson Hammar; Kari H. Nordberg, eds. Circulation of Knowledge: Explorations in the History of Knowledge. Isis, 111(1):146– 148, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fressoz:2020:BRP

[20] Jean-Baptiste Fressoz. Book review: Paul Warde. The Invention of Sustainability: Nature and Destiny, c. 1500–1870. Isis, 111(1):148–149, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Burman:2020:BRT

[21] Jeremy Trevelyan Burman, Jennifer L. Bazar, and Frederic Weizmann. Book review: Theodore M. Porter. Genetics in the Madhouse: The Unknown History of Human Heredity. Isis, 111(1):149–151, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Coste:2020:BRM

[22] Joël Coste. Book review: Mirko D. Grmek. Pathological Realities: Essays on Disease, Experiments, and History. Isis, 111(1):151–152, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Proust:2020:BRJ

[23] Christine Proust. Book review: Jöran Friberg; Farouk N. H. Al-Rawi. New Mathematical Cuneiform Texts. Isis, 111(1):152–154, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Raphals:2020:BRF

[24] Lisa Raphals. Book review: Florence Bretelle-Establet; Stéphane Schmitt, eds. Pieces and Parts in Scientific Texts. Isis, 111(1):154–155, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Israelowich:2020:BRE

[25] Ido Israelowich. Book review: Evelyne Samama. La médecine de guerre en Grèce ancienne. Isis, 111(1):??, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ledger:2020:BRA

[26] Marissa L. Ledger and Piers D. Mitchell. Book review: Ann Olga Koloski-Ostrow. The Archaeology of Sanitation in Roman Italy: Toilets, Sewers, and Water Systems. Isis, 111(1):156–158, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bruzzone:2020:BRD

[27] Raffaella Bruzzone. Book review: David Bennett. Medicine and Pharmacy in Byzantine Hospitals: a Study of the Extant Formularies. Isis, 111(1):

158–159, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fransen:2020:BRD

[28] Sietske Fransen. Book review: Dániel Margócsy; Mark Somos; Stephen N. Joffe. The Fabrica of Andreas Vesalius: a Worldwide Descriptive Census, Ownership, and Annotations of the 1543 and 1555 Editions. Isis, 111(1): 159–160, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Baumgartner:2020:BRA

[29] Ingrid Baumgärtner. Book review: Axelle Chassagnette. Savoir géographique et cartographie dans l’espace germanique protestant (1520– 1620). Isis, 111(1):161–162, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rommevaux-Tani:2020:BRS

[30] Sabine Rommevaux-Tani. Book review: Shin Higashi. Penser les mathématiques au XVIe siècle. Isis, 111(1):162–163, March 2020. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Braakenhielm:2020:BRM

[31] Carl-Reinhold Bråkenhielm. Book review: Michael E. Hobart. The Great Rift: Literacy, Numeracy, and the Religion–Science Divide. Isis, 111(1): 163–165, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pimentel:2020:BRC

[32] Juan Pimentel. Book review: Cristiano Zanetti. Janello Torriani and the Spanish Empire: a Vitruvian Artisan at the Dawn of the Scientific Revolution. Isis, 111(1):165–166, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vetter:2020:BRC

[33] Jeremy Vetter. Book review: Cameron B. Strang. Frontiers of Science: Imperialism and Natural Knowledge in the Gulf South Borderlands, 1500– 1850. Isis, 111(1):166–167, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Oldridge:2020:BRT

[34] Darren Oldridge. Book review: Todd H. J. Pettigrew; Stephanie M. Pettigrew; Jacques A. Bailly, eds. The Major Works of John Cotta: The Short Discovery (1612) and The Trial of Witchcraft (1616). Isis, 111(1):167– 169, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hatch:2020:BRP

[35] Robert Alan Hatch. Book review: Paula Findlen, ed. Empires of Knowledge: Scientific Networks in the Early Modern World. Isis, 111(1):169– 170, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shapiro:2020:BRT

[36] Barbara J. Shapiro. Book review: Thomas Harrison. The Ark of Studies. Isis, 111(1):170–171, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Romeiras:2020:BRN

[37] Francisco Malta Romeiras. Book review: Natacha Fabbri; Federica Favino, eds. Copernicus Banned: The Entangled Matter of the Anti-Copernican Decree of 1616. Isis, 111(1):172–173, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bernardi:2020:BRO

[38] Gabriella Bernardi. Book review: Ovidio Montalbani. Del vento e delle comete: Speculazioni accademiche. Isis, 111(1):173–174, March 2020. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Harper:2020:BRV

[39] Kristine C. Harper. Book review: Victoria C. Slonosky. Climate in the Age of Empire: Weather Observers in Colonial Canada. Isis, 111(1):174– 175, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jackson:2020:BRD

[40] Noel Jackson. Book review: Dahlia Porter. Science, Form, and the Problem of Induction in British Romanticism. Isis, 111(1):175–176, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roque:2020:BRP

[41] Ricardo Roque. Book review: Paul Turnbull. Science, Museums, and Collecting the Indigenous Dead in Colonial Australia. Isis, 111(1):177– 178, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Keene:2020:BRA

[42] Melanie Keene. Book review: Angela G. Ray; Paul Stob, eds. Thinking Together: Lecturing, Learning, and Difference in the Long Nineteenth Century. Isis, 111(1):178–179, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schultze:2020:BRH

[43] Sven Schultze. Book review: Heinz-Elmar Tenorth. Wilhelm von Humboldt: Bildungspolitik und Universitätsreform. Isis, 111(1):179–180, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Selcer:2020:BRN

[44] Perrin Selcer. Book review: Nicholas B. Breyfogle, ed. Eurasian Environments: Nature and Ecology in Imperial Russian and Soviet History. Isis, 111(1):181–182, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fiss:2020:BRC

[45] Andrew Fiss. Book review: Christopher Hollings; Ursula Martin; Adrian Rice. Ada Lovelace: The Making of a Computer Scientist. Isis, 111(1): 182–183, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sha:2020:BRS

[46] Richard C. Sha. Book reviews: Sidney Perkowitz; Eddy von Mueller, eds. Frankenstein: How A Monster Became an Icon: The Science and Enduring Allure of Mary Shelley’s Creation; Mary Shelley; David Guston; Ed Finn; Jason Scott Robert; Charles E. Robinson, eds. Frankenstein: Annotated for Scientists, Engineers, and Creators of All Kinds. Isis, 111 (1):183–186, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pieters:2020:BRA

[47] Toine Pieters. Book review: Anna Lindemann. Sigmund Freud, das “Cocain” und die Morphinisten: Ein Beitrag zur Geschichte der wissenschaftlichen und klinischen Praxis im Umgang mit Suchtmitteln (1850– 1890). Isis, 111(1):186–187, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Turbil:2020:BRJ

[48] Cristiano Turbil. Book review: Jennifer Wallis. Investigating the Body in the Victorian Asylum: Doctors, Patients, and Practices. Isis, 111(1):187– 188, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hung:2020:BRR

[49] Kuang-Chi Hung. Book reviews: Ronald Scott Vasile. William Stimpson and the Golden Age of American Natural History; Christopher Kemp. The Lost Species: Great Expeditions in the Collections of Natural History Museums. Isis, 111(1):188–190, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

GassoMiracle:2020:BRR

[50] Eulàlia Gassó Miracle. Book review: Robert-Jan Wille. Mannen van de microscoop: De laboratorium biologie op veldtocht in Nederland en Indië, 1840–1910. Isis, 111(1):191–192, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sauer:2020:BRE

[51] Tilman Sauer. Book review: Eric Johnson. Anxiety and the Equation: Understanding Boltzmann’s Entropy. Isis, 111(1):192–193, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Weiss:2020:BRE

[52] Martin P. M. Weiss. Book review: Emil Bessels. Polaris: The Chief Scientist’s Recollections of the American North Pole Expedition, 1871– 73. Isis, 111(1):193–194, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Walter:2020:BRC

[53] Scott A. Walter. Book review: Claudia E. Graf-Grossmann. Marcel Grossmann: For the Love of Mathematics. Isis, 111(1):194–195, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Singh:2020:BRP

[54] Rajinder Singh. Book review: Purabi Mukherji; Atri Mukhopadhyay. History of the Calcutta School of Physical Sciences. Isis, 111(1):195–196, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Eckert:2020:BRR

[55] Michael Eckert. Book review: Roy MacLeod; Russell G. Egdell; Elizabeth Bruton, eds. For Science, King, and Country: The Life and Legacy of Henry Moseley. Isis, 111(1):196–197, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Manton:2020:BRC

[56] John Manton. Book review: Clapperton Chakanetsa Mavhunga. The Mobile Workshop: The Tsetse Fly and African Knowledge Production. Isis, 111(1):198–199, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gliboff:2020:BRG

[57] Sander Gliboff. Book review: Gerd B. Müller, ed. Vivarium: Experimental, Quantitative, and Theoretical Biology at Vienna’s Biologische Versuchsanstalt. Isis, 111(1):199–200, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stanley:2020:BRJ

[58] Matthew Stanley. Book review: Jan Guichelaar. Willem de Sitter: Einstein’s Friend and Opponent. Isis, 111(1):200–201, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McLarty:2020:BRJ

[59] Colin McLarty. Book review: Jan von Plato. Saved from the Cellar: Gerhard Gentzen’s Shorthand Notes on Logic and Foundations of Mathematics. Isis, 111(1):??, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Seely:2020:BRT

[60] Bruce Seely. Book review: Thomas C. Lassman. Edward Condon’s Cooperative Vision: Science, Industry, and Innovation in Modern America. Isis, 111(1):203–204, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Walsh:2020:BRJ

[61] Lynda Walsh. Book reviews: Jared S. Buss. Willy Ley: Prophet of the Space Age; Lisa Vox. Existential Threats: American Apocalyptic Beliefs in the Technological Era. Isis, 111(1):204–206, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Krygier:2020:BRM

[62] John Krygier. Book review: Matthew H. Edney. Cartography: The Ideal and Its History. Isis, 111(1):207–208, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nieuwland:2020:BRI

[63] Ilja Nieuwland. Book review: Ina Heumann; Holger Stoecker; Marco Tamborini; Mareike Vennen. Dinosaurierfragmente: Zur Geschichte der Tendaguru-Expedition und ihrer Objekte, 1906–2018. Isis, 111(1):208– 209, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gaudilliere:2020:BRS

[64] Jean-Paul Gaudillière. Book review: Scott H. Podolsky. The Antibiotic Era: Reform, Resistance, and the Pursuit of a Rational Therapeutics. Isis, 111(1):209–211, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Whitesides:2020:BRA

[65] Greg Whitesides. Book review: Audra J. Wolfe. Freedom’s Laboratory: The Cold War Struggle for the Soul of Science. Isis, 111(1):??, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Weidman:2020:BRM

[66] Nadine Weidman. Book review: Matthew Oram. The Trials of Psychedelic Therapy: LSD Psychotherapy in America. Isis, 111(1):212–213, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Forstner:2020:BRC

[67] Christian Forstner. Book review: Chris Talbot, ed. David Bohm: Causality and Chance, Letters to Three Women. Isis, 111(1):213–214, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mold:2020:BRC

[68] Alex Mold. Book review: Cynthia Carson Bisbee; Paul Bisbee; Erika Dyck; Patrick Farrell; James Sexton; James W. Spisak, eds. Psychedelic Prophets: The Letters of Aldous Huxley and Humphry Osmond. Isis, 111 (1):214–215, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Inkpen:2020:BRP

[69] Dani K. Inkpen. Book review: Philip W. Clements. Science in an Extreme Environment: The 1963 American Mount Everest Expedition. Isis, 111(1): 215–216, March 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2020:FCb

[70] Anonymous. Front cover. Isis, 111(2):??, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2020:FMa

[71] Anonymous. Front matter. Isis, 111(2):??, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bakker:2020:MCM

[72] Frederik Bakker and Carla Rita Palmerino. Motion to the center or motion to the whole? Plutarch’s views on gravity and their influence on Galileo. Isis, 111(2):217–238, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Meerwijk:2020:VID

[73] Maurits Bastiaan Meerwijk. Viral imagery of dengue fever in the age of bacteriology. Isis, 111(2):239–263, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heng:2020:CMB

[74] Wen Heng. From the China Medical Board to the China Foundation: The network of interlocking patronage and China’s new scientific community, 1920s–1930s. Isis, 111(2):264–283, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Evans:2020:MUB

[75] Andrew D. Evans. “Most unusual” beauty contests: Nordic photographic competitions and the construction of a public for German race science, 1926–1935. Isis, 111(2):284–309, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mukharji:2020:RDC

[76] Projit Bihari Mukharji, Myrna Perez Sheldon, Elise K. Burton, Sebastián Gil-Riaño, Terence Keel, Emily Merchant, Wangui Muigai, Ahmed Ragab, and Suman Seth. A roundtable discussion on collecting demographics data. Isis, 111(2):310–353, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2020:HSS

[77] Anonymous. History of Science Society Annual Meeting, 2019. Isis, 111 (2):354–361, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Martins:2020:ACK

[78] Roberto de Andrade Martins, Cibelle Celestino Silva, and Maria Elice Brzezinski Prestes. Anna Carolina Krebs Pereira Regner (1947– 2020). Isis, 111(2):362–364, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Henry:2020:NSP

[79] John Henry. Nicolas Steno (1638–1686): a polymath reassessed. Isis, 111(2):365–367, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hale:2020:MMS

[80] Piers J. Hale. The making of a man of science: Darwin’s development in a transformative time. Isis, 111(2):368–370, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barth:2020:BRP

[81] Howard G. Barth. Book review: Patrick Forterre; trans. by Teresa Lavender Fagan. Microbes from Hell. Isis, 111(2):371–372, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Channell:2020:BRE

[82] David F. Channell. Book review: Eric Schatzberg. Technology: Critical History of a Concept. Isis, 111(2):372–373, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gossin:2020:BRH

[83] Pamela Gossin. Book review: Helge Kragh; Malcolm S. Longair, eds. The Oxford Handbook of the History of Modern Cosmology. Isis, 111(2): 373–374, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cahan:2020:BRJ

[84] David Cahan. Book review: John Gascoigne. Science and the State: From the Scientific Revolution to World War II. Isis, 111(2):375–376, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Amelin:2020:BRJ

[85] Olov Amelin. Book review: Johan Kärnfelt; Karl Grandin; Solveig Jülich, eds. Knowledge in Motion: The Royal Swedish Academy of Sciences and the Making of Modern Society. Isis, 111(2):376–377, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Esposito:2020:BRA

[86] Maurizio Esposito. Book review: Amanda Jo Goldstein. Sweet Science: Romantic Materialism and the New Logics of Life. Isis, 111(2):377–378, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ohlers:2020:BRJ

[87] R. Clinton Ohlers. Book review: Jeff Hardin; Ronald L. Numbers; Ronald A. Binzley, eds. The Warfare between Science and Religion: The Idea That Wouldn’t Die. Isis, 111(2):379–380, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Verburgt:2020:BRW

[88] Lukas M. Verburgt. Book review: William Boos; Florence S. Boos, ed. Metamathematics and the Philosophical Tradition. Isis, 111(2):380–381, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Harrison:2020:BRE

[89] Thomas Harrison. Book review: Ewen Bowie, ed. Herodotus: Narrator, Scientist, Historian. Isis, 111(2):382–383, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tieleman:2020:BRW

[90] Teun Tieleman. Book review: Werner Albert Golder, ed. Celsus und die antike Wissenschaft. Isis, 111(2):383–384, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Runciman:2020:BRC

[91] David Runciman. Book review: C. Philipp E. Nothaft. Walcher of Malvern, De lunationibus and De Dracone: Study, Edition, Translation, and Commentary. Isis, 111(2):384–385, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hayden:2020:BRS

[92] Judy A. Hayden. Book review: Sheila J. Nayar. Renaissance Responses to Technological Change. Isis, 111(2):386–387, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jacobson:2020:BRT

[93] Nicholas A. Jacobson. Book review: Todd Timberlake; Paul Wallace. Finding Our Place in the Solar System: The Scientific Story of the Copernican Revolution. Isis, 111(2):387–388, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Omodeo:2020:BRM

[94] Pietro Daniel Omodeo and Omar Del Nonno. Book review: Manuel Mertens. Magic and Memory in Giordano Bruno: The Art of a Heroic Spirit. Isis, 111(2):388–390, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Reeves:2020:BRH

[95] Eileen Reeves. Book review: Horst Bredekamp; trans. by Mitch Cohen. Galileo’s Thinking Hand: Mannerism, Anti-Mannerism, and the Virtue of Drawing in the Foundation of Early Modern Science. Isis, 111(2):390– 391, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brixius:2020:BRR

[96] Dorit Brixius. Book review: Raphaële Garrod; Paul J. Smith, eds. Natural History in Early Modern France: The Poetics of an Epistemic Genre. Isis, 111(2):391–392, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rankin:2020:BRM

[97] William Rankin. Book review: Margaret E. Schotte. Sailing School: Navigating Science and Skill, 1550–1800. Isis, 111(2):392–394, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Oosterhoff:2020:BRJ

[98] Richard Oosterhoff. Book review: James Dougal Fleming. The Mirror of Information in Early Modern England: John Wilkins and the Universal Character. Isis, 111(2):394–395, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nummedal:2020:BRW

[99] Tara Nummedal. Book review: William R. Newman. Newton the Alchemist: Science, Enigma, and the Quest for Nature’s “Secret Fire”. Isis, 111(2):395–396, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stewart:2020:BRM

[100] Larry Stewart. Book review: Margaret C. Jacob. The Secular Enlightenment. Isis, 111(2):396–397, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Norrgren:2020:BRA

[101] Hilde Norrgrén. Book review: Ane Ohrvik. Medicine, Magic and Art in Early Modern Norway: Conceptualizing Knowledge. Isis, 111(2):398– 399, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anderson:2020:BRA

[102] Katharine Anderson. Book review: Adriana Craciun; Mary Terrall, eds. Curious Encounters: Voyaging, Collecting, and Making Knowledge in the Long Eighteenth Century. Isis, 111(2):399–400, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marks:2020:BRS

[103] Jonathan Marks. Book review: Suman Seth. Difference and Disease: Medicine, Race, and the Eighteenth-Century British Empire. Isis, 111(2): 400–401, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kabat:2020:BRJ

[104] Alan R. Kabat. Book review: Jürgen Goldstein; Anne Janusch. Georg Forster: Voyager, Naturalist, Revolutionary. Isis, 111(2):401–402, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Torrens:2020:BRD

[105] Hugh Torrens. Book review: David Philip Miller. The Life and Legend of James Watt: Collaboration, Natural Philosophy, and the Improvement of the Steam Engine. Isis, 111(2):402–403, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Seitz:2020:BRM

[106] Jonathan Seitz. Book review: Maria Pia Donato, ed. Medicine and the Inquisition in the Early Modern World. Isis, 111(2):404–405, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sheppard:2020:BRB

[107] Kathleen Sheppard. Book review: Brent Maner. Germany’s Ancient Pasts: Archaeology and Historical Interpretation since 1700. Isis, 111(2): 405–406, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Labbe:2020:BRH

[108] Morgane Labbé. Book review: Heinrich Hartmann; trans. by Ellen Yutzy Glebe. The Body Populace: Military Statistics and Demography in Europe before the First World War. Isis, 111(2):406–407, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hobbins:2020:BRM

[109] Peter Hobbins. Book review: Michael Worboys; Julie-Marie Strange; Neil Pemberton. The Invention of the Modern Dog: Breed and Blood in Victorian Britain. Isis, 111(2):407–409, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pearn:2020:BRP

[110] Alison M. Pearn. Book review: Patrick Armstrong. Alfred Russel Wallace. Isis, 111(2):409–410, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Benson:2020:BRA

[111] Etienne S. Benson. Book review: Abigail Woods; Michael Bresalier; Angela Cassidy; Rachel Mason Dentinger. Animals and the Shaping of Modern Medicine: One Health and Its Histories. Isis, 111(2):410–411, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Plotnitsky:2020:BRP

[112] Arkady Plotnitsky. Book review: Peter Roquette. The Riemann Hypothesis in Characteristic p in Historical Perspective. Isis, 111(2):411–412, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Serina:2020:BRT

[113] Florent Serina. Book review: Thibaud Trochu. William James: Une autre histoire de la psychologie. Isis, 111(2):412–413, June 2020. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Derksen:2020:BRA

[114] Maarten Derksen. Book review: Andrew J. Vinchur. The Early Years of Industrial and Organizational Psychology. Isis, 111(2):??, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Turner:2020:BRJ

[115] Anthony Turner. Book review: Jonathan Betts. Marine Chronometers at Greenwich: a Catalogue of Marine Chronometers at the National Maritime Museum, Greenwich. Isis, 111(2):415–416, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Freire:2020:BRK

[116] Olival Freire Jr. Book review: Klaus Hentschel. Photons: The History and Mental Models of Light Quanta. Isis, 111(2):416–417, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nichols:2020:BRD

[117] Tiffany Nichols. Book review: Daniel Kennefick. No Shadow of a Doubt: The 1919 Eclipse That Confirmed Einstein’s Theory of Relativity. Isis, 111(2):417–418, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Eckes:2020:BRD

[118] Christophe Eckes. Book review: David Aubin. L’élite sous la mitraille: Les normaliens, les mathématiques et la Grande Guerre 1900–1925. Isis, 111(2):418–419, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wu:2020:BRE

[119] Harry Yi-Jui Wu. Book review: Emily Baum. The Invention of Madness: State, Society, and the Insane in Modern China. Isis, 111(2):420–421, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Johnson:2020:BRS

[120] Jeffrey Allan Johnson. Book review: Seth C. Rasmussen, ed. Igniting the Chemical Ring of Fire: Historical Evolution of the Chemical Communities of the Pacific Rim. Isis, 111(2):421–422, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Freidenfelds:2020:BRJ

[121] Lara Freidenfelds. Book review: Jacqueline H. Wolf. Cesarean Section: an American History of Risk, Technology, and Consequence. Isis, 111(2): 422–423, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cowles:2020:BRJ

[122] Henry M. Cowles. Book review: John L. Rudolph. How We Teach Science: What’s Changed, and Why It Matters. Isis, 111(2):424–425, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bocking:2020:BRR

[123] Stephen Bocking. Book review: Robert E. Kohler. Inside Science: Stories from the Field in Human and Animal Science. Isis, 111(2):425–426, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Corn:2020:BRL

[124] Joseph J. Corn. Book review: Lee Vinsel. Moving Violations: Automobiles, Experts, and Regulations in the United States. Isis, 111(2):426–428, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Priestley:2020:BRM

[125] Rebecca Priestley. Book review: Mary Harris. Rocks, Radio, and Radar: The Extraordinary Scientific, Social, and Military Life of Elizabeth Alexander. Isis, 111(2):428–429, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Burton:2020:BRH

[126] Elise K. Burton. Book review: Hugo Iltis. Race, Genetics, and Science: Resisting Racism in the 1930s. Isis, 111(2):429–430, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Herran:2020:BRJ

[127] Néstor Herran. Book review: John Krige, ed. How Knowledge Moves: Writing the Transnational History of Science and Technology. Isis, 111 (2):430–433, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hoffenberg:2020:BRA

[128] Peter H. Hoffenberg. Book review: Arthur P. Molella; Scott Gabriel Knowles, eds. World’s Fairs in the Cold War: Science, Technology, and the Culture of Progress. Isis, 111(2):433–434, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rubinson:2020:BRJ

[129] Paul Rubinson. Book review: J. Pat Brown; B. C. D. Lipton; Michael Morisy, eds. Scientists under Surveillance: The FBI Files. Isis, 111(2): 434–435, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ellis:2020:BRB

[130] Heather Ellis. Book review: Ben Barres. The Autobiography of a Transgender Scientist. Isis, 111(2):??, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hamblin:2020:BRS

[131] Jacob Darwin Hamblin. Book review: Simone Turchetti. Greening the Alliance: The Diplomacy of NATO’s Science and Environmental Initiatives. Isis, 111(2):437–438, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rheinberger:2020:BRO

[132] Hans-Jörg Rheinberger. Book review: Onur Erdur. Die epistemologischen Jahre: Philosophie und Biologie in Frankreich, 1960–1980. Isis, 111(2): 438–439, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mody:2020:BRS

[133] Cyrus C. M. Mody. Book review: Subrata Dasgupta. The Second Age of Computer Science: From Algol Genes to Neural Nets. Isis, 111(2): 439–440, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ebeling:2020:BRB

[134] Mary F. E. Ebeling. Book review: Bruno J. Strasser. Collecting Experiments: Making Big Data Biology. Isis, 111(2):440–441, June 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2020:FCc

[135] Anonymous. Front cover. Isis, 111(3):??, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2020:FMb

[136] Anonymous. Front matter. Isis, 111(3):??, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Greene:2020:IRI

[137] Jeremy A. Greene, Victor Braitberg, and Gabriella Maya Bernadett. Innovation on the Reservation: Information technology and health systems research among the Papago Tribe of Arizona, 1965–1980. Isis, 111(3):443– 470, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kromhout:2020:UEG

[138] Melle Jan Kromhout. The unmusical ear: Georg Simon Ohm and the mathematical analysis of sound. Isis, 111(3):471–492, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Scheffler:2020:BBH

[139] Robin Wolfe Scheffler. Brightening biochemistry: Humor, identity, and scientific work at the Sir William Dunn Institute of Biochemistry, 1923– 1931. Isis, 111(3):493–514, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ossmer:2020:NDP

[140] Carola Ossmer. Normal development: The photographic dome and the children of the Yale Psycho-Clinic. Isis, 111(3):515–541, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Henry:2020:PSC

[141] John Henry. Primary and secondary causation in Samuel Clarke’s and Isaac Newton’s theories of gravity. Isis, 111(3):542–561, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Berryman:2020:HAP

[142] Sylvia Berryman. How Archimedes proposed to move the Earth. Isis, 111 (3):562–567, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rader:2020:ICP

[143] Karen Rader. Introduction: The changing pedagogical landscapes of history of science and the “two cultures”. Isis, 111(3):568–575, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Powers:2020:HCC

[144] John C. Powers. The history of chemistry in chemical education. Isis, 111 (3):576–581, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gamito-Marques:2020:PHS

[145] Daniel Gamito-Marques. In praise of a historical storytelling approach in science education. Isis, 111(3):582–587, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marsh:2020:CCH

[146] Allison Marsh and Bethany Johnson. Crash course history of science: Popular science for general education? Isis, 111(3):588–594, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dennis:2020:BHL

[147] David Brandon Dennis, R. A. Lawson, and Jessica M. Pisano. Bringing history into the lab: a new approach to scientific learning in general education. Isis, 111(3):595–605, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Uchacz:2020:REM

[148] Tianna Helena Uchacz. Reconstructing early modern artisanal epistemologies and an “undisciplined” mode of inquiry. Isis, 111(3):606–613, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bowcutt:2020:CTB

[149] Frederica Bowcutt and Tamara Caulkins. Co-teaching botany and history: An interdisciplinary model for a more inclusive curriculum. Isis, 111(3): 614–622, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hamilton:2020:HES

[150] Vivien Hamilton and Daniel M. Stoebel. History in the education of scientists: Encouraging judgment and social action. Isis, 111(3):623–630, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DiMoia:2020:EAM

[151] John P. DiMoia and Hiromi Mizuno. Eloge: Aaron S. Moore (1972–2019). Isis, 111(3):631–633, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pangburn:2020:EPH

[152] Kris Pangburn. Eloge: Peter Hanns Reill (1938–2019). Isis, 111(3):634– 635, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rosner:2020:EBG

[153] David Rosner. Eloge: Barbara Gutmann Rosenkrantz (1923–2014). Isis, 111(3):636–640, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fleck:2020:LE

[154] Robert Fleck. Letter to the Editor. Isis, 111(3):641–642, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bowler:2020:MN

[155] Peter J. Bowler. Material nature. Isis, 111(3):643–646, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bennett:2020:BRA

[156] Jim Bennett. Book review: Anthony Turner. Mathematical Instruments in the Collections of the Bibliothèque Nationale de France. Isis, 111(3):647– 648, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Castelao-Lawless:2020:BRG

[157] Teresa Castelão-Lawless. Book reviews: Gaston Bachelard, Atomistic Intuitions: An Essay on Classification; Translated and with an introduction by Roch C. Smith. (SUNY Series in Contemporary French Thought.) xxiv + 127 pp., notes, index. Albany: SUNY Press, 2018. $80 (cloth). Paperback and e-book available. Hans-Jörg Rheinberger. The Hand of the Engraver: Albert Flocon Meets Gaston Bachelard. Translated by Kate Sturge. (SUNY Series, Intersections: Philosophy and Critical Theory.) xiv + 111 pp., illus., notes, bibl., index. Albany: SUNY Press, 2018. $75 (cloth). Paperback and e-book available. Isis, 111(3):648–650, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Arnold:2020:BRE

[158] David Arnold. Book review: Ernst Homburg; Elisabeth Vaupel, eds. Hazardous Chemicals: Agents of Risk and Change, 1800–2000. Isis, 111(3): 651–652, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Birkhead:2020:BRR

[159] Tim Birkhead. Book review: Roger J. Lederer. The Art of the Bird: The History of Ornithological Art through Forty Artists. Isis, 111(3):652–653, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Spiller:2020:BRM

[160] James A. Spiller. Book review: Michael Bravo. North Pole: Nature and Culture. Isis, 111(3):653–654, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ankeny:2020:BRM

[161] Rachel A. Ankeny. Book review: Michel Anctil. Luminous Creatures: The History and Science of Light Production in Living Organisms. Isis, 111 (3):654–655, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Perilli:2020:BRA

[162] Lorenzo Perilli. Book review: Alexander Jones; Liba Taub, eds. The Cambridge History of Science. Volume 1: Ancient Science. Isis, 111(3): 656–657, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Steele:2020:BRK

[163] John Steele. Book review: Kathryn Stevens. Between Greece and Babylonia: Hellenistic Intellectual History in Cross-Cultural Perspective. Isis, 111(3):657–658, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Runstedler:2020:BRN

[164] Curtis Runstedler. Book review: E. Nicolaidis, ed. Greek Alchemy from Late Antiquity to Early Modernity. Isis, 111(3):659–660, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Falk:2020:BRJ

[165] Seb Falk. Book review: Josefina Rodríguez–Arribas; Charles Burnett; Silke Ackermann; Ryan Szpiech, eds. Astrolabes in Medieval Cultures. Isis, 111(3):660–661, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schmitz-Esser:2020:BRD

[166] Romedio Schmitz-Esser. Book review: David J. Collins, ed. The Sacred and the Sinister: Studies in Medieval Religion and Magic. Isis, 111(3): 661–662, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Smoller:2020:BRC

[167] Laura Ackerman Smoller. Book review: Carlos Steel; Steven Vanden Broecke; David Juste; Shlomo Sela, eds. The Astrological Autobiography of a Medieval Philosopher: Henry Bate’s Nativitas (1280–81). Isis, 111 (3):662–663, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bynum:2020:BRJ

[168] W. F. Bynum. Book review: Juhani Norri (Compiler). Dictionary of Medical Vocabulary in English, 1375–1550: Body Parts, Sicknesses, Instruments, and Medicinal Preparations. Isis, 111(3):664–665, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Turpin:2020:BRA

[169] Adriana Turpin. Book review: Angelica Groom. Exotic Animals in the Art and Culture of the Medici Court in Florence. Isis, 111(3):665–666, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schuster:2020:BRD

[170] John A. Schuster. Book review: Domenico Bertoloni Meli. Mechanism: a Visual, Lexical, and Conceptual History. Isis, 111(3):666–667, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hendriksen:2020:BRS

[171] Marieke Hendriksen. Book review: Simon Werrett. Thrifty Science: Making the Most of Materials in the History of Experiment. Isis, 111(3):668– 669, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hamel:2020:BRC

[172] Jürgen Hamel and Claire Conklin Sabel. Book review: Chantal Grell, ed. Correspondance de Johannes Hevelius.Volume 1: Prolégomènes critiques; Chantal Grell, ed. Correspondance de Johannes Hevelius. Volume 2: Correspondance avec la cour de France et ses agents, avec un dossier sur la querelle de la comète de 1664–1665. Isis, 111(3):669–671, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hamlin:2020:BRC

[173] Christopher Hamlin. Book review: Christopher Braider. Experimental Selves: Person and Experience in Early Modern Europe. Isis, 111(3):671– 672, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Boantza:2020:BRB

[174] Victor D. Boantza. Book review: Benjamin Wardhaugh. Gunpowder and Geometry: The Life of Charles Hutton: Pit Boy, Mathematician, and Scientific Rebel. Isis, 111(3):672–674, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barnett:2020:BRD

[175] Lydia Barnett. Book review: Deborah R. Coen. Climate in Motion: Science, Empire, and the Problem of Scale. Isis, 111(3):674–675, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Osborne:2020:BRP

[176] Michael A. Osborne. Book review: Pete Minard. PAll Things Harmless, Useful, and Ornamental: Environmental Transformation through Species Acclimatization, from Colonial Australia to the World. Isis, 111(3):675– 677, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gilman:2020:BRA

[177] Sander L. Gilman. Book review: Amelia Bonea; Melissa Dickson; Sally Shuttleworth; Jennifer Wallis. Anxious Times: Medicine and Modernity in Nineteenth-Century Britain. Isis, 111(3):677–678, September 2020. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sivils:2020:BRR

[178] Matthew Wynn Sivils. Book review: Richard J. King. Ahab’s Rolling Sea: a Natural History of Moby-Dick. Isis, 111(3):678–679, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hatfield:2020:BRD

[179] Gary Hatfield. Book review: David E. Leary. The Routledge Guidebook to James’s. Isis, 111(3):679–680, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Levitt:2020:BRO

[180] Theresa Levitt. Book review: Oliver Hochadel; Agustí Nieto-Galan, eds. Urban Histories of Science: Making Knowledge in the City, 1820–1940. Isis, 111(3):??, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Peckhaus:2020:BRB

[181] Volker Peckhaus. Book review: Brendan Dooley, ed. The Continued Exercise of Reason: Public Addresses by George Boole. Isis, 111(3):682–683, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Guerraggio:2020:BRG

[182] Angelo Guerraggio. Book review: Giorgio Israel, ed. Correspondence of Luigi Cremona (1830–1903): Conserved in the Department of Mathematics, “Sapienza” Università di Roma. Isis, 111(3):683–684, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lehmann:2020:BRE

[183] Philipp N. Lehmann. Book review: Else Wegener-Köppen, ed. Wladimir Köppen–Scholar for Life/Wladimir Köppen —- Ein Gelehrtenleben für die Meteorologie. Isis, 111(3):684–685, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Huang:2020:BRJ

[184] Hsiang-Fu Huang. Book review: Joshua Nall. News from Mars: Mass Media and the Forging of a New Astronomy, 1860–1910. Isis, 111(3):685– 686, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lewis:2020:BRP

[185] Daniel Lewis. Book review: Paolo Palmieri, trans. A Translation of Luigi Paolucci’s On Birdsong: Phenomenology, Animal Psychology, and Biology. Isis, 111(3):686–688, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dyck:2020:BRA

[186] Erika Dyck. Book review: Andrew Scull. Psychiatry and Its Discontents. Isis, 111(3):688–689, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Campos:2020:BRJ

[187] Luis Campos. Book review: James Mahaffey. Atomic Adventures: Secret Islands, Forgotten N-Rays, and Isotopic Murder — A Journey into the Wild World of Nuclear Science. xxxiii + 363 pp., illus., bibl., index. New York/London: Pegasus Books, 2017. $29.95 (cloth). E-book and paperback available. Isis, 111(3):689–690, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Greco:2020:BRR

[188] Cinzia Greco. Book review: Robin Wolfe Scheffler. A Contagious Cause: The American Hunt for Cancer Viruses and the Rise of Molecular Medicine. ix + 379 pp., bibl., index. Chicago/London: University of Chicago Press, 2019. $40 (paper). Hardback and e-book available. Isis, 111 (3):690–691, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lowy:2020:BRM

[189] Ilana Löwy. Book review: Mathias Grote. Membranes to Molecular Machines: Active Matter and the Remaking of Life. 286 pp., bibl., index. Chicago/London: University of Chicago Press, 2019. $45 (cloth). E-book available. Isis, 111(3):692–693, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Howell:2020:BRA

[190] Yvonne Howell. Book review: Anya Bernstein. The Future of Immortality: Remaking Life and Death in Contemporary Russia. Isis, 111(3):693–694, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Prasad:2020:BRB

[191] Amit Prasad. Book review: Banu Subramaniam. Holy Science: The Biopolitics of Hindu Nationalism. Isis, 111(3):694–695, September 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2020:FCd

[192] Anonymous. Front cover. Isis, 111(4):??, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2020:FMc

[193] Anonymous. Front matter. Isis, 111(4):??, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Martin:2020:SCC

[194] Joseph D. Martin. The simple and courageous course: Industrial patronage of basic research at the University of Chicago, 1945–1953. Isis, 111 (4):697–716, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Foks:2020:CFI

[195] Freddy Foks. Constructing the field in interwar social anthropology: Power, personae, and paper technology. Isis, 111(4):717–739, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Finocchiaro:2020:VGM

[196] Maurice A. Finocchiaro. Vincenzo Galilei’s musicology and Galileo’s science: Methodological comparison and contrast. Isis, 111(4):740–758, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jacobs:2020:MOP

[197] Noortje Jacobs. A moral obligation to proper experimentation: Research ethics as epistemic filter in the aftermath of World War II. Isis, 111 (4):759–780, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hui:2020:EI

[198] Alexandra Hui and Matthew Lavine. Editors’ introduction. Isis, 111 (4):781–782, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bouk:2020:MC

[199] Dan Bouk. Materializing COVID. Isis, 111(4):783–786, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Aronowitz:2020:LLV

[200] Robert Aronowitz. Learning to live with the virus. Isis, 111(4):787–790, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vargha:2020:EYT

[201] Dora Vargha. Epidemic years: a third look. Isis, 111(4):791–794, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brazelton:2020:RHM

[202] Mary Augusta Brazelton. Recovering hygienic modernity in the world of COVID-19. Isis, 111(4):795–799, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tresch:2020:SSP

[203] John Tresch. Sickness and sweetness and power. Isis, 111(4):800–804, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hamlin:2020:WCT

[204] Christopher Hamlin. So who cares? Taking Livingston out of Africa. Isis, 111(4):805–808, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marcus:2020:RPA

[205] Hannah Marcus. Revisiting the plague in the age of Galileo. Isis, 111 (4):809–813, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Smocovitis:2020:SBP

[206] Vassiliki Betty Smocovitis. Someday a big plague will come: Laurie Garrett’s the coming plague: newly emerging diseases in a world out of balance. Isis, 111(4):814–817, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Steere-Williams:2020:MMP

[207] Jacob Steere-Williams. A “menace’ or a martyr to the public’s health? Isis, 111(4):818–821, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Warner:2020:RGG

[208] John Harley Warner. Rereading the gospel of germs during a pandemic. Isis, 111(4):822–825, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Thompson:2020:FDC

[209] Courtney E. Thompson. Finding Deborah: Centering patients and placing emotion in the history of disease. Isis, 111(4):826–829, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bonhomme:2020:TPC

[210] Edna Bonhomme. Troubling (post)colonial histories of medicine: Toward a praxis of the human. Isis, 111(4):830–833, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anderson:2020:WWL

[211] Warwick Anderson. The way we live now? Isis, 111(4):834–837, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Weiner:2020:EEI

[212] Douglas R. Weiner, Lloyd Ackert, Stephen C. Brain, Loren R. Graham, and Paul Josephson. Eloge: Eduard Izrailevich Kolchinskii (1944–2020). Isis, 111(4):838–839, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ash:2020:BRD

[213] Mitchell G. Ash. Book review: David Cahan’s Helmholtz: History of Science in European History. Isis, 111(4):840–844, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Phillips:2020:WWM

[214] Christopher J. Phillips. Who wants to be a mathematician? Isis, 111 (4):845–848, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Klein:2020:BRO

[215] Wouter Klein. Book reviews: Ole Peter Grell; Andrew Cunningham; Jon Arrizabalaga, eds. “It All Depends on the Dose”: Poisons and Medicines in European History. (The History of Medicine in Context.) xiii + 244 pp., figs., tables, index. New York/London: Routledge, 2018. $155 (cloth); ISBN 978-1-138-69761-4. E-book available; Frederick W. Gibbs. Poison, Medicine, and Disease in Late Medieval and Early Modern Europe. xvii + 313 pp., bibl., index. New York/London: Routledge, 2018. $155 (cloth); ISBN 978-1-4724-2039-8. Isis, 111(4):849–851, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Guardasole:2020:BRP

[216] Alessia Guardasole. Book review: Petros Bouras-Vallianatos; Barbara Zipser, eds. Brill’s Companion to the Reception of Galen. Isis, 111(4): 851–852, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Margocsy:2020:BRP

[217] Dániel Margócsy. Book reviews: Patrick Manning; Daniel Rood, eds. Global Scientific Practice in an Age of Revolutions, 1750–1850; Patrick Manning; Mat Savelli, eds. Global Transformations in the Life Sciences, 1945–1980; Patrick Manning; Abigail Owen, eds. Knowledge in Translation: Global Patterns of Scientific Exchange, 1000–1800 CE. Isis, 111 (4):852–855, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hess:2020:BRC

[218] Volker Hess. Book review: Carla Bittel; Elaine Leong; Christine von Oertzen, eds. Working with Paper: Gendered Practices in the History of Knowledge. Isis, 111(4):855–856, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schatzberg:2020:BRD

[219] Eric Schatzberg. Book review: David F. Channell. The Rise of Engineering Science: How Technology Became Scientific. Isis, 111(4):857–858, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lanzoni:2020:BRR

[220] Susan Lanzoni. Book review: Roger Smith. The Sense of Movement: An Intellectual History. Isis, 111(4):858–859, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Alberti:2020:BRT

[221] Fay Bound Alberti. Book review: Thomas Dodman. What Nostalgia Was: War, Empire, and the Time of a Deadly Emotion. Isis, 111(4):859–860, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sadowsky:2020:BRO

[222] Jonathan Sadowsky. Book review: Owen Whooley. On the Heels of Ignorance: Psychiatry and the Politics of Not Knowing. Isis, 111(4):861–862, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Suarez-Diaz:2020:BRE

[223] Edna Suárez-Díaz. Book review: Esha Shah. Who Is the Scientist-Subject? Affective History of the Gene. Isis, 111(4):862–863, December 2020. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Suh:2020:BRA

[224] Soyoung Suh. Book review: Angela Ki Che Leung; Izumi Nakayama, eds. Gender, Health, and History in Modern East Asia. Isis, 111(4):863–864, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Simeoni:2020:BRT

[225] Luca Simeoni. Book review: Theokritos Kouremenos. Plato’s Forms, Mathematics, and Astronomy. Isis, 111(4):864–866, December 2020. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Husson:2020:BRJ

[226] Matthieu Husson. Book review: Jacqueline Feke. Ptolemy’s Philosophy: Mathematics as a Way of Life. Isis, 111(4):866–867, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Petit:2020:BRK

[227] Caroline Petit. Book review: Keith Andrew Stewart. Galen’s Theory of Black Bile: Hippocratic Tradition, Manipulation, Innovation; P. N. Singer; Philip J. van der Eijk, eds. Galen: Works on Human Nature. Volume 1: Mixtures (De temperamentis). Isis, 111(4):867–869, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Savage-Smith:2020:BRL

[228] Emilie Savage-Smith. Book review: Lucia Raggetti. Īsā ibn Alī’s Book on the Useful Properties of Animal Parts: Edition, Translation, and Study of a Fluid Tradition. Isis, 111(4):869–870, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nothaft:2020:BRJ

[229] C. Philipp E. Nothaft. Book review: José Chabás. Computational Astronomy in the Middle Ages: Sets of Astronomical Tables in Latin. Isis, 111 (4):870–871, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cohen:2020:BRJ

[230] Esther Cohen. Book review: Julie Orlemanski. Symptomatic Subjects: Bodies, Medicine, and Causation in the Literature of Late Medieval England. Isis, 111(4):871–872, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marcus:2020:BRS

[231] Hannah Marcus. Book review: Sharon T. Strocchia. Forgotten Healers: Women and the Pursuit of Health in Late Renaissance Italy. Isis, 111 (4):873–874, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barth:2020:BRM

[232] Howard G. Barth. Book review: Michael Brooks. The Quantum Astrologer’s Handbook: a History of the Renaissance Mathematics That Birthed Imaginary Numbers, Probability, and the New Physics of the Universe. Isis, 111(4):874–875, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rankin:2020:BRH

[233] Alisha Rankin. Book review: Hannah Murphy. A New Order of Medicine: The Rise of Physicians in Reformation Nuremberg. Isis, 111(4):875–876, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Winterbottom:2020:BRM

[234] Anna Winterbottom. Book review: Matthew James Crawford; Joseph M. Gabriel, eds. Drugs on the Page: Pharmacopoeias and Healing Knowledge in the Early Modern Atlantic World. Isis, 111(4):877–878, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Valleriani:2020:BRP

[235] Matteo Valleriani. Book review: Pietro Greco. Galileo Galilei, the Tuscan Artist. Isis, 111(4):878–879, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Silver:2020:BRM

[236] Sean Silver. Book reviews: Margaret Willes. The Curious World of Samuel Pepys and John Evelyn; John Dixon Hunt. John Evelyn: a Life of Domesticity. Isis, 111(4):879–881, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanBerkel:2020:BRA

[237] Klaas van Berkel. Book review: Andrea Strazzoni. Dutch Cartesianism and the Birth of Philosophy of Science: From Regius to ’s Gravesande. Isis, 111(4):881–882, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brentjes:2020:BRC

[238] Sonja Brentjes. Book review: Cornel Zwierlein. Imperial Unknowns: The French and British in the Mediterranean, 1650-1750. Isis, 111(4):883– 884, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Keblusek:2020:BRE

[239] Marika Keblusek. Book review: Eva Dolezel; Rainer Godel; Andreas Pecar; Holger Zaunstöck, eds. Ordnen-Vernetzen-Vermitteln: Kunst- und Naturalienkammern der Frühen Neuzeit als Lehr- und Lernorte. Isis, 111 (4):884–885, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rouphail:2020:BRJ

[240] Robert M. Rouphail. Book review: Jeremy Adelman, ed. Empire and the Social Sciences: Global Histories of Knowledge. Isis, 111(4):886–887, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bhattacharya:2020:BRS

[241] Nandini Bhattacharya. Book review: Stephen Snelders. Leprosy and Colonialism: Suriname under Dutch Rule, 1750-1950. Isis, 111(4):887–888, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ronda:2020:BRR

[242] Margaret Ronda. Book review: Ron Broglio. Beasts of Burden: Biopolitics, Labor, and Animal Life in British Romanticism. Isis, 111(4):888– 889, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hawkins:2020:BRA

[243] Spencer Hawkins. Book review: Alison E. Martin. Nature Translated: Alexander von Humboldt’s Works in Nineteenth-Century Britain. Isis, 111 (4):890–891, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Keller:2020:BRA

[244] Richard C. Keller. Book review: Amir A. Afkhami. A Modern Contagion: Imperialism and Public Health in Iran’s Age of Cholera. Isis, 111(4):891– 892, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DAntonio:2020:BRJ

[245] Patricia D’Antonio. Book review: Jessica Wang. Mad Dogs and Other New Yorkers: Rabies, Medicine, and Society in an American Metropolis, 1840–1920. Isis, 111(4):892–894, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Siddiqi:2020:BRC

[246] Asif A. Siddiqi. Book review: Catherine L. Newell. Destined for the Stars: Faith, the Future, and America’s Final Frontier. Isis, 111(4):894–895, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mishra:2020:BRM

[247] Saurabh Mishra. Book review: Marilyn Bailey Ogilvie. For the Birds: American Ornithologist Margaret Morse Nice. Isis, 111(4):895–896, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Noble:2020:BRI

[248] Brian Noble. Book review: Ilja Nieuwland. American Dinosaur Abroad: a Cultural History of Carnegie’s Plaster Diplodocus. Isis, 111(4):896–898, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kox:2020:BRJ

[249] A. J. Kox. Book review: Jaume Navarro, ed. Ether and Modernity: The Recalcitrance of an Epistemic Object in the Early Twentieth Century. Isis, 111(4):898–899, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jacobsen:2020:BRJ

[250] Anja Skaar Jacobsen. Book review: Jörg Resag. Feynman and His Physics: The Life and Science of an Extraordinary Man. (Springer Biographies.) xii + 319 pp., figs., bibl. Cham, Switzerland: Springer, 2018. EUR 32.69 (cloth); ISBN 978-3-319-96835-3. E-book available. Isis, 111(4):899–900, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rasmussen:2020:BRJ

[251] Seth C. Rasmussen. Book review: John F. Marra. Hot Carbon: Carbon-14 and a Revolution in Science. Isis, 111(4):900–901, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gioielli:2020:BRD

[252] Robert Gioielli. Book review: Dolores L. Augustine. Taking on Technocracy: Nuclear Power in Germany, 1945 to the Present. Isis, 111(4): 902–903, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lajus:2020:BRP

[253] Julia Lajus. Book review: Perrin Selcer. The Postwar Origins of the Global Environment: How the United Nations Built Spaceship Earth. Isis, 111 (4):903–904, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hepler-Smith:2020:BRJ

[254] Evan Hepler-Smith. Book review: Joshua P. Howe, ed. Making Climate Change History: Documents from Global Warming’s Past. Isis, 111(4): 904–906, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bargheer:2020:SSP

[255] Stefan Bargheer. Simone schleper. planning for the planet: environmental expertise and the international union for conservation of nature and natural resources, 1960–1980. Isis, 111(4):906–907, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wisnioski:2020:BRD

[256] Matthew Wisnioski. Book review: David Cateforis; Steven Duval; Shepherd Steiner, eds. Hybrid Practices: Art in Collaboration with Science and Technology in the Long 1960s. Isis, 111(4):907–908, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jankovic:2020:BRR

[257] Vladimir Jankovic’. Book review: Rachel Emma Rothschild. Poisonous Skies: Acid Rain and the Globalization of Pollution. Isis, 111(4):909–910, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pols:2020:BRA

[258] Hans Pols. Book review: Allan V. Horwitz. PTSD: A Short History. Isis, 111(4):910–911, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2020:CIV

[259] Anonymous. Contents of Isis, volume 111. Isis, 111(4):912–914, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2020:ICB

[260] Anonymous. Isis current bibliography of the history of science and its cultural influences 2020. Isis, 111(S1):1–317, December 2020. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2021:FCa

[261] Anonymous. Front cover. Isis, 112(1):??, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2021:FMa

[262] Anonymous. Front matter. Isis, 112(1):??, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Curry:2021:TRS

[263] Helen Anne Curry. Taxonomy, race science, and Mexican maize. Isis, 112 (1):1–21, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Prkachin:2021:SBB

[264] Yvan Prkachin. “The sleeping beauty of the brain”: Memory, MIT, Montreal, and the origins of neuroscience. Isis, 112(1):22–44, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cooley:2021:GRM

[265] Mackenzie Cooley. The giant remains: Mesoamerican natural history, medicine, and cycles of empire. Isis, 112(1):45–67, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Michel:2021:MSS

[266] Nicolas Michel. Mathematical selves and the shaping of mathematical modernism: Conflicting epistemic ideals in the emergence of enumerative geometry (1864–1893). Isis, 112(1):68–92, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

White:2021:MGS

[267] Rebekah C. White, Giles E. M. Gasper, Tom C. B. McLeish, Brian K. Tanner, Joshua S. Harvey, Sigbjørn O. Sønnesyn, Laura K. Young, and Hannah E. Smithson. Magnifying grains of sand, seeds, and blades of grass: Optical effects in Robert Grosseteste’s De iride (On the Rainbow) (circa 1228–1230). Isis, 112(1):93–107, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

GrafvonHardenberg:2021:KLP

[268] Wilko Graf von Hardenberg. Knowing the littoral: Perception and representation of terraqueous spaces in a global perspective. Isis, 112(1): 108–110, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Papelitzky:2021:SWC

[269] Elke Papelitzky. Shallow water at China’s coast: Depicting dangers on early modern Chinese maps. Isis, 112(1):111–121, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pastore:2021:SSW

[270] Christopher L. Pastore. The science of shallow waters: Connecting and classifying the early modern Atlantic. Isis, 112(1):122–129, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

GrafvonHardenberg:2021:MSS

[271] Wilko Graf von Hardenberg. Making a stable sea: The littorals of Eighteenth-Century Europe and the origins of a spatial concept. Isis, 112(1):130–140, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bhattacharyya:2021:RPS

[272] Debjani Bhattacharyya. A river is not a pendulum: Sediments of science in the world of tides. Isis, 112(1):141–149, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tolmacheva:2021:RKS

[273] Marina Tolmacheva. Reflections on the knowledge of shallow waters. Isis, 112(1):150–153, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Carrier:2021:LE

[274] Richard C. Carrier. Letter to the Editor. Isis, 112(1):154–155, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cohen:2021:SSM

[275] H. Floris Cohen. Science shaping modernity: Stephen Gaukroger’s fourvolume series completed. Isis, 112(1):156–163, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Camprubi:2021:WPH

[276] Lino Camprubí. Water and planetary history. Isis, 112(1):164–167, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

NortonGreene:2021:BRM

[277] Ann Norton Greene. Book review: Margery Fee. Polar Bear. Isis, 112(1): 168–169, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Howkins:2021:BRP

[278] Adrian Howkins. Book review: Peter G. Knight. Glacier: Nature and Culture. Isis, 112(1):169–170, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Karchmer:2021:BRH

[279] Eric I. Karchmer. Book review: Hilary A. Smith. Forgotten Disease: Illnesses Transformed in Chinese Medicine. Isis, 112(1):170–171, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jiang:2021:BRA

[280] Lijing Jiang. Book review: Anna Marie Roos. Goldfish. Isis, 112(1):171– 173, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Montgomery:2021:BRH

[281] Georgina M. Montgomery. Book reviews: Helena Ekerholm; Karl Grandin; Christer Nordlund; Patience A. Schell, eds. Understanding Field Science Institutions; Michael J. Lannoo. This Land Is Your Land: The Story of Field Biology in America. Isis, 112(1):173–174, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rocke:2021:BRG

[282] Alan Rocke. Book review: Glen E. Rodgers. Traveling with the Atom: a Scientific Guide to Europe and Beyond. Isis, 112(1):175–176, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Li:2021:BRE

[283] Lan A. Li. Book review: Eric Huntington. Creating the Universe: Depictions of the Cosmos in Himalayan Buddhism. Isis, 112(1):176–177, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Raina:2021:BRJ

[284] Dhruv Raina. Book review: Jobin M. Kanjirakkat; Gordon McOuat; Sundar Sarukkai, eds. Science and Narratives of Nature: East and West. Isis, 112(1):177–179, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hanegraaff:2021:BRE

[285] Wouter J. Hanegraaff. Book review: Esteban Law. Das Corpus Hermeticum-Wirkungsgeschichte: Transzendenz, Immanenz, Ethik. Isis, 112(1):179–180, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Beretta:2021:BRA

[286] Marco Beretta. Book review: Alison M. Roberts. Hathor’s Alchemy: The Ancient Egyptian Roots of the Hermetic Art. Isis, 112(1):??, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McHugh:2021:BRK

[287] James McHugh. Book review: Katelynn Robinson. The Sense of Smell in the Middle Ages: a Source of Certainty. Isis, 112(1):182–183, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jacquart:2021:BRH

[288] Danielle Jacquart. Book review: Herbert L. Kessler; Richard G. Newhauser; Arthur J. Russell, eds. Optics, Ethics, and Art in the Thirteenth and Fourteenth Centuries: Looking into Peter of Limoge’s “Moral Treatise on the Eye”. Isis, 112(1):183–184, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rosen:2021:BRP

[289] Mark Rosen. Book review: Paolo Galluzzi; trans. by Jonathan Mandelbaum. The Italian Renaissance of Machines. Isis, 112(1):184–185, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sepkoski:2021:BRL

[290] David Sepkoski. Book review: Lydia Barnett. After the Flood: Imagining the Global Environment in Early Modern Europe. Isis, 112(1):186–187, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Canizares-Esguerra:2021:BRJ

[291] Jorge Cañizares-Esguerra. Book review: Jaime Marroquín Arredondo; Ralph Bauer, eds. Translating Nature: Cross-Cultural Histories of Early Modern Science. Isis, 112(1):187–188, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Paradis:2021:BRH

[292] Swann Paradis. Book review: Hanna Roman. The Language of Nature in Buffon’s Histoire naturelle. Isis, 112(1):189–190, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Burkett:2021:BRJ

[293] Andrew Burkett. Book review: Joel Faflak, ed. Marking Time: Romanticism and Evolution. Isis, 112(1):190–192, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barton:2021:BRR

[294] Patricia Barton. Book review: Rohan Deb Roy. Malarial Subjects: Empire, Medicine, and Nonhumans in British India, 1820–1909. Isis, 112(1): 192–193, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dunk:2021:BRP

[295] James Dunk. Book review: Peter H. Hoffenberg. A Science of Our Own: Exhibitions and the Rise of Australian Public Science. Isis, 112(1):193– 194, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heggie:2021:BRG

[296] Vanessa Heggie. Book review: Gillen D’Arcy Wood. Land of Wondrous Cold: The Race to Discover Antarctica and Unlock the Secrets of Its Ice. Isis, 112(1):194–195, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anker:2021:BRM

[297] Peder Anker. Book review: Marilyn Fischer. Jane Addams’s Evolutionary Theorizing: Constructing “Democracy and Social Ethics”. Isis, 112(1): 195–196, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Goderle:2021:BRJ

[298] Wolfgang Göderle. Book review: Jan Surman. Universities in Imperial Austria, 1848–1918: a Social History of a Multilingual Space. Isis, 112 (1):197–198, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Johnson:2021:BRV

[299] Andi Johnson. Book review: Vanessa Heggie. Higher and Colder: a History of Extreme Physiology and Exploration. Isis, 112(1):198–199, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wilson:2021:BRJ

[300] Daniel J. Wilson. Book review: Jan Peter Verhave. A Constant State of Emergency: Paul de Kruif: Microbe Hunter and Health Activist. Isis, 112(1):200–201, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kapila:2021:BRD

[301] Shruti Kapila. Book review: Daniela Finzi; Herman Westerink, eds. Dora, Hysteria, and Gender: Reconsidering Freud’s Case Study. Isis, 112(1): 201–202, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Creager:2021:BRM

[302] Angela N. H. Creager. Book review: M. Norton Wise. Aesthetics, Industry, and Science: Hermann von Helmholtz and the Berlin Physical Society. Isis, 112(1):202–203, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wise:2021:BRM

[303] M. Norton Wise. Book review: Michael D. Gordin. Einstein in Bohemia. Isis, 112(1):203–204, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Adams:2021:BRP

[304] Mark B. Adams. Book review: Peter J. Bowler. A History of the Future: Prophets of Progress from H. G. Wells to Isaac Asimov. Isis, 112(1):205– 206, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Conklin:2021:BRP

[305] Alice L. Conklin. Book review: Paul-André Rosental. A Human Garden: French Policy and the Transatlantic Legacies of Eugenic Experimentation. Isis, 112(1):206–207, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Campbell:2021:BRN

[306] Nancy D. Campbell. Book review: Nadine Ehlers; Shiloh Krupar. Deadly Biocultures: The Ethics of Life-Making. Isis, 112(1):208–209, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hersch:2021:BRR

[307] Matthew Hersch. Book review: Robert Serber; introduction by Richard Rhodes. The Los Alamos Primer: The First Lectures on How to Build an Atomic Bomb. Isis, 112(1):209–210, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Raby:2021:BRS

[308] Megan Raby. Book review: Sabine Clarke. Science at the End of Empire: Experts and the Development of the British Caribbean, 1940–62. Isis, 112(1):210–211, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Whitesides:2021:BRG

[309] John Gregory Whitesides. Book review: Gerson S. Sher. From Pugwash to Putin: a Critical History of U.S.-Soviet Scientific Cooperation. Isis, 112(1):212–213, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Habinek:2021:BRM

[310] Lianne Habinek, Hannah Murphy, By Harold J. Cook, Projit Bihari Mukharji, Myrna Perez Sheldon, Elise K. Burton, Sebastián Gil-Ria no, Terence Keel, Emily Merchant, Wangui Muigai, Ahmed Ragab, Suman Seth, Robert K. Merton, James A. Secord, and Kapil Raj. Book review: Mark Dennis Robinson. The Market in Mind: How Financialization Is Shaping Neuroscience, Translational Medicine, and Innovation in Biotechnology. Isis, 112(1):213–214, March 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2021:FCb

[311] Anonymous. Front cover. Isis, 112(2):??, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2021:FMb

[312] Anonymous. Front matter. Isis, 112(2):??, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bijsterveld:2021:SSS

[313] Karin Bijsterveld. Slicing sound: Speaker identification and sonic skills at the Stasi, 1966-1989. Isis, 112(2):215–241, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Levitin:2021:NRP

[314] Dmitri Levitin. Newton on the rules of philosophizing and hypotheses: New evidence, new conclusions. Isis, 112(2):242–265, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Polat:2021:MHP

[315] Bican Polat. Mental hygiene, psychoanalysis, and interwar psychology: The making of the maternal deprivation hypothesis. Isis, 112(2):266– 290, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sattelmacher:2021:IRR

[316] Anja Sattelmacher, Mario Schulze, and Sarine Waltenspül. Introduction: Reusing research film and the Institute for Scientific Film. Isis, 112(2): 291–298, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Reiss:2021:SCE

[317] Christian Reiß. Shooting chicken embryos: The making of Ludwig Gräper’s embryological films, 1911–1940. Isis, 112(2):299–306, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Olszynko-Gryn:2021:FFE

[318] Jesse Olszynko-Gryn. Filming fly eggs: Time-lapse cinematography as an intermedial practice. Isis, 112(2):307–314, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Leyssen:2021:RMR

[319] Sigrid Leyssen. Remaking “Michotte”: Reusing and remaking moving images in the history of perception research. Isis, 112(2):315–325, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Scholz:2021:DNE

[320] Juliane Scholz. Duplicating nature and elements of subjectivity in The Ethology of the Greylag Goose. Isis, 112(2):326–334, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Naser:2021:BAR

[321] Torsten Näser. Between aspiration and reality: Folklore film work in times of upheaval. Isis, 112(2):335–341, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Waltenspul:2021:RNF

[322] Sarine Waltenspül. Reusable and nonreusable films: From ballistic films to the encyclopaedia cinematographica. Isis, 112(2):342–351, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sattelmacher:2021:SZC

[323] Anja Sattelmacher. Shuffled Zeppelin clips: The flight and crash of LZ 129 Hindenburg in the archives. Isis, 112(2):352–360, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schulze:2021:MMI

[324] Mario Schulze. Mobilizing moving images: Reusing a German flow film of the 1920s for U.S. science education in the Cold War. Isis, 112(2): 361–369, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gaycken:2021:FEU

[325] Oliver Gaycken. Film ist. (1–6): an exemplary use of the encyclopaedia cinematographica. Isis, 112(2):370–378, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lindqvist:2021:ENW

[326] Svante Lindqvist. Eloge: Neale W. Watson (1934–2019). Isis, 112(2): 379–381, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ungureanu:2021:CHS

[327] James C. Ungureanu. The complicated history of science and religion. Isis, 112(2):382–388, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Naramore:2021:BRA

[328] Sarah Naramore. Book review: Amy Koerber. From Hysteria to Hormones: a Rhetorical History. Isis, 112(2):389–390, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DeBont:2021:BRJ

[329] Raf De Bont. Book review: Jürgen Renn. The Evolution of Knowledge: Rethinking Science for the Anthropocene. Isis, 112(2):390–391, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stoffel:2021:BRM

[330] Jean-François Stoffel. Book review: Michel Blay. Critique de l’histoire des sciences. Isis, 112(2):391–392, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Suarez-Diaz:2021:BRM

[331] Edna Suárez-Díaz. Book review: M. Susan Lindee. Rational Fog: Science and Technology in Modern War. Isis, 112(2):393–394, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Matala:2021:LDF

[332] Saara Matala. Larrie d. ferreiro. Bridging the Seas: The Rise of Naval Architecture in the Industrial Age, 1800-2000.. Isis, 112(2):394–395, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Flow:2021:BRK

[333] Christian Flow. Book review: Kassandra J. Miller; Sarah L. Symons, eds. Down to the Hour: Short Time in the Ancient Mediterranean and Near East. Isis, 112(2):395–397, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Carson:2021:ALM

[334] Sarah Carson. Anne Lawrence-mathers. Medieval Meteorology: Forecasting the Weather from Aristotle to the Almanac. Isis, 112(2):397–398, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Galle:2021:BRK

[335] Karl Galle. Book review: Katharina N. Piechocki. Cartographic Humanism: The Making of Early Modern Europe. Isis, 112(2):399–400, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Reeves:2021:BRW

[336] Eileen Reeves. Book review: William R. Shea. Conversations with Galileo. Isis, 112(2):400–401, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rankin:2021:BRB

[337] Alisha Rankin. Book review: Benjamin Breen. The Age of Intoxication: Origins of the Global Drug Trade. Isis, 112(2):401–403, June 2021. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roos:2021:BRM

[338] Anna Marie Roos. Book review: Mordechai Feingold; Giulia Giannini, eds. The Institutionalization of Science in Early Modern Europe. Isis, 112(2):403–404, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Maass:2021:BRS

[339] Yvonne Maaß. Book review: Silke Förschler; Anne Mariss, eds. Akteure, Tiere, Dinge: Verfahrensweisen der Naturgeschichte in der Frühen Neuzeit. Isis, 112(2):404–406, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Maass:2021:BRB

[340] Yvonne Maaß. Book review: Bettina Dietz. Das System der Natur: Die kollaborative Wissenskultur der Botanik im 18. Jahrhundert. Isis, 112(2): 406–407, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sysling:2021:BRJ

[341] Fenneke Sysling. Book review: James Poskett. Materials of the Mind: Phrenology, Race, and the Global History of Science, 1815-1920.. Isis, 112(2):407–408, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

OSullivan:2021:BRJ

[342] Tanya O’Sullivan. Book review: James C. Ungureanu. Science, Religion, and the Protestant Tradition: Retracing the Origins of Conflict. Isis, 112 (2):409–410, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vanderstraeten:2021:BRJ

[343] Raf Vanderstraeten. Book review: Joris Vandendriessche. Medical Societies and Scientific Culture in Nineteenth-Century Belgium. Isis, 112(2): 410–411, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rieppel:2021:BRJ

[344] Lukas Rieppel. Book review: Jennifer L. Derr. The Lived Nile: Environment, Disease, and Material Colonial Economy in Egypt. Isis, 112(2): 411–412, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brain:2021:BRM

[345] Robert Brain. Book review: Michael Rossi. The Republic of Color: Science, Perception, and the Making of Modern America. Isis, 112(2):412– 414, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lim:2021:BRT

[346] Samson Lim. Book review: Trais Pearson. Sovereign Necropolis: The Politics of Death in Semi-Colonial Siam. Isis, 112(2):414–415, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cowles:2021:BRL

[347] Henry M. Cowles. Book review: Lukas Rieppel. Assembling the Dinosaur: Fossil Hunters, Tycoons, and the Making of a Spectacle. Isis, 112(2):415– 417, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hochadel:2021:BRT

[348] Oliver Hochadel. Book review: Tanya O’Sullivan. Geographies of City Science: Urban Lives and Origin Debates in Late Victorian Dublin. Isis, 112(2):417–418, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lindee:2021:BRJ

[349] Susan Lindee. Book review: Jeffrey Womack. Radiation Evangelists: Technology, Therapy, and Uncertainty at the Turn of the Century. Isis, 112(2):418–419, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Donato:2021:BRF

[350] Maria Pia Donato. Book review: Francisco Malta Romeiras. Jesuits and the Book of Nature: Science and Education in Modern Portugal. Isis, 112(2):419–420, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wu:2021:BRB

[351] Shellen Xiao Wu. Book review: On Barak. Powering Empire: How Coal Made the Middle East and Sparked Global Carbonization. Isis, 112(2): 421–422, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Guerra:2021:BRE

[352] Francesco Guerra. Book review: Erasmo Recami. The Majorana Case: Letters, Documents, Testimonies, 400 pp., bibl., app. Singapore: World Scientific, 2020. £40 (paper); ISBN 978-981-12-0701-3. Isis, 112(2):422– 423, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Meister:2021:BRZ

[353] Anna-Maria Meister. Book review: Zeynep Çelik Alexander; John May, eds. Design Technics: Archaeologies of Architectural Practice. Isis, 112(2): 423–425, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rappaport:2021:BRS

[354] Erika Rappaport. Book review: Sarah Besky. Tasting Qualities: The Past and Future of Tea. Isis, 112(2):425–426, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hsiung:2021:BRM

[355] Hansun Hsiung. Book review: Miriam Kingsberg Kadia. Into the Field: Human Scientists of Transwar Japan. Isis, 112(2):426–428, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wang:2021:BRA

[356] Zuoyue Wang. Book review: Arunabh Ghosh. Making It Count: Statistics and Statecraft in the Early People’s Republic of China. Isis, 112(2):428– 429, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shapira:2021:BRE

[357] Michal Shapira. Book review: Erika Lorraine Milam. Creatures of Cain: The Hunt for Human Nature in Cold War America. Isis, 112(2):429– 430, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Helbig:2021:BRK

[358] Phillip Helbig. Book review: Klaus Mauersberger; Monika Schulz-Fieguth, eds. Hans-Jürgen Treder: Ein Porträt. Isis, 112(2):431–432, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Helbig:2021:BRM

[359] Phillip Helbig. Book review: Michael Kaasch; Joachim Kaasch; Torsten K. D. Himmel, eds. Biologie in der DDR: Beiträge zur 24. Jahrestagung der DGGTB in Greifswald 2015.. Isis, 112(2):432–433, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Raby:2021:BRA

[360] Megan Raby. Book review: Andra B. Chastain; Timothy Lorek, eds. Itineraries of Expertise: Science, Technology, and the Environment in Latin America’s Long Cold War. Isis, 112(2):433–434, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lubar:2021:BRE

[361] Steven Lubar. Book review: Elena Canadelli; Marco Beretta; Laura Ronzon, eds. Behind the Exhibit: Displaying Science and Technology at World’s Fairs and Museums in the Twentieth Century. Isis, 112(2):??, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Krige:2021:BRA

[362] John Krige. Book review: Abena Dove Osseo-Asare. Atomic Junction: Nuclear Power in Africa after Independence. Isis, 112(2):436–437, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rogers:2021:BRS

[363] Molly Rogers, Karin Bijsterveld, Anja Sattelmacher, Mario Schulze, Sarine Waltenspül, Jesse Olszynko-Gryn, Svante Lindqvist, James C. Ungureanu, and Jean-François Stoffel. Book review: Shawn Michelle Smith. Photographic Returns: Racial Justice and the Time of Photography. Isis, 112(2):437–438, June 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2021:FCc

[364] Anonymous. Front cover. Isis, 112(3):??, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2021:FMc

[365] Anonymous. Front matter. Isis, 112(3):??, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Engelmann:2021:ESO

[366] Lukas Engelmann. An epidemic for sale: Observation, modification, and commercial circulation of the Danysz virus, 1890–1910. Isis, 112(3):439– 460, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Burton:2021:RCM

[367] Elise K. Burton. Rethinking collaboration: Medical research and working relationships at the Iranian Pasteur Institute. Isis, 112(3):461–483, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rubinson:2021:MSA

[368] Paul Rubinson. Mistress of the sciences, asylum of Liberty: Joseph Priestley, human rights, and science in the early U.S. Republic. Isis, 112(3):484– 506, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Keller:2021:WSP

[369] Vera Keller. A “wild swing to phantsy”: The philosophical gardener and emergent experimental philosophy in the Seventeenth-Century Atlantic world. Isis, 112(3):507–530, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Soderfeldt:2021:TWM

[370] Ylva Söderfeldt. The truth within: Making medical knowledge in the Hay Fever Association of Heligoland, 1899–1909. Isis, 112(3):531–547, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bray:2021:UIT

[371] Francesca Bray. Underground inspirations: Tuber sciences and their histories. Isis, 112(3):548–563, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2021:HSS

[372] Anonymous. History of science society virtual forum, 2020. Isis, 112 (3):564–572, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marcos:2021:HVF

[373] Patrícia Martins Marcos, Sarah E. Naramore, Myrna Perez Sheldon, Sarah Pickman, Sarah A. Qidwai, and Kathleen Sheppard. HSS virtual forum: Futures series. Isis, 112(3):573–581, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marcus:2021:EHP

[374] Alan I. Marcus. Eloge: Howard P. Segal (1948–2020). Isis, 112(3):582– 583, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ducheyne:2021:LE

[375] Steffen Ducheyne. Letter to the Editor. Isis, 112(3):??, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Henry:2021:LE

[376] John Henry. Letter to the Editor. Isis, 112(3):584–585, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Li:2021:CDH

[377] Lan A. Li. Crafting digital histories of science: a review and tour of Secrets of Craft and Nature in Renaissance France. Isis, 112(3):586–589, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Topp:2021:BRJ

[378] Leslie Topp. Book review: Jane L. Stevens Crawshaw; Irena Benyovsky Latin; Kathleen Vongsathorn, eds. Tracing Hospital Boundaries: Integration and Segregation in Southeastern Europe and Beyond, 1050–1970. Isis, 112(3):590–591, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roth:2021:BRL

[379] Cassia Roth. Book review: Lara Freidenfelds. The Myth of the Perfect Pregnancy: a History of Miscarriage in America. Isis, 112(3):591–592, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Linker:2021:BRJ

[380] Beth Linker. Book review: Jaipreet Virdi. Hearing Happiness: Deafness Cures in History. Isis, 112(3):592–593, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Whitney:2021:BRS

[381] Kristoffer Whitney. Book review: Sebastian Vehlken. Zootechnologies: a Media History of Swarm Research. Isis, 112(3):594–595, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schaefer:2021:BRA

[382] Donovan O. Schaefer. Book review: Andrew Jewett. Science under Fire: Challenges to Scientific Authority in Modern America. Isis, 112(3):595– 596, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jones:2021:BRH

[383] Alexander Jones. Book review: Hermann Hunger; John Steele. The Babylonian Astronomical Compendium MUL.APIN. Isis, 112(3):597–598, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Polloni:2021:BRD

[384] Nicola Polloni. Book review: Dag Nikolaus Hasse; Amos Bertolacci, eds. The Arabic, Hebrew and Latin Reception of Avicenna’s Physics and Cosmology. Isis, 112(3):598–599, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gamini:2021:BRH

[385] Amir-Mohammad Gamini. Book review: Hanif Ghalandari. ’Abd al-Jabbār al-Kharaqī’s Muntaha al-Idrāk fi Taqāsīm al-Aflāk (Ultimate Comprehension of the Subdivision of Celestial Spheres: The First Comprehensive hay’a Work on Ptolemaic Cosmology). Isis, 112(3):599–600, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kedar:2021:BRN

[386] Yael Kedar. Book review: Nikolaus Egel, ed. Roger Bacon, Opus Tertium. Isis, 112(3):600–602, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mezes:2021:BRF

[387] Ádám Mézes. Book review: Francesco Paolo de Ceglia. The Body of Evidence: Corpses and Proofs in Early Modern European Medicine. Isis, 112 (3):602–603, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Robison:2021:BRC

[388] Kira Robison. Book review: Christian K. Kleinbub. Michelangelo’s Inner Anatomies. Isis, 112(3):603–604, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hahn:2021:BRF

[389] Barbara Hahn. Book review: Frances E. Dolan. Digging the Past: How and Why to Imagine Seventeenth-Century Agriculture. Isis, 112(3):605– 606, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Atcil:2021:BRH

[390] Abdurrahman Atçil. Book review: Harun Küçük. Science without Leisure: Practical Naturalism in Istanbul, 1660-1732.. Isis, 112(3):606– 607, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Reeves:2021:BRP

[391] Eileen Reeves. Book review: Per Pippin Aspaas; László Kontler. Maximilian Hell (1720-92) and the Ends of Jesuit Science in Enlightenment Europe. Isis, 112(3):607–609, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Caulkins:2021:BRB

[392] Tamara Caulkins. Book review: Bruno Belhoste; trans. by Susan Emanuel. Paris Savant: Capital of Science in the Age of Enlightenment. Isis, 112(3):609–610, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dysart:2021:BRJ

[393] Taylor Dysart. Book review: Juan Carlos González Espitia. Sifilografía: A History of the Writerly Pox in the Eighteenth-Century Hispanic World. Isis, 112(3):610–611, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pearn:2021:BRC

[394] Alison Pearn. Book review: Charles H. Smith; James T. Costa; David A. Collard, eds. An Alfred Russel Wallace Companion. Isis, 112(3):612–613, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Braun:2021:BRA

[395] Lundy Braun. Book review: Arleen Marcia Tuchman. Diabetes: a History of Race and Disease. Isis, 112(3):613–614, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fairman:2021:BRR

[396] Julie A. Fairman. Book review: Ronald S. Coddington. Faces of Civil War Nurses. Isis, 112(3):614–616, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Grafe:2021:BRC

[397] Sophia Gräfe. Book review: Christian Reiß. Der Axolotl: Ein Labortier im Heimaquarium 1864-1914.. Isis, 112(3):616–617, September 2021. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Siddiqi:2021:BRN

[398] Asif Siddiqi. Book review: Nikolai Krementsov. With and Without Galton: Vasilii Florinskii and the Fate of Eugenics in Russia. Isis, 112(3):617–619, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hahn:2021:BRP

[399] Andre Michael Hahn. Book review: Peter Ayres. Women and the Natural Sciences in Edwardian Britain: In Search of Fellowship. Isis, 112(3):619– 620, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Madruga:2021:BRU

[400] Catarina Madruga. Book review: Ulrike Kirchberger; Brett M. Bennett, eds. Environments of Empire: Networks and Agents of Ecological Change. Isis, 112(3):620–621, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hentschel:2021:BRJ

[401] Klaus Hentschel. Book review: Jost Lemmerich. Max von Laue-Furchtlos und treu: Eine Biographie des Nobelpreisträgers für Physik. Isis, 112 (3):622–623, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Meyer-Spasche:2021:BRA

[402] Rita Meyer-Spasche. Book review: Arne Schirrmacher. Establishing Quantum Physics in Göttingen: David Hilbert, Max Born, and Peter Debye in Context, 1900-1926.. Isis, 112(3):623–624, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Seow:2021:BRE

[403] Victor Seow. Book review: Eugenia Lean. Vernacular Industrialism in China: Local Innovation and Translated Technologies in the Making of a Cosmetics Empire, 1900-1940.. Isis, 112(3):624–626, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shelangoskie:2021:BRP

[404] Susan Shelangoskie. Book review: Paul J. Nahin. Hot Molecules, Cold Electrons: From the Mathematics of Heat to the Development of the Trans-Atlantic Cable. Isis, 112(3):626–627, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Merchant:2021:BRE

[405] Emily Klancher Merchant. Book review: Emmanuel Didier; trans. by Priya Vari Sen. America by the Numbers: Quantification, Democracy, and the Birth of National Statistics. Isis, 112(3):627–628, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wolffram:2021:BRR

[406] Heather Wolffram. Book review: Richard Noakes. Physics and Psychics: The Occult and the Sciences in Modern Britain. Isis, 112(3):628–629, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pickman:2021:BRE

[407] Sarah M. Pickman. Book review: Elizabeth Baigent; André Reyes Novaes, eds. Geographers: Biobibliographical Studies. Isis, 112(3):630–631, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Monty:2021:BRT

[408] Adrian Monty and Linda Marie Richards. Book review: Trisha T. Pritikin; introduction by Karen Dorn Steele. The Hanford Plaintiffs: Voices from the Fight for Atomic Justice. Isis, 112(3):631–632, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Maher:2021:BRB

[409] Neil M. Maher. Book review: Brian C. Odom; Stephen P. Waring, eds. NASA and the Long Civil Rights Movement. Isis, 112(3):632–634, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bell:2021:BRA

[410] Sarah A. Bell. Book review: Antonio Badia. The Information Manifold: Why Computers Can’t Solve Algorithmic Bias and Fake News. Isis, 112 (3):634–635, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roberts:2021:BRP

[411] Peder Roberts. Book review: Peder Anker. The Power of the Periphery: How Norway Became an Environmental Pioneer for the World. Isis, 112 (3):635–636, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hewitt:2021:BRD

[412] Kim Hewitt. Book review: Danielle Giffort. Acid Revival: The Psychedelic Renaissance and the Quest for Medical Legitimacy. Isis, 112(3):637–638, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mathew:2021:BRP

[413] John Mathew. Book review: Pankaj Sekhsaria. Instrumental Lives: an Intimate Biography of an Indian Laboratory. Isis, 112(3):638–639, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hepler-Smith:2021:BRI

[414] Evan Hepler-Smith. Book review: Isabel Malaquias; Peter J. T. Morris, eds. Perspectives on Chemical Biography in the Twenty-First Century. Isis, 112(3):639–640, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mitra:2021:BRD

[415] Durba Mitra. Book review: Dwaipayan Banerjee. Enduring Cancer: Life, Death, and Diagnosis in Delhi. Isis, 112(3):641–642, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Doyle:2021:BRJ

[416] Richard Doyle, Lukas Engelmann, Vera Keller, Asif Siddiqi, Klaus Hentschel, and John Henry. Book review: Jennifer Johung. Vital Forms: Biological Art, Architecture, and the Dependencies of Life. Isis, 112(3): 642–643, September 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2021:FCd

[417] Anonymous. Front cover. Isis, 112(4):??, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2021:FMd

[418] Anonymous. Front matter. Isis, 112(4):??, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hegesh:2021:MGA

[419] Noa Hegesh. Mind the gap: Acoustical answers to cosmological concerns in First-Century B.C.E. China. Isis, 112(4):645–669, December 2021. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mogilner:2021:MEP

[420] Marina Mogilner. Mediterraneanizing Europe: The project of subaltern race and the postimperial search for hybridity. Isis, 112(4):670–693, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bolman:2021:PMR

[421] Brad Bolman. Pig mentations: Race and face in radiobiology. Isis, 112 (4):694–716, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Verburgt:2021:WFB

[422] Lukas M. Verburgt. The Works of Francis Bacon: A Victorian classic in the history of science. Isis, 112(4):717–736, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Duppe:2021:HWS

[423] Till Düppe. How Western science corrupts class consciousness: East Germany’s presence at IIASA. Isis, 112(4):737–759, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hsiung:2021:IEM

[424] Hansun Hsiung and Elena Serrano. Introduction: Epistemologies of the match. Isis, 112(4):760–765, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Azzolini:2021:SAM

[425] Monica Azzolini. Are the stars aligned? Matchmaking and astrology in Early Modern Italy. Isis, 112(4):766–775, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Serrano:2021:FPB

[426] Elena Serrano. A feminist physiology: B. J. Feijoo (1676–1764) and his advice for those in love. Isis, 112(4):776–785, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hsiung:2021:HEC

[427] Hansun Hsiung. From harmony to eHarmony: Charles Fourier, social science, and the management of love. Isis, 112(4):786–794, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bittel:2021:CCP

[428] Carla Bittel. Cranial compatibility: Phrenology, measurement, and marriage assessment. Isis, 112(4):795–803, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bouk:2021:LPK

[429] Dan Bouk. Love is a problem of knowledge. Isis, 112(4):804–806, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Liu:2021:TBN

[430] Yan Liu. Traversing boundaries: New digital project to study drugs in Asian history. Isis, 112(4):807–810, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hisey:2021:OAR

[431] Aimee Dávila Hisey. Open-access resource for instructors and students of history of science in Latin America and the Caribbean. Isis, 112(4):811– 814, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Naramore:2021:BRS

[432] Sarah E. Naramore. Book review: Sean Silver, The Mind Is a Collection: a Born-digital museum of eighteenth-century thought. Isis, 112(4):815– 818, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Zakar:2021:BRK

[433] Adrien Zakar. Book review: Kären Wigen; Caroline Winterer, eds. Time in Maps: From the Age of Discovery to Our Digital Era. Isis, 112(4):819– 820, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cook:2021:BRF

[434] Harold J. Cook. Book review: Fokko Jan Dijksterhuis; Andreas Weber; Huib J. Zuidervaart, eds. Locations of Knowledge in Dutch Contexts. Isis, 112(4):820–822, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Callison:2021:BRE

[435] Candis Callison. Book review: Etienne S. Benson. Surroundings: a History of Environments and Environmentalisms. Isis, 112(4):822–823, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Merrills:2021:BRF

[436] Andy Merrills. Book review: Felipe Rojas. The Pasts of Roman Anatolia: Interpreters, Traces, Horizons. Isis, 112(4):823–824, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Khalek:2021:BRA

[437] Nancy Khalek. Book review: Ahmed Ragab. Piety and Patienthood in Medieval Islam. Isis, 112(4):824–826, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heaney:2021:BRS

[438] Christopher Heaney. Book review: Steven R. Gullberg. Astronomy of the Inca Empire: Use and Significance of the Sun and Night Sky. Isis, 112 (4):826–827, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Piorko:2021:BRB

[439] Megan Piorko. Book review: Bruce T. Moran. Paracelsus: an Alchemical Life. Isis, 112(4):827–828, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mellyn:2021:BRJ

[440] Elizabeth W. Mellyn. Book review: John Christopoulos. Abortion in Early Modern Italy. Isis, 112(4):829–830, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hanson:2021:BRB

[441] Marta Hanson. Book review: He Bian. Know Your Remedies: Pharmacy and Culture in Early Modern China. Isis, 112(4):830–831, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Inglehart:2021:BRM

[442] Ashley J. Inglehart. Book review: Marina Paola Banchetti-Robino. The Chemical Philosophy of Robert Boyle: Mechanism, Chymical Atoms, and Emergence. Isis, 112(4):831–832, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schneider:2021:BRF

[443] Jeremy Robin Schneider. Book review: Francesco G. Sacco. Real, Mechanical, Experimental: Robert Hooke’s Natural Philosophy. Isis, 112(4): 833–834, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Berger:2021:BRA

[444] Susanna Berger. Book review: Anna Marie Roos. Martin Lister and His Remarkable Daughters: The Art of Science in the Seventeenth Century. Isis, 112(4):834–835, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hayton:2021:BRT

[445] Darin Hayton. Book review: Timothy Paul Grove. Christ Came Forth from India: Georgian Astrological Texts of the Seventeenth, Eighteenth, and Nineteenth Centuries. Isis, 112(4):835–836, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Podgorny:2021:BRS

[446] Irina Podgorny. Book review: Sophie Brockmann. The Science of Useful Nature in Central America: Landscapes, Networks, and Practical Enlightenment, 1784-1838.. Isis, 112(4):836–838, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lehman:2021:BRY

[447] Christine Lehman. Book review: Yann Piot. Jean-Antoine Nollet, artisan expérimentateur: Un discours technique au Xviiie siècle.. Isis, 112(4):838– 839, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Issa:2021:BRA

[448] Rana Issa. Book review: Andreas Feldtkeller; Uta Zeuge-Buberl. Networks of Knowledge: Epistemic Entanglement Initiated by American Protestant Missionary Presence in Nineteenth-Century Syria. Isis, 112(4):839–841, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jin:2021:BRC

[449] Xiaoxing Jin. Book review: Curtis N. Johnson. Darwin’s “Historical Sketch”: an Examination of the “Preface” to the Origin of Species. xxx + 472 pp., bibl., index. Oxford: Oxford University Press, 2019. £22.99 (cloth); ISBN 978-0-19-088293-8. Isis, 112(4):841–842, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shih:2021:BRR

[450] Ashanti Shih. Book review: Robert J. Mayhew; Charles W. J. Withers, eds. Geographies of Knowledge: Science, Scale, and Spatiality in the Nineteenth Century. Isis, 112(4):842–843, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Welk-Joerger:2021:BRA

[451] Nicole Welk-Joerger. Book review: Antoine Traisnel. Capture: American Pursuits and the Making of a New Animal Condition. Isis, 112(4):844– 845, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hung:2021:BRK

[452] Kuang-Chi Hung. Book review: Ken Thompson. Darwin’s Most Wonderful Plants: a Tour of His Botanical Legacy. Isis, 112(4):845–846, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roque:2021:BRA

[453] Ricardo Roque. Book review: Ashley Elizabeth Kerr. Sex, Skulls, and Citizens: Gender and Racial Science in Argentina (1860-1910).. Isis, 112 (4):847–848, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Holterhoff:2021:BRD

[454] Kate Holterhoff. Book review: David Sepkoski. Catastrophic Thinking: Extinction and the Value of Diversity from Darwin to the Anthropocene. Isis, 112(4):848–849, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McGuire:2021:BRV

[455] Coreen McGuire. Book review: Viktoria Tkaczyk; Mara Mills; Alexandra Hui, eds. Testing Hearing: The Making of Modern Aurality. Isis, 112 (4):849–851, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Monaldi:2021:BRD

[456] Daniela Monaldi. Book review: Dean Rickles. Covered with Deep Mist: The Development of Quantum Gravity (1916-1956).. Isis, 112(4):851– 852, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Newell:2021:BRM

[457] Catherine L. Newell. Book review: Matthew Shindell. The Life and Science of Harold C. Urey. Isis, 112(4):852–853, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Carson:2021:BRJ

[458] Sarah Carson. Book review: James Rodger Fleming. First Woman: Joanne Simpson and the Tropical Atmosphere. Isis, 112(4):853–855, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McDonald:2021:BRK

[459] Garret J. McDonald. Book review: Karena Kalmbach. The Meanings of a Disaster: Chernobyl and Its Afterlives in Britain and France. Isis, 112 (4):855–856, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Azzolini:2021:CIV

[460] Monica Azzolini, Noa Hegesh, and Dan Bouk. Contents of Isis, volume 112. Isis, 112(4):857–859, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2021:ICB

[461] Anonymous. Isis current bibliography of the history of science and its cultural influences 2021. Isis, 112(S1):1–300, December 2021. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2022:FCa

[462] Anonymous. Front cover. Isis, 113(1):??, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2022:FMa

[463] Anonymous. Front matter. Isis, 113(1):??, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kang:2022:CNP

[464] Hyeok Hweon Kang. Cooking niter, prototyping nature: Saltpeter and artisanal experiment in Korea, 1592–1635. Isis, 113(1):1–21, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brock:2022:SSR

[465] Claire Brock. Surgery, success, and the role of the patient in cleft palate operations, circa 1800–1930. Isis, 113(1):22–44, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kraemer:2022:BSH

[466] Fabian Kraemer, Kärin Nickelsen, and Dana von Suffrin. Botany and the science of history: Nature, culture, and the origins of civilization, circa 1850–1900. Isis, 113(1):45–62, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bos:2022:JAA

[467] Colin Bos. John Augustus Abayomi Cole and the search for an African science, 1885–1898. Isis, 113(1):63–84, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lorenat:2022:OHN

[468] Jemma Lorenat. An okapi hypothesis: Non-Euclidean geometry and the professional expert in American mathematics. Isis, 113(1):85–107, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brinitzer:2022:IWF

[469] Cameron Brinitzer and Etienne Benson. Introduction: What is a field? Transformations in fields, fieldwork, and field sciences since the Mid-Twentieth Century. Isis, 113(1):108–113, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Benson:2022:PHF

[470] Etienne Benson. The post-heroic field. Isis, 113(1):114–120, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Milam:2022:MPF

[471] Erika Lorraine Milam. Making place in the field. Isis, 113(1):121–127, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stark:2022:R

[472] Laura Stark. Reservations. Isis, 113(1):128–136, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dent:2022:WHF

[473] Rosanna Dent. Whose home is the field? Isis, 113(1):137–143, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Brinitzer:2022:GF

[474] Cameron Brinitzer. Generating fields. Isis, 113(1):144–150, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Helmreich:2022:FF

[475] Stefan Helmreich. Flipping the field. Isis, 113(1):151–156, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Greene:2022:HRF

[476] Mott T. Greene and George Gale. Henry R. Frankel (1944–2019). Isis, 113(1):157–159, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Macrakis:2022:E

[477] Kristie Macrakis. To the Editors. Isis, 113(1):??, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bijsterveld:2022:R

[478] Karin Bijsterveld. In reply. Isis, 113(1):??, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Giovannetti-Singh:2022:WHC

[479] Gianamar Giovannetti-Singh and Dániel Margócsy. Werner Herzog; Clive Oppenheimer (directors). Fireball: Visitors from Darker Worlds. Isis, 113(1):162–164, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Freidenfelds:2022:BRN

[480] Lara Freidenfelds. Book review: Nick Hopwood; Rebecca Flemming; Lauren Kassell, eds. Reproduction: Antiquity to the Present Day. Isis, 113 (1):165–166, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Messac:2022:BRR

[481] Luke Messac. Book review: Rebekah Lee. Health, Healing, and Illness in African History. Isis, 113(1):166–167, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schirvar:2022:BRS

[482] Sam Schirvar. Book review: Stephen Cave; Kanta Dihal; Sarah Dillon, eds. AI Narratives: a History of Imaginative Thinking about Intelligent Machines. Isis, 113(1):167–169, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DeLue:2022:BRJ

[483] Rachael Z. DeLue. Book review: A. Joan Saab. Objects of Vision: Making Sense of What We See. Isis, 113(1):169–170, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cooley:2022:BRR

[484] Mackenzie Cooley. Book review: Rocío G. Sumillera; Jan Surman; Katharina Kühn, eds. Translation in Knowledge, Knowledge in Translation. Isis, 113(1):170–172, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sepkoski:2022:BRD

[485] David Sepkoski. Book review: Dipesh Chakrabarty. The Climate of History in a Planetary Age.; carolyn merchant. The Anthropocene and the Humanities: From Climate Change to a New Age of Sustainability. Isis, 113(1):172–175, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gamini:2022:BRS

[486] Amir-Mohammad Gamini. Book review: Sajjad Nikfahm; Fateme Savadi, eds. Nas.īr al-Dīn Muh.ammad al-T .ūsi’s al-Risāla al-Mu’īniyya (The Mu’īniyya Treatise) and Its Supplement. Isis, 113(1):175–177, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Archambeau:2022:BRD

[487] Nicole Archambeau. Book review: Diana Luft. Medieval Welsh Medical Texts. Volume 1: The Recipes. Isis, 113(1):177–178, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Piorko:2022:BRJ

[488] Megan Piorko. Book review: Jennifer M. Rampling. The Experimental Fire: Inventing English Alchemy, 1300-1700.. Isis, 113(1):178–179, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fan:2022:BRR

[489] Fa ti Fan. Book review: Rienk Vermij. Thinking on Earthquakes in Early Modern Europe: Firm Beliefs on Shaky Ground. Isis, 113(1):180–181, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wilde:2022:BRP

[490] Lisa Wilde. Book review: Philip Beeley; Yelda Nasifoglu; Benjamin Wardhaugh, eds. Reading Mathematics in Early Modern Europe: Studies in the Production, Collection, and Use of Mathematical Books. Isis, 113(1):181– 182, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Moser:2022:BRP

[491] Lena Moser. Book review: Phillip Reid. The Merchant Ship in the British Atlantic, 1600-1800: Continuity and Innovation in a Key Technology. Isis, 113(1):183–184, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lund:2022:BRS

[492] Mary Ann Lund. Book review: Susan Wells. Robert Burton’s Rhetoric: An Anatomy of Early Modern Knowledge. Isis, 113(1):184–185, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roos:2022:BRL

[493] Anna Marie Roos. Book review: Lawrence M. Principe. The Transmutations of Chymistry: Wilhelm Homberg and the Academie Royal des Sciences. Isis, 113(1):185–186, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Simmons:2022:BRS

[494] Anna Simmons. Book review: Stefanie Gänger. A Singular Remedy: Cinchona across the Atlantic World, 1750-1820.. Isis, 113(1):186–187, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anthony:2022:BRU

[495] Patrick Anthony. Book review: Ursula Klein. Technoscience in History: Prussia, 1750-1850.. Isis, 113(1):187–189, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kallendorf:2022:BRD

[496] Hilaire Kallendorf. Book review: Dániel Bárth. The Exorcist of Sombor: The Mentality of an Eighteenth-Century Franciscan Friar. Isis, 113(1): 189–190, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Samuel:2022:BRC

[497] Geoffrey Samuel. Book review: C. Pierce Salguero, ed. Buddhism and Medicine: An Anthology of Modern and Contemporary Sources. Isis, 113 (1):190–191, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Durbach:2022:BRE

[498] Nadja Durbach. Book review: Elizabeth A. Williams. Appetite and Its Discontents: Science, Medicine, and the Urge to Eat, 1750-1950.. Isis, 113(1):192–193, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Heggie:2022:BRN

[499] Vanessa Heggie. Book review: Nanna Katrine Lüders Kaalund. Explorations in the Icy North: How Travel Narratives Shaped Arctic Science in the Nineteenth Century. Isis, 113(1):193–194, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pyenson:2022:BRC

[500] Lewis Pyenson. Book review: Christa Jungnickel; Russell McCormmach. The Second Physicist: On the History of Theoretical Physics in Germany. Isis, 113(1):194–196, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bokulich:2022:BRH

[501] Alisa Bokulich and Federica Bocchi. Book review: Henry M. Cowles. The Scientific Method: An Evolution of Thinking from Darwin to Dewey. Isis, 113(1):196–197, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Doroshow:2022:BRE

[502] Deborah Doroshow. Book review: Elliott Bowen. In Search of Sexual Health: Diagnosing and Treating Syphilis in Hot Springs, Arkansas, 1890-1940.. Isis, 113(1):197–199, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cowie:2022:BRM

[503] Helen Cowie. Book review: Mary Anne Andrei. Nature’s Mirror: How Taxidermists Shaped America’s Natural History Museums and Saved Endangered Species. Isis, 113(1):199–200, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fitzgerald:2022:BRJ

[504] Deborah Fitzgerald. Book review: Jonathan Rees. The Chemistry of Fear: Harvey Wiley’s Fight for Pure Food. Isis, 113(1):200–201, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Somsen:2022:BRI

[505] Geert Somsen. Book review: Ilaria Scaglia. The Emotions of Internationalism: Feeling International Cooperation in the Alps in the Interwar Period. Isis, 113(1):202–203, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tonn:2022:BRN

[506] Jenna Tonn. Book review: Nara B. Milanich. Paternity: The Elusive Quest for the Father. Isis, 113(1):203–204, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Meister:2022:BRD

[507] Anna-Maria Meister. Book review: Daniel A. Barber. Modern Architecture and Climate: Design before Air Conditioning. Isis, 113(1):204–205, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Determann:2022:BRJ

[508] Jörg Matthias Determann. Book review: Jacqueline Mitton; Simon Mitton. Vera Rubin: a Life. Isis, 113(1):206–207, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barth:2022:BRJ

[509] Howard G. Barth. Book review: Joachim Schummer; Tom Børsen, eds. Ethics of Chemistry: From Poison Gas to Climate Engineering. Isis, 113 (1):207–208, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Meade:2022:BRM

[510] Ruselle Meade. Book review: Morris Low. Visualizing Nuclear Power in Japan: a Trip to the Reactor. Isis, 113(1):209–210, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McDonald:2022:BRJ

[511] Garret J. McDonald. Book review: Jinee Lokaneeta. The Truth Machines: Policing, Violence, and Scientific Interrogations in India. Isis, 113(1): 210–211, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Whitesides:2022:BRN

[512] Greg Whitesides. Book review: Naomi Oreskes. Science on a Mission: How Military Funding Shaped What We Do and Don’t Know about the Ocean. Isis, 113(1):211–213, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mold:2022:BRI

[513] Alex Mold. Book review: Ido Hartogsohn. American Trip: Set, Setting, and the Psychedelic Experience in the Twentieth Century. Isis, 113(1): 213–214, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Walker:2022:BRK

[514] Paul Walker. Book review: Kenny Cupers; Catharina Gabrielsson; Helena Mattsson, eds. Neoliberalism on the Ground: Architecture and Transformation from the 1960s to the Present. Isis, 113(1):214–215, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lourdusamy:2022:BRP

[515] John B. Lourdusamy. Book review: Pankaj Sekhsaria. Nanoscale: Society’s Deep Impact on Science, Technology, and Innovation in India. Isis, 113(1):216–217, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Baum:2022:BRJ

[516] Emily Baum. Book review: Judith Farquhar; Lili Lai. Gathering Medicines: Nation and Knowledge in China’s Mountain South. Isis, 113 (1):217–218, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ungureanu:2022:BRL

[517] James C. Ungureanu, Claire Brock, Cameron Brinitzer, Etienne Benson, and Stefan Helmreich. Book review: Leonardo Ambasciano. An Unnatural History of Religions, Academia, Post-Truth, and the Quest for Scientific Knowledge. Isis, 113(1):219–220, March 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2022:FCb

[518] Anonymous. Front cover. Isis, 113(2):??, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2022:FMb

[519] Anonymous. Front matter. Isis, 113(2):??, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Espahangizi:2022:SGM

[520] Kijan Espahangizi. Science in glass: Material pathologies in laboratory research, glassware standardization, and the (un)natural history of a modern material, 1900s–1930s. Isis, 113(2):221–244, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Antonello:2022:AKT

[521] Alessandro Antonello. Antarctic krill and the temporalities of oceanic abundance, 1930s–1960s. Isis, 113(2):245–265, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Casper:2022:PDS

[522] Stephen T. Casper. Punch-drunk slugnuts: Violence and the vernacular history of disease. Isis, 113(2):266–288, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Janosikova:2022:USD

[523] Magdaléna Jánosíková. United in scholarship, divided in practice: (re)translating smallpox and measles for Seventeenth-Century Jews. Isis, 113(2):289–309, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bae:2022:EFA

[524] Kyoungjin Bae. Export furniture and artisanal translation in eighteenthcentury canton. Isis, 113(2):310–330, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Flack:2022:DDL

[525] Andrew Flack. Dark degenerations: Life, light, and transformation beneath the Earth, 1840–circa 1900. Isis, 113(2):331–351, June 2022. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Guldi:2022:CED

[526] Jo Guldi. The climate emergency demands a new kind of history: Pragmatic approaches from science and technology studies, text mining, and affiliated disciplines. Isis, 113(2):352–365, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Verburgt:2022:IRH

[527] Lukas M. Verburgt and Elske de Waal. Introduction: Rethinking history of science in the Anthropocene. Isis, 113(2):366–376, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Renn:2022:HSG

[528] Jürgen Renn. From the history of science to geoanthropology. Isis, 113(2): 377–385, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pesa:2022:PAV

[529] Iva Pesa. A planetary anthropocene? Views from Africa. Isis, 113(2): 386–395, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Simon:2022:ESS

[530] Zoltán Boldizsár Simon and Julia Adeney Thomas. Earth system science, Anthropocene historiography, and three forms of human agency. Isis, 113 (2):396–406, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Coen:2022:BHE

[531] Deborah R. Coen and Fredrik Albritton Jonsson. Between history and Earth system science. Isis, 113(2):407–416, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Weinstein:2022:E

[532] Galina Weinstein. To the Editors. Isis, 113(2):417–418, June 2022. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rowe:2022:R

[533] David E. Rowe. In reply. Isis, 113(2):??, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2022:EN

[534] Anonymous. Editors’ note. Isis, 113(2):??, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Siddiqi:2022:WNS

[535] Asif Siddiqi. We need to see things. Isis, 113(2):420–424, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dave:2022:BRK

[536] Kinjal Dave. Book review: Katherine McKittrick. Dear Science and Other Stories. Isis, 113(2):425–426, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pulkkinen:2022:BRP

[537] Karoliina Pulkkinen. Book review: Philip Ball. The Elements: A Visual History of Their Discovery. Isis, 113(2):426–427, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Svatek:2022:BRR

[538] Petra Svatek. Book review: Rasmus Grønfeldt Winther. When Maps Become the World. Isis, 113(2):427–428, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anderson:2022:BRA

[539] Bethany G. Anderson. Book review: Anna Reser; Leila McNeill. Forces of Nature: The Women Who Changed Science. Isis, 113(2):428–429, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DeGroot:2022:BRC

[540] Jean De Groot. Book review: Christopher Byrne. Aristotle’s Science of Matter and Motion. Isis, 113(2):430–431, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marrone:2022:BRD

[541] Steven P. Marrone. Book review: David Pingree. Picatrix: A Medieval Treatise on Astral Magic. Isis, 113(2):431–432, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lewis:2022:BRP

[542] Orly Lewis. Book review: Petros Bouras-Vallianatos. Innovation in Byzantine Medicine: The Writings of John Zacharias Aktouarios (c. 1275–c. 1330). Isis, 113(2):432–433, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rodriguez-Arribas:2022:BRS

[543] J. Rodríguez-Arribas. Book review: Seb Falk. The Light Ages: The Surprising Story of Medieval Science. Isis, 113(2):434–435, June 2022. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Taylor-Poleskey:2022:BRA

[544] Molly Taylor-Poleskey. Book review: Alisha Rankin. The Poison Trials: Wonder Drugs, Experiments, and the Battle for Authority in Renaissance Science. Isis, 113(2):435–436, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Macuglia:2022:BRH

[545] Daniele Macuglia. Book review: Hannah Marcus. Forbidden Knowledge: Medicine, Science, and Censorship in Early Modern Italy. Isis, 113(2): 436–437, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marcos:2022:BRK

[546] Patrícia Martins Marcos. Book review: Kalle Kananoja. Healing Knowledge in Atlantic Africa: Medical Encounters, 1500–1850. Isis, 113(2):

438–439, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gilby:2022:BRT

[547] Emma Gilby. Book review: Timothy M. Harrison. Coming To: Consciousness and Natality in Early Modern England. Isis, 113(2):439–440, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Smith:2022:BRZ

[548] Leonard Smith. Book review: Zachary Dorner. Merchants of Medicines: The Commerce and Coercion of Health in Britain’s Long Eighteenth Century. Isis, 113(2):440–441, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Boantza:2022:BRJ

[549] Victor D. Boantza. Book review: Joan Steigerwald. Experimenting at the Boundaries of Life: Organic Vitality in Germany around 1800. Isis, 113(2):442–443, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Caulkins:2022:BRJ

[550] Tamara Caulkins. Book review: Jeremy Zallen. American Lucifers: The Dark History of Artificial Light, 1750–1865. Isis, 113(2):443–444, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blatchford:2022:BRS

[551] Barrie Blatchford. Book review: Steven Turner. The Science of James Smithson: Discoveries from the Smithsonian Founder. Isis, 113(2):445– 446, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

dAvignon:2022:BRP

[552] Robyn d’Avignon. Book review: Pratik Chakrabarti. Inscriptions of Nature: Geology and the Naturalization of Antiquity. Isis, 113(2):446–447, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ziemer:2022:BRD

[553] Hansjakob Ziemer. Book review: David Trippett; Benjamin Walton, eds. Nineteenth-Century Opera and the Scientific Imagination. Isis, 113(2): 447–449, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Eder:2022:BRE

[554] Sandra Eder. Book review: Elizabeth Reis. Bodies in Doubt: An American History of Intersex. Isis, 113(2):449–450, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Richards:2022:BRB

[555] Evelleen Richards. Book review: Bill Jenkins. Evolution before Darwin: Theories of the Transmutation of Species in Edinburgh, 1804–1834. Isis, 113(2):450–451, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Raia:2022:BRJ

[556] Courtenay Raia. Book review: James Mussell; Graeme Gooday, eds. A Pioneer of Connection: Recovering the Life and Work of Oliver Lodge. Isis, 113(2):452–454, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Naramore:2022:BRR

[557] Sarah E. Naramore. Book review: Rob Boddice. Humane Professions: The Defense of Experimental Medicine, 1876–1914. Isis, 113(2):454–455, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sparks:2022:BRJ

[558] Tabitha Sparks. Book review: Jacob Steere-Williams. The Filth Disease: Typhoid Fever and the Practices of Epidemiology in Victorian England. Isis, 113(2):456–457, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blankholm:2022:BRS

[559] Joseph Blankholm. Book review: Stephen P. Weldon. The Scientific Spirit of American Humanism. Isis, 113(2):457–458, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Park:2022:BRN

[560] Hyung Wook Park. Book review: Neeraja Sankaran. A Tale of Two Viruses: Parallels in the Research Trajectories of Tumor and Bacterial Viruses. Isis, 113(2):458–459, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wolfe:2022:BRM

[561] Audra J. Wolfe. Book review: Mark Solovey. Social Science for What? Battles over Public Funding for the “Other Sciences” at the National Science Foundation. Isis, 113(2):??, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Macuglia:2022:BRG

[562] Daniele Macuglia. Book review: Giovanni Battimelli; Giovanni Ciccotti; Pietro Greco. Computer Meets Theoretical Physics: The New Frontier of Molecular Simulation. Isis, 113(2):461–462, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rogaski:2022:BRX

[563] Ruth Rogaski. Book review: Xiaoping Fang. China and the Cholera Pandemic: Restructuring Society under Mao. Isis, 113(2):462–463, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

MacDonald:2022:BRJ

[564] Helen MacDonald. Book review: John Troyer. Technologies of the Human Corpse. Isis, 113(2):463–465, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kirsh:2022:BRN

[565] Nurit Kirsh. Book review: Noah Tamarkin. Genetic Afterlives: Black Jewish Indigeneity in South Africa. Isis, 113(2):465–466, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hamilton:2022:BRJ

[566] Vivien Hamilton, Stephen T. Casper, Iva Pesa, Zoltán Boldizsár Simon, and Julia Adeney Thomas. Book review: John H. Evans. The Human Gene Editing Debate. Isis, 113(2):466–467, June 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2022:FCc

[567] Anonymous. Front cover. Isis, 113(3):??, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2022:FMc

[568] Anonymous. Front matter. Isis, 113(3):??, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Slagstad:2022:BMC

[569] Ketil Slagstad. Bureaucratizing medicine: Creating a gender identity clinic in the welfare state. Isis, 113(3):469–490, September 2022. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DeBont:2022:MAZ

[570] Raf De Bont. Moving across the zoo-field border: Heini Hediger in Congo. Isis, 113(3):491–512, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anduaga:2022:ENJ

[571] Aitor Anduaga. Epistemic network: The Jesuits and tropical cyclone prediction, 1860–1900. Isis, 113(3):513–536, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Zhang:2022:PRL

[572] Xue Zhang. The plurality of reception: Latitude and longitude in early modern China, 1700–1900. Isis, 113(3):537–558, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Roberts:2022:BEI

[573] Alexandre M. Roberts. Byzantine engagement with Islamicate alchemy. Isis, 113(3):559–580, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kumar:2022:ISH

[574] Prakash Kumar. Introduction: Seeds and the history of science. Isis, 113 (3):581–587, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shopov:2022:FBE

[575] Aleksandar Shopov. Flower breeding in early modern Istanbul: a science of seeds. Isis, 113(3):588–596, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Saraiva:2022:BSA

[576] Tiago Saraiva. Black science: Amílcar Cabral’s agricultural survey and the seeds of African decolonization. Isis, 113(3):597–609, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Curry:2022:HSH

[577] Helen Anne Curry. Hybrid seeds in history and historiography. Isis, 113 (3):610–617, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Laveaga:2022:RSD

[578] Gabriela Soto Laveaga. Rogue seeds in disturbed fields. Isis, 113(3):618– 624, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fullilove:2022:FFV

[579] Courtney Fullilove. “Famine foods” and the values of biodiversity preservation in Israel–Palestine. Isis, 113(3):625–636, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

deChadarevian:2022:HSS

[580] Soraya de Chadarevian and John Krige. History of Science Society annual meeting, 2021. Isis, 113(3):637–644, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Liepold:2022:EMK

[581] Annka Liepold. Eloge: Maya Karin Peterson (1980–2021). Isis, 113(3): 645–646, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ratcliff:2022:BRM

[582] Jessica Ratcliff. Book review: Maia Wellington Gahtan; Eva-Maria Troelenberg, eds. Collecting and Empires: an Historical and Global Perspective. Isis, 113(3):647–648, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Culver:2022:BRS

[583] Lawrence Culver. Book review: Sam White; Christian Pfister; Franz Mauelshagen, eds. The Palgrave Handbook of Climate History. Isis, 113 (3):648–649, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hayes:2022:BRV

[584] Emily Hayes. Book review: Veronica della Dora. The Mantle of the Earth: Genealogies of a Geographical Metaphor. Isis, 113(3):649–651, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Canizares-Esguerra:2022:BRJ

[585] Jorge Cañizares-Esguerra. Book review: Juan Pimentel. Fantasmas de la ciencia española. Isis, 113(3):651–652, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lindemann:2022:BRK

[586] Mary Lindemann. Book review: Kira Robison. Healers in the Making: Students, Physicians, and Medical Education in Medieval Bologna (1250– 1550). Isis, 113(3):652–654, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nielsen:2022:BRP

[587] Wendy C. Nielsen. Book review: Philip Steadman. Renaissance Fun: The Machines behind the Scenes. Isis, 113(3):654–655, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Caulkins:2022:BRW

[588] Tamara Caulkins. Book review: Wolfram Koeppe, ed. Making Marvels: Science and Splendor at the Courts of Europe. Isis, 113(3):655–657, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Scott:2022:BRR

[589] Heidi V. Scott. Book review: Ricardo Padrón. The Indies of the Setting Sun: How Early Modern Spain Mapped the Far East as the Transpacific West. Isis, 113(3):657–658, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barth:2022:BRG

[590] Howard G. Barth. Book review: Gary Patterson. Chemistry in Seventeenth-Century New England. Isis, 113(3):658–659, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Velocci:2022:BRG

[591] Beans Velocci. Book review: Greta LaFleur. The Natural History of Sexuality in Early America. Isis, 113(3):659–660, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Moore:2022:BRS

[592] Alison Downham Moore. Book review: Sean Quinlan. Morbid Undercurrents: Medical Subcultures in Postrevolutionary France. Isis, 113(3):

661–662, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lequin:2022:BRA

[593] Mathilde Lequin. Book review: Andreas Mayer. The Science of Walking: Investigations into Locomotion in the Long Nineteenth Century. Isis, 113 (3):662–663, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pearson:2022:BRD

[594] Jessica Lynne Pearson. Book review: David A. Guba, Jr. Taming Cannabis: Drugs and Empire in Nineteenth-Century France. Isis, 113 (3):663–664, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hoegaerts:2022:BRE

[595] Josephine Hoegaerts. Book review: Edward J. Gillin. Sound Authorities: Scientific and Musical Knowledge in Nineteenth-Century Britain. Isis, 113 (3):664–666, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Blumenthal:2022:BRC

[596] Susanna L. Blumenthal. Book review: Courtney E. Thompson. An Organ of Murder: Crime, Violence, and Phrenology in Nineteenth-Century America. Isis, 113(3):666–667, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Verburgt:2022:BRW

[597] Lukas M. Verburgt. Book review: William J. Ashworth. The Trinity Circle: Anxiety, Intelligence, and Knowledge Creation in Nineteenth-Century England. Isis, 113(3):667–669, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Chajes:2022:BRC

[598] Julie Chajes. Book review: Courtenay Raia. The New Prometheans: Faith, Science, and the Supernatural Mind in the Victorian. Isis, 113 (3):669–670, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Thompson:2022:BRD

[599] Courtney E. Thompson. Book review: Diarmid A. Finnegan. The Voice of Science: British Scientists on the Lecture Circuit in Gilded Age America.

Isis, 113(3):670–671, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ning:2022:BRJ

[600] Zhang Ning. Book review: John Bellamy Foster. The Return of Nature: Socialism and Ecology. Isis, 113(3):672–673, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Benson:2022:BRS

[601] Etienne Benson. Book review: Sara B. Pritchard; Carl A. Zimring. Technology and the Environment in History. Isis, 113(3):673–674, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Camprubi:2022:BRK

[602] Lino Camprubí. Book review: Karl S. Matlin; Jane Maienschein; Rachel A. Ankeny, eds. Why Study Biology by the Sea? Isis, 113(3):674–676, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Moore:2022:BRJ

[603] Taylor M. Moore. Book review: Joelle M. Abi-Rached. ‘As.fūriyyeh: a History of Madness, Modernity, and War in the Middle East. Isis, 113 (3):676–677, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tomar:2022:BRB

[604] Koyna Tomar. Book review: Bharat Jayram Venkat. At the Limits of Cure. Isis, 113(3):677–678, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hudson:2022:BRG

[605] Robert Hudson. Book review: George E. Smith; Raghav Seth. Brownian Motion and Molecular Reality: a Study in Theory-Mediated Measurement. Isis, 113(3):678–680, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gurevitch:2022:BRC

[606] Eric Moses Gurevitch. Book review: Caterina Guenzi. Words of Destiny: Practicing Astrology in North India. Isis, 113(3):680–681, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tomes:2022:BRH

[607] Nancy Tomes. Book review: Harry Yi-Jui Wu. Mad by the Millions: Mental Disorders and the Early Years of the World Health Organization. Isis, 113(3):681–682, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tatarchenko:2022:BRE

[608] Ksenia Tatarchenko. Book review: Elena Aronova. Scientific History: Experiments in History and Politics from the Bolshevik Revolution to the End of the Cold War. Isis, 113(3):682–684, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Merck:2022:BRH

[609] Ashton W. Merck. Book review: Hans Radder. From Commodification to the Common Good: Reconstructing Science, Technology, and Society. Isis, 113(3):684–685, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Herran:2022:BRA

[610] Nestor Herran. Book review: Ana Romero de Pablos. Las primeras centrales nucleares españolas: Actores, políticas y tecnologías. Isis, 113(3): 685–686, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vanderHorst:2022:BRM

[611] Frank C. P. van der Horst, Hannah Murphy, Harold J. Cook, Efthymios Nicolaidis, Eudoxie Delli, Nikolaos Livanos, Kostas Tampakis, George Vlahakis, Tiago Saraiva, Gabriela Soto Laveaga, and Helen Anne Curry. Book review: Marga Vicedo. Intelligent Love: The Story of Clara Park, Her Autistic Daughter, and the Myth of the Refrigerator Mother. Isis, 113 (3):686–687, September 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2022:FCd

[612] Anonymous. Front cover. Isis, 113(4):??, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2022:FMd

[613] Anonymous. Front matter. Isis, 113(4):??, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Grjebine:2022:DMR

[614] Liv Grjebine. A Darwinian murder: The role of the Barré-Lebiez affair in the diffusion of Darwinism in Nineteenth-Century France. Isis, 113 (4):689–709, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sanhueza-Cerda:2022:SLK

[615] Carlos Sanhueza-Cerda. Stabilizing local knowledge: The installation of a meridian circle at the National Astronomical Observatory of Chile (1908– 1913). Isis, 113(4):710–727, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Abou-Nemeh:2022:DCS

[616] Catherine Abou-Nemeh. Daring to conjecture in Seventeenth- and Eighteenth-Century sciences. Isis, 113(4):728–746, December 2022. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tamborini:2022:CMK

[617] Marco Tamborini. The circulation of morphological knowledge: Understanding “Form” across disciplines in the Twentieth and Twenty-First centuries. Isis, 113(4):747–766, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anderson:2022:PHH

[618] Warwick Anderson and James Dunk. Planetary health histories: Toward new ecologies of epidemiology? Isis, 113(4):767–788, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pence:2022:CLE

[619] David Evan Pence. Closing the loop: Ewald von Kleist and the origins of the Leyden jar. Isis, 113(4):789–796, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wu:2022:IRM

[620] Shellen X. Wu. Introduction: Redrawing the map of science in modern China. Isis, 113(4):797–804, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wu:2022:ULU

[621] Shellen X. Wu. Using LoGart to uncover a new spatiality of science in China. Isis, 113(4):805–815, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lavelle:2022:PSA

[622] Peter B. Lavelle. Placing the science of agriculture in Early Twentieth-Century China. Isis, 113(4):816–828, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Zhang:2022:ACN

[623] Jiajing Zhang. The art of compromise: New maps in local gazetteers of the late qing dynasty. Isis, 113(4):829–840, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kohler:2022:EFC

[624] Robert E. Kohler, Lynn K. Nyhart, and Arnold Thackray. Eloge: Frances Coulborn Kohler (1938–2021). Isis, 113(4):841–846, December 2022. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Grafton:2022:ENS

[625] Anthony Grafton and Jed Buchwald. Eloge: Noel Swerdlow (1941–2021). Isis, 113(4):847–853, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fallan:2022:EGO

[626] Kjetil Fallan. Eloge: Gabriele Oropallo (1976–2021). Isis, 113(4):854– 855, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2022:E

[627] Anonymous. Erratum. Isis, 113(4):??, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Abnet:2022:BRT

[628] Dustin Abnet. Book review: Teun Koetsier. The Ascent of GIM, the Global Intelligent Machine: a History of Production and Information Machines. Isis, 113(4):857–858, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Few:2022:BRR

[629] Martha Few. Book review: Richard Jones. Wasp. Isis, 113(4):858–859, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rojas:2022:BRT

[630] Felipe Rojas. Book review: Tobias Esch; Josef Mühlenbrock, eds. Irrtümer und Fälschungen der Archäologie: Begleitband zur Sonderausstellung; Monica Baggio; Elisa Bernard; Monica Salvadori; Luca Zamparo, eds. Anthropology of Forgery: a Multidisciplinary Approach to the Study of Archaeological Fakes; Cécile Michel; Michael Friedrich, eds. Fakes and Forgeries of Written Artefacts from Ancient Mesopotamia to Modern China. Isis, 113(4):859–861, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Oikonomopoulou:2022:BRJ

[631] Katerina Oikonomopoulou. Book review: Jeff Loveland. The European Encyclopedia: From 1650 to the Twenty-First Century. Isis, 113(4):861– 863, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marsella:2022:BRD

[632] Jamie Marsella. Book review: Donald A. Barr. Health Disparities in the United States: Social Class, Race, Ethnicity, and the Social Determinants of Health. Isis, 113(4):863–864, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Irby:2022:BRA

[633] Georgia Irby. Book review: Ágnes Darab. The Anecdotal Narration and Encyclopedic Thought of Pliny the Elder’s “Naturalis Historia”. Isis, 113 (4):865–866, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lang:2022:BRM

[634] Sarah Lang. Book review: Matteo Soranzo. Giovanni Aurelio Augurello (1441–1524) and Renaissance Alchemy: a Critical Edition of “Chrysopoeia” and Other Alchemical Poems, with an Introduction, English Translation and Commentary. Isis, 113(4):866–867, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bigelow:2022:BRM

[635] Alison Bigelow. Book review: Mauricio Nieto Olarte. Exploration, Religion and Empire in the Sixteenth-Century Ibero-Atlantic World: a New Perspective on the History of Modern Science. Isis, 113(4):867–869, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Piorko:2022:BRT

[636] Megan Piorko. Book review: Tara Nummedal. Anna Zieglerin and the Lion’s Blood: Alchemy and End Times in Reformation Germany. Isis, 113 (4):869–870, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Caulkins:2022:BRR

[637] Tamara Caulkins. Book review: Rienk H. Vermij; Ida H. Stamhuis; Fokko Jan Dijksterhuis; C. A. Davids, eds. Rethinking Stevin, Stevin Rethinking: Constructions of a Dutch Polymath. Isis, 113(4):870–871, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Akerman:2022:BRA

[638] James R. Akerman. Book review: Alex Hidalgo. Trail of Footprints: a History of Indigenous Maps from Viceregal Mexico. Isis, 113(4):872–873, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Slaton:2022:BRB

[639] Amy Slaton. Book review: Barbara Hahn. Technology in the Industrial Revolution. Isis, 113(4):873–875, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wahlberg:2022:BRL

[640] Ayo Wahlberg. Book review: Laurence Monnais. The Colonial Life of Pharmaceuticals: Medicines and Modernity in Vietnam. Isis, 113(4):875– 876, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McKay:2022:BRT

[641] Ramah McKay. Book review: Tom Scott-Smith. On an Empty Stomach: Two Hundred Years of Hunger Relief. Isis, 113(4):876–877, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Samanta:2022:BRC

[642] Samiparna Samanta. Book review: Chinmay Tumbe. The Age of Pandemics, 1817–1920: How They Shaped India and the World. Isis, 113 (4):878–879, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Loo:2022:BRN

[643] Tina Loo. Book review: Nancy Langston. Climate Ghosts: Migratory Species in the Anthropocene. Isis, 113(4):879–880, December 2022. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mulberger:2022:BRJ

[644] Annette Mülberger. Book review: James Elwick. Making a Grade: Victorian Examinations and the Rise of Standardized Testing. Isis, 113(4): 881–882, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kennedy:2022:BRA

[645] Devin Kennedy. Book review: Andrew Fiss. Performing Math: a History of Communication and Anxiety in the American Mathematics Classroom. Isis, 113(4):882–883, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dutta:2022:BRS

[646] Arindam Dutta. Book review: Suvobrata Sarkar. Let There Be Light: Engineering, Entrepreneurship, and Electricity in Colonial Bengal, 1880– 1945. Isis, 113(4):884–885, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Abrams:2022:BRT

[647] Ellen Abrams. Book review: Trevor Pearce. Pragmatism’s Evolution: Organism and Environment in American Philosophy. Isis, 113(4):885–886, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Woods:2022:BTB

[648] Rebecca J. H. Woods. Bert theunissen. Beauty or Statistics: Practice and Science in Dutch Livestock Breeding, 1900–2000. Isis, 113(4):887– 888, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fuechtner:2022:BRD

[649] Veronika Fuechtner. Book review: Durba Mitra. Indian Sex Life: Sexuality and the Colonial Origins of Modern Social Thought. Isis, 113(4): 888–889, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Smith:2022:BRM

[650] Sean Morey Smith. Book review: Manuel Barcia. The Yellow Demon of Fever: Fighting Disease in the Nineteenth-Century Transatlantic Slave Trade. Isis, 113(4):889–890, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Radin:2022:BRJ

[651] Joanna Radin. Book review: Jenny Bangham. Blood Relations: Transfusion and the Making of Human Genetics. Isis, 113(4):891–892, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tracy:2022:BRV

[652] Sarah Tracy. Book review: Victoria Lee. The Arts of the Microbial World: Fermentation Science in Twentieth-Century Japan. Isis, 113(4):892–893, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hogarth:2022:BRK

[653] Rana Hogarth. Book review: Keith Wailoo. Pushing Cool: Big Tobacco, Racial Marketing, and the Untold Story of the Menthol Cigarette. Isis, 113 (4):894–895, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hubbard:2022:BRK

[654] Jennifer Hubbard. Book review: Kristin A. Wintersteen. The Fishmeal Revolution: The Industrialization of the Humboldt Current Ecosystem. Isis, 113(4):895–897, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Wolfe:2022:BRJ

[655] Audra J. Wolfe. Book review: Jacob Darwin Hamblin. The Wretched Atom: America’s Global Gamble with Peaceful Nuclear Technology. Isis, 113(4):897–898, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Campos:2022:BRA

[656] Luis A. Campos. Book review: Alex Wellerstein. Restricted Data: The History of Nuclear Secrecy in the United States. Isis, 113(4):898–899, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gannett:2022:BRS

[657] Lisa Gannett. Book review: Soraya de Chadarevian. Heredity under the Microscope: Chromosomes and the Study of the Human Genome. Isis, 113 (4):900–901, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Murphy:2022:CIV

[658] Hannah Murphy, Harold J. Cook, Efthymios Nicolaidis, Eudoxie Delli, Nikolaos Livanos, Kostas Tampakis, and George Vlahakis. Contents of Isis, volume 113. Isis, 113(4):902–904, December 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Murphy:2022:ICB

[659] Hannah Murphy, Harold J. Cook, Efthymios Nicolaidis, Eudoxie Delli, Nikolaos Livanos, Kostas Tampakis, and George Vlahakis. Isis current bibliography of the history of science and its cultural influences 2022. Isis, 113(S1):1–266, ???? 2022. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2023:FCa

[660] Anonymous. Front cover. Isis, 114(1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2023:FMa

[661] Anonymous. Front matter. Isis, 114(1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hui:2023:EN

[662] Alexandra Hui and Matthew Lavine. Editors’ note. Isis, 114(1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Terrall:2023:AIF

[663] Mary Terrall. African indigo in the French Atlantic: Michel Adanson’s encounter with Senegal. Isis, 114(1):2–24, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Clever:2023:BAF

[664] Iris Clever. Biometry against fascism: Geoffrey morant, race, and antiracism in Twentieth-Century Physical anthropology. Isis, 114(1):25–49, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Secord:2023:ISR

[665] James A. Secord. Inventing the scientific revolution. Isis, 114(1):50– 76, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Madison:2023:TWB

[666] Paige Madison. Tug-of-war: Bones and stones as scientific objects in postcolonial Indonesia. Isis, 114(1):77–98, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Xie:2023:APO

[667] Xin zhe Xie. Administration of perception: Observing and transcribing dead bodies in the forensic methodology of Qing China (1644–1912). Isis, 114(1):99–122, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

dellaDora:2023:RSC

[668] Veronica della Dora. From the Radio Shack to the Cosmos: Listening to Sputnik during the International Geophysical Year (1957–1958). Isis, 114(1):123–149, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Park:2023:MOS

[669] Katharine Park. The myth of the “One-Sex” body. Isis, 114(1):150– 175, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bengtsson:2023:EPF

[670] Keith R. Bengtsson and Kristin Johnson. Eloge: Paul Farber (1944–2021). Isis, 114(1):176–181, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hanson:2023:ENS

[671] Marta Hanson, Michael Nylan, and Hilary A. Smith. Eloge: Nathan Sivin (1931–2022). Isis, 114(1):182–186, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ambasciano:2023:LE

[672] Leonardo Ambasciano. Letter to the Editors. Isis, 114(1):187–188, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schmidt:2023:PG

[673] Susanne Schmidt. The prenatal gaze. Isis, 114(1):189–192, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lang:2023:BRL

[674] Philippa Lang. Book review: Liba Taub, ed. The Cambridge Companion to Ancient Greek and Roman Science. Isis, 114(1):193–194, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pearl:2023:BRL

[675] Sharrona Pearl. Book review: Lisa Devriese, ed. The Body as a Mirror of the Soul: Physiognomy from Antiquity to the Renaissance. Isis, 114(1): 194–195, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Yale:2023:BRS

[676] Elizabeth Yale. Book review: Sarah Neville. Early Modern Herbals and the Book Trade: English Stationers and the Commodification of Botany. Isis, 114(1):195–196, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

DeAngelis:2023:BRM

[677] Simone De Angelis. Book review: Martin Mulsow, ed. Urban Gottfried Bucher (1679–1724). Isis, 114(1):197–198, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Duncan:2023:BRL

[678] Ian Duncan. Book review: Lynn Festa. Fiction without Humanity: Person, Animal, Thing in Early Enlightenment Literature and Culture. Isis, 114(1):198–199, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Suzuki:2023:BRW

[679] Akihito Suzuki. Book review: Wei Yu Wayne Tan. Blind in Early Modern Japan: Disability, Medicine, and Identity. Isis, 114(1):199–200, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Menon:2023:BRP

[680] Minakshi Menon. Book review: Paul A. Elliot. Erasmus Darwin’s Gardens: Medicine, Agriculture and the Enlightenment Sciences. Isis, 114(1): 201–202, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Naramore:2023:BRS

[681] Sarah Naramore. Book review: Susan H. Brandt. Women Healers: Gender, Authority, and Medicine in Early Philadelphia. Isis, 114(1):202–203, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ramirez:2023:BRC

[682] Paul Ramírez. Book review: Christina Ramos. Bedlam in the New World: a Mexican Madhouse in the Age of Enlightenment. Isis, 114(1):203–204, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Levine:2023:BRK

[683] Philippa Levine. Book review: Kristin D. Hussey. Imperial Bodies in London: Empire, Mobility, and the Making of British Medicine, 1880– 1914. Isis, 114(1):205–206, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Simpson:2023:BRL

[684] Thomas Simpson. Book review: Lachlan Fleetwood. Science on the Roof of the World: Empire and the Remaking of the Himalaya. Isis, 114(1): 206–207, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Myburgh:2023:BRA

[685] Brittany Myburgh. Book review: Andrei Pop. A Forest of Symbols: Art, Science, and Truth in the Long Nineteenth Century. Isis, 114(1):207– 208, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bix:2023:BRK

[686] Amy Sue Bix. Book review: Kat Jungnickel. Bikes and Bloomers: Victorian Women Inventors and Their Extraordinary Cycle Wear. Isis, 114(1): 209–210, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shelton:2023:BRJ

[687] Laura M. Shelton. Book review: Jennifer Koshatka Seman. Borderlands Curanderos: The Worlds of Santa Teresa Urrea and Don Pedrito Jaramillo. Isis, 114(1):210–211, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Willoughby:2023:BRJ

[688] Urmi Engineer Willoughby. Book review: Jeannie N. Shinozuka. Biotic Borders: Transpacific Plant and Insect Migration and the Rise of Anti-Asian Racism in America, 1890–1950. Isis, 114(1):211–213, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barth:2023:BRC

[689] Howard G. Barth. Book review: Carmen J. Giunta; Vera V. Mainz; Gregory S. Girolami, eds. 150 Years of the Periodic Table: a Commemorative Symposium, (Perspectives on the History of Chemistry.) 460 pp., illus., index. Berlin: Springer, 2021. $149.99 (cloth); ISBN 978-3-030-67909-5. Isis, 114(1):213–214, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Soentgen:2023:BRA

[690] Jens Soentgen. Book review: Alexander Klose; Benjamin Steininger. Erdöl: Ein Atlas der Petromoderne. Isis, 114(1):214–215, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Podgorny:2023:BRC

[691] Irina Podgorny. Book review: Caitlin Donahue Wylie. Preparing Dinosaurs: The Work behind the Scenes. Isis, 114(1):215–216, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Koester:2023:BRF

[692] Elizabeth Koester. Book review: Frank W. Stahnisch; Erna Kurbegović, eds. Psychiatry and the Legacies of Eugenics: Historical Studies of Alberta and Beyond, 488 pp., app., notes, bibl., index. Edmonton: AU Press, 2020. $41.95 (paper); ISBN 978-1-77199-265-7. Isis, 114(1):217–218, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Mart:2023:BRP

[693] Michelle Mart. Book review: E. Paul Durrenberger. The Dawn of Industrial Agriculture in Iowa: Anthropology, Literature, and History. Isis, 114(1):218–219, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Petrov:2023:BRT

[694] Victor Petrov. Book review: Thomas Haigh; Paul E. Ceruzzi. A New History of Modern Computing, 544 pp., notes, bibl., index. Cambridge, Mass./London: MIT Press, 2021. $40 (paper); ISBN 978-0-262-54290-6.

Isis, 114(1):219–220, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Schlombs:2023:BRD

[695] Corinna Schlombs. Book review: Douglas M. O’Reagan. Taking Nazi Technology: Allied Exploitation of German Science after the Second World War, 296 pp., notes, bibl., index. Baltimore: Johns Hopkins University Press, 2021. $34.95 (paper); ISBN 978-1-4214-3984-6. Isis, 114(1):220– 222, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Suarez-Diaz:2023:BRE

[696] Edna Suárez-Díaz. Book review: Elise K. Burton. Genetic Crossroads: The Middle East and the Science of Human Heredity. Isis, 114(1):222– 223, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kaplan:2023:BRR

[697] Judith Kaplan. Book review: Randy Allen Harris. The Linguistics Wars: Chomsky, Lakoff, and the Battle over Deep Structure. Isis, 114(1):223– 224, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bhattacharya:2023:BRS

[698] Arnav Bhattacharya. Book review: Steven Epstein. The Quest for Sexual Health: How an Elusive Ideal Has Transformed Science, Politics, and Everyday Life. Isis, 114(1):225–226, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Durrani:2023:BRH

[699] Haris A. Durrani. Book review: Hugh R. Slotten. Beyond Sputnik and the Space Race: The Origins of Global Satellite Communications. Isis, 114(1): 226–227, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sutter:2023:BRC

[700] Paul S. Sutter. Book review: Christine Keiner. Deep Cut: Science, Power, and the Unbuilt Interoceanic Canal. Isis, 114(1):227–228, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rogers:2023:BRJ

[701] Emily Lim Rogers. Book review: Jan Nisbet. Pain and Shock in America: Politics, Advocacy, and the Controversial Treatment of People with Disabilities. Isis, 114(1):229–230, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Charenko:2023:BRR

[702] Melissa Charenko. Book review: Roy Plotnick. Explorers of Deep Time: Paleontologists and the History of Life. Isis, 114(1):230–231, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hanemaayer:2023:BRE

[703] Ariane Hanemaayer, Hannah Murphy, By Harold J. Cook, Efthymios Nicolaidis, Eudoxie Delli, Nikolaos Livanos, Kostas Tampakis, and George Vlahakis. Book review: Emma Bedor Hiland. Therapy Tech: The Digital Transformation of Mental Healthcare. Isis, 114(1):231–232, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2023:FCb

[704] Anonymous. Front cover. Isis, 114(2):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2023:FMb

[705] Anonymous. Front matter. Isis, 114(2):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Robinson:2023:WWW

[706] Morgan J. Robinson. When a wonder is not a wonder: Swahili, translation, and the communication of knowledge. Isis, 114(2):233–248, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Donaghy:2023:SPE

[707] Paige Donaghy. The secrets of the placenta in European anatomy and midwifery, 1560-1700. Isis, 114(2):249–271, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Degroot:2023:OSS

[708] Dagomar Degroot. One small step for man, one giant leap for Moon microbes? Interpretations of risk and the limits of quarantine in NASA’s Apollo Program. Isis, 114(2):272–298, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Joseph:2023:WNM

[709] Jacques Joseph. Whither natural magic? Science, witchcraft, and the decline of magic in Henry More. Isis, 114(2):299–316, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Parry:2023:DBA

[710] Alexander I. Parry. Delivering bacteriology to the American homemaker: Correspondence education, kitchen experiments, and public health, 1890-1930. Isis, 114(2):317–340, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Raphael:2023:TCT

[711] Renée Raphael. Toward a critical transatlantic history of early modern mining: Depiction, reality, and readers’ expectations in Álvaro Alonso Barba’s 1640 El arte de los metales. Isis, 114(2):341–358, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tkaczyk:2023:IRR

[712] Viktoria Tkaczyk and Christine von Oertzen. Introduction: Reconsidering the resources of epistemic tools. Isis, 114(2):359–365, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gribenski:2023:SAP

[713] Fanny Gribenski and David Pantalony. Sounding acoustic precision: Tuning forks and cast steel’s Nineteenth-Century Euro–American networks. Isis, 114(2):366–372, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tkaczyk:2023:ESA

[714] Viktoria Tkaczyk. Early sound archiving and the making of scientific resources. Isis, 114(2):373–379, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

vonOertzen:2023:PKS

[715] Christine von Oertzen. Paper knowledge and statistical precision. Isis, 114(2):380–386, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Felten:2023:MS

[716] Sebastian Felten. Money supplies. Isis, 114(2):387–392, ???? 2023. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Dupre:2023:NGT

[717] Sven Dupré. The nature of glass: Technologies of transparency, materials on the move. Isis, 114(2):393–399, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marchand:2023:KSM

[718] Suzanne Marchand. On kaolin and its substitutes; or, the management of whiteness. Isis, 114(2):400–406, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Werrett:2023:DSU

[719] Simon Werrett. Diverse shapes: Used goods as material resources in early modern sciences. Isis, 114(2):407–412, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sterne:2023:ETP

[720] Jonathan Sterne. Eloge: Trevor Pinch (1952–2021). Isis, 114(2):413– 415, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Tresch:2023:EBL

[721] John Tresch. Eloge: Bruno Latour (1947–2022). Isis, 114(2):416–420, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Cohen:2023:OMS

[722] H. Floris Cohen. The origins of modern science in the classroom. Isis, 114(2):421–423, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kingsland:2023:BRM

[723] Sharon Kingsland. Book review: Michel Morange; translated by Teresa Lavender Fagan and Joseph Muise. A History of Biology. Isis, 114(2): 424–425, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Murphy:2023:PVC

[724] Hannah Murphy. Book review: Paula S. De Vos. Compound Remedies: Galenic Pharmacy from the Ancient Mediterranean to New Spain. Isis, 114(2):425–426, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stuart:2023:JCB

[725] Michael T. Stuart. Book review: Jimena Canales. Bedeviled: A Shadow History of Demons in Science. Isis, 114(2):426–427, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lawrence-Mathers:2023:SSE

[726] Anne Lawrence-Mathers. Book review: Sacha Stern, ed. Calendars in the Making: The Origins of Calendars from the Roman Empire to the Later Middle Ages. Isis, 114(2):428–429, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lehleiter:2023:MCP

[727] Christine Lehleiter. Book review: Mackenzie Cooley. The Perfection of Nature: Animals, Breeding, and Race in the Renaissance. Isis, 114(2): 429–430, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sauter:2023:GOA

[728] Michael J. Sauter. Book review: Günther Oestmann. The Astronomical Clock of Strasbourg Cathedral: Function and Significance. Isis, 114(2): 431–432, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lynn:2023:MWM

[729] Michael R. Lynn. Book review: Mark A. Waddell. Magic, Science, and Religion in Early Modern Europe. Isis, 114(2):432–433, ???? 2023. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Carolino:2023:VNB

[730] Luís Miguel Carolino. Book review: Víctor Navarro Brotons. Jerónimo Muñoz: Matemáticas, cosmología y humanismo en la época del Renacimiento. Isis, 114(2):433–434, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fara:2023:AWM

[731] Patricia Fara. Book review: Alexander Wragge-Morley. Aesthetic Science: Representing Nature in the Royal Society of London, 1650–1720. Isis, 114 (2):434–436, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Buning:2023:GGS

[732] Marius Buning. Book review: Graeme Gooday; Steven Wilf, eds. Patent Cultures: Diversity and Harmonization in Historical Perspective. Isis, 114(2):436–437, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lindsay:2023:CIR

[733] Debra Lindsay. Book review: Christoph Irmscher; Richard J. King, eds. Audubon at Sea: The Coastal and Transatlantic Adventures of John James Audubon. Isis, 114(2):437–438, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bruisch:2023:CGG

[734] Katja Bruisch. Book review: Catherine Gibson. Geographies of Nationhood: Cartography, Science, and Society in the Russian Imperial Baltic. Isis, 114(2):438–439, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sen:2023:EIH

[735] A. Tunç Sen. Book review: Ekmeleddin Ihsanoglu. The House of Sciences: The First Modern University in the Muslim World. Isis, 114(2): 440–441, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hale:2023:RMW

[736] Piers J. Hale. Book review: Roger M. White; M. J. S. Hodge; Gregory Radick. Darwin’s Argument by Analogy: From Artificial to Natural Selection. Isis, 114(2):441–442, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McCook:2023:LKW

[737] Stuart McCook. Book review: Luke Keogh. The Wardian Case: How a Simple Box Moved Plants and Changed the World. Isis, 114(2):442– 443, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Chinnici:2023:OWN

[738] Ileana Chinnici. Book review: Omar W. Nasim. The Astronomer’s Chair: a Visual and Cultural History. Isis, 114(2):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Teter:2023:HJK

[739] Magda Teter. Book review: Hillel J. Kieval. Blood Inscriptions: Science, Modernity, and Ritual Murder at Europe’s Fin de Siècle. Isis, 114(2): 444–446, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Maerker:2023:TCC

[740] Anna Maerker. Book review: Tinne Claes. Corpses in Belgian Anatomy, 1860-1914: Nobody’s Dead. Isis, 114(2):446–447, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

An:2023:NKM

[741] Sooyoung An. Book review: Nicholas K. Menzies. Ordering the Myriad Things: From Traditional Knowledge to Scientific Botany in China. Isis, 114(2):447–449, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

LaFay:2023:MMS

[742] Elaine LaFay. Book review: Martin Mahony; Samuel Randalls, eds. Weather, Climate, and the Geographical Imagination: Placing Atmospheric Knowledges. Isis, 114(2):449–450, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bras:2023:CRD

[743] Anatole Le Bras. Book review: Camille Robcis. Disalienation: Politics, Philosophy, and Radical Psychiatry in Postwar France. Isis, 114(2):450– 451, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Minor:2023:GWS

[744] Adriana Minor. Book review: Greg Whitesides. Science and American Foreign Relations since World War II. Isis, 114(2):452–453, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Loh:2023:JRI

[745] Kah Seng Loh. Book review: Jo Robertson. The International Campaign against Leprosy, 1948–2005. Isis, 114(2):453–454, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Reverby:2023:SHD

[746] Susan M. Reverby. Book review: Sydney A. Halpern. Dangerous Medicine: The Story behind Human Experiments with Hepatitis. Isis, 114(2):455– 456, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McDonald:2023:TKA

[747] Garret J. McDonald. Book review: Togzhan Kassenova. Atomic Steppe: How Kazakhstan Gave Up the Bomb. Isis, 114(2):456–457, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Domingues:2023:EKW

[748] Amanda Domingues. Book review: Emily K. Wilson. Mildred Trotter and the Invisible Histories of Physical and Forensic Anthropology. Isis, 114(2): 457–458, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Scholthof:2023:GJM

[749] Karen-Beth G. Scholthof. Book review: Gregory J. Morgan. Cancer Virus Hunters: a History of Tumor Virology. Isis, 114(2):459–460, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Steingart:2023:JAS

[750] Alma Steingart. Book review: Janet Abbate; Stephanie Dick, eds. Abstractions and Embodiments: New Histories of Computing and Society. Isis, 114(2):460–461, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hersch:2023:AJC

[751] Matthew Hersch. Book review: Andrew L. Jenks. Collaboration in Space and the Search for Peace on Earth. Isis, 114(2):461–462, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Aronowitz:2023:JGD

[752] Robert Aronowitz. Book review: Jeremy A. Greene. The Doctor Who Wasn’t There: Technology, History, and the Limits of Telehealth. Isis, 114(2):463–464, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kessler:2023:MPM

[753] Elizabeth A. Kessler. Book review: Maura Phillips Mackowski. Life in Space: NASA Life Sciences Research during the Late Twentieth Century. Isis, 114(2):464–465, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ray:2023:EDJ

[754] Subhadeepta Ray. Book review: Elizabeth D. Jones. Ancient DNA: The Making of a Celebrity Science. Isis, 114(2):466–467, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Widmalm:2023:OHC

[755] Sven Widmalm, Hannah Murphy, By Harold J. Cook, Efthymios Nicolaidis, Eudoxie Delli, Nikolaos Livanos, Kostas Tampakis, and George Vlahakis. Book review: Olof Hallonsten. The Campaign: How a European Big Science Facility Ended Up on the Peripheral Farmlands of Southern Sweden. Isis, 114(2):467–468, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2023:FCc

[756] Anonymous. Front cover. Isis, 114(3):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2023:FMc

[757] Anonymous. Front matter. Isis, 114(3):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Moore:2023:NHT

[758] Taylor M. Moore. An (un)Natural history: Tracing the magical rhinoceros Horn in Egypt. Isis, 114(3):469–489, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Luo:2023:AFA

[759] Anin Luo. Animals, film, audiences: Regulating cruelty and morality through science and law in interwar Britain. Isis, 114(3):490–512, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Rich:2023:MMS

[760] Miriam Rich. Monstrosity in medical science: Race-making and teratology in the Nineteenth-Century United States. Isis, 114(3):513–536, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Gillin:2023:SCS

[761] Edward J. Gillin and Fanny Gribenski. Science in the Church: The sacred spaces of sonorous experiment and the formation of modern acoustics. Isis, 114(3):537–558, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barr:2023:EMP

[762] Zachary Barr. Ernst Mach’s popular science. Isis, 114(3):559–577, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lucier:2023:CCC

[763] Oliver Lucier. Climate conscious: Caribbean commodities and Holdridge life zones, 1940s-1960s. Isis, 114(3):578–598, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Felten:2023:EMR

[764] Sebastian Felten and Renée Raphael. Early modern resources: An introduction. Isis, 114(3):599–603, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barnett:2023:EPE

[765] Lydia Barnett. Eco-prospecting in early modern wetlands. Isis, 114(3): 604–610, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Chen:2023:EPM

[766] BuYun Chen. “To embrace and protect”: Managing wind, water, and trees in the Ryukyu Kingdom. Isis, 114(3):611–618, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

deAvila:2023:NPE

[767] Nydia Pineda de Ávila. Negotiating pearls in the early Jesuit missions of California. Isis, 114(3):619–625, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Felten:2023:MMG

[768] Sebastian Felten. Managing mineral growth in early modern mining. Isis, 114(3):626–630, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bouley:2023:PMF

[769] Bradford Bouley. A plague of meat: Food, politics, and warfare in Early Modern Italy. Isis, 114(3):631–637, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Raphael:2023:MMC

[770] Renée Raphael. Mining mercury for the common good: Debating the public good and wealth in Huancavelica. Isis, 114(3):638–645, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Opitz:2023:HSS

[771] Donald L. Opitz and Jaipreet Virdi. History of science society annual meeting, 2022. Isis, 114(3):646–653, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Jarrick:2023:RBW

[772] Arne Kent Jarrick. Book review: Rens Bod. World of Patterns: A Global History of Knowledge, translated by Leston Buell. Isis, 114(3):654–655, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Donato:2023:NTD

[773] Maria Pia Donato. Book review: Neil Tarrant. Defining Nature’s Limits: The Roman Inquisition and the Boundaries of Science. Isis, 114(3):655– 657, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Swan:2023:JND

[774] Claudia Swan. Book review: Jennifer Nelson. Disharmony of the Spheres: The Europe of Holbein’s “Ambassadors”. Isis, 114(3):657–658, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Echterholter:2023:MAB

[775] Anna Echterhölter. Book review: Marisa Anne Bass; Anne Goldgar; Hanneke Grootenboer; Claudia Swan. Conchophilia: Shells, Art, and Curiosity in Early Modern Europe. Isis, 114(3):658–660, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Amster:2023:JKS

[776] Ellen Amster. Book review: Justin K. Stearns. Revealed Sciences: The Natural Sciences in Islam in Seventeenth-Century Morocco. Isis, 114(3): 660–661, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Serrano:2023:MCM

[777] Elena Serrano. Book review: Montserrat Cabré; María Cruz de Carlos, eds. Maria Sibylla Merian y Alida Withoos: Mujeres, Arte y Ciencia en la Edad Moderna. Isis, 114(3):661–662, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Bond:2023:PDG

[778] Dean W. Bond. Book review: Phil Dodds. The Geographies of Enlightenment Edinburgh. Isis, 114(3):662–663, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Walsh:2023:APR

[779] Sarah Walsh. Book review: Alexander Pavuk. Respectably Catholic and Scientific: Evolution and Birth Control between the World Wars. Isis, 114(3):664–665, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Marshall:2023:MCH

[780] Nancy Rose Marshall. Book review: Matthew C. Hunter. Painting with Fire: Sir Joshua Reynolds, Photography, and the Temporally Evolving Chemical Object. Isis, 114(3):665–666, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sharafi:2023:CCS

[781] Mitra Sharafi. Book review: E. Claire Cage. The Science of Proof: Forensic Medicine in Modern France. Isis, 114(3):666–668, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Pols:2023:MBG

[782] Hans Pols. Book review: Moritz von Brescius. German Science in the Age of Empire: Enterprise, Opportunity and the Schlagintweit Brothers. Isis, 114(3):668–669, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Barton:2023:RJN

[783] Ruth Barton. Book review: Roland Jackson; Nicola Jackson; Daniel Brown, eds. The Poetry of John Tyndall. Isis, 114(3):669–670, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Coleman:2023:MWM

[784] Deirdre Coleman. Book review: Matthew Wale. Making Entomologists: How Periodicals Shaped Scientific Communities in Nineteenth-Century Britain. Isis, 114(3):670–671, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Varanda:2023:SCP

[785] Jorge Varanda. Book review: Samuël Coghe. Population Politics in the Tropics: Demography, Health, and Transimperialism in Colonial Angola. Isis, 114(3):672–673, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Runcie:2023:GLC

[786] Sarah Runcie. Book review: Guillaume Lachenal; Cheryl Smeall. The Doctor Who Would Be King. Isis, 114(3):673–674, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Shen:2023:VSC

[787] Grace Yen Shen. Book review: Victor Seow. Carbon Technocracy: Energy Regimes in Modern East Asia. Isis, 114(3):674–676, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Lorenat:2023:KHP

[788] Jemma Lorenat. Book review: Karen Hunger Parshall. The New Era in American Mathematics, 1920–1950. Isis, 114(3):676–677, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Curry:2023:MRB

[789] Helen Anne Curry. Book review: Marci R. Baranski. The Globalization of Wheat: a Critical History of the Green Revolution. Isis, 114(3):677– 679, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Freeborn:2023:NJE

[790] Alfred Freeborn. Book review: Noortje Jacobs. Ethics by Committee: a History of Reasoning Together about Medicine, Science, Society, and the State. Isis, 114(3):679–680, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kumbhar:2023:RTS

[791] Kiran Kumbhar. Book review: Renny Thomas. Science and Religion in India: Beyond Disenchantment. Isis, 114(3):680–681, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Adler:2023:SMO

[792] Antony Adler. Book review: Samantha Muka. Oceans under Glass: Tank Craft and the Sciences of the Sea. Isis, 114(3):681–683, ???? 2023. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Finkelstein:2023:SPD

[793] Maura Finkelstein. Book review: Sarah Pinto. Daughters of Parvati: Women and Madness in Contemporary India. Isis, 114(3):683–684, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Siddiqi:2023:JVS

[794] Asif Siddiqi, Hannah Murphy, By Harold J. Cook, Efthymios Nicolaidis, Eudoxie Delli, Nikolaos Livanos, Kostas Tampakis, and George Vlahakis. Book review: Janet Vertesi. Shaping Science: Organizations, Decisions, and Culture on NASA’s Teams. Isis, 114(3):684–685, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2023:FCd

[795] Anonymous. Front cover. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2023:FMd

[796] Anonymous. Front matter. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Anonymous:2023:C

[797] Anonymous. Contents. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sankaran:2023:A

[798] Neeraja Sankaran and Stephen P. Weldon. Acknowledgments. Isis, 114 (S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Sankaran:2023:STC

[799] Neeraja Sankaran and Stephen P. Weldon. Scholarship in the time of COVID-19: An introduction to the IsisCB special issue on pandemics. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Engelmann:2023:SIE

[800] Lukas Engelmann. A short introduction into the English-language historiography of epidemiology. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

daSilva:2023:EID

[801] Matheus Alves Duarte da Silva and Jules Skotnes-Brown. Emerging infectious diseases and disease emergence: Critical, ontological and epistemological approaches. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Vargha:2023:VP

[802] Dora Vargha and Imogen Wilkins. Vaccination and pandemics. Isis, 114 (S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Engelmann:2023:CCS

[803] Lukas Engelmann. Coinfection, comorbidity, and syndemics: On the edges of epidemic historiography. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Stark:2023:MMT

[804] James Stark. Making microbes: Theorizing the invisible in historical scholarship. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Canavan:2023:HLR

[805] Barbara Canavan. Historical literature related to zoonoses and pandemics. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Honigsbaum:2023:SFP

[806] Mark Honigsbaum. The “Spanish” flu and the pandemic imaginary. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Kanazawa:2023:PRS

[807] Reiko Kanazawa. Pandemic responses and the strengths of health systems: a review of global AIDS historiography in light of COVID-19. Isis, 114 (S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

McGovern:2023:EIS

[808] Michael F. McGovern and Keith A. Wailoo. Epidemic inequities: Social and racial inequality in the history of pandemics. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Parisi:2023:LLR

[809] Valentina Parisi and Kavita Sivaramakrishnan. The limits of linearity: Recasting histories of epidemics in the global south. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Flemming:2023:PAM

[810] Rebecca Flemming. Pandemics in the Ancient Mediterranean World. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Varlik:2023:PMI

[811] Nükhet Varlik. Plague in the Mediterranean and Islamicate World. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Peckham:2023:EHE

[812] Robert Peckham and Mei Li. Epidemic histories in East Asia. Isis, 114 (S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Neelakantan:2023:HPS

[813] Vivek Neelakantan. History of pandemics in Southeast Asia: a return of national anxieties? Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Chakraborty:2023:CRS

[814] Arnab Chakraborty. COVID-19 response in South Asia: Case studies from India, Sri Lanka, and Pakistan. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Diener:2023:EPP

[815] Leander Diener and Flurin Condrau. The European perspective on pandemics. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Ragas:2023:HPL

[816] José Ragas. History of pandemics in Latin America. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Conforti:2023:HEB

[817] Maria Conforti. History of epidemics: a bibliographical essay on secondary sources in Italian and on Italy. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Fangerau:2023:SHW

[818] Heiner Fangerau, Ulrich Koppitz, and Alfons Labisch. A survey of historical works on pandemics in the German language. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Hamilton:2023:CWI

[819] Emily Hamilton, Hannah Murphy, Harold J. Cook, Efthymios Nicolaidis, Eudoxie Delli, Nikolaos Livanos, Kostas Tampakis, and George Vlahakis. Conclusion: What it means. Isis, 114(S1):??, ???? 2023. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Guichelaar:2018:WSE

[820] Jan Guichelaar. Willem de Sitter: Einstein’s Friend and Opponent. Springer Biographies. Springer International Publishing, Cham, Switzerland, 2018. ISBN 3-319-98336-9 (hardcover), 3-319-98337-7 (e-book), 3-319-98338-5. ISSN 2365-0613 (print), 2365-0621 (electronic). xxii + 278 + 71 pp. LCCN QB36.S6 G85 2018. URL http://link.springer.com/

10.1007/978-3-319-98337-0.